### v1

# Mental Disorders Detection in the Era of Large Language Models

Abstract
This paper compares the effectiveness of traditional machine learning methods, encoder-based models, and large language models (LLMs) on the task of detecting depression and anxiety. Five datasets were considered, each differing in format and the method used to define the target pathology class. We tested AutoML models based on linguistic features, several variations of encoder-based Transformers such as BERT, and state-of-the-art LLMs as pathology classification models. The results demonstrated that LLMs outperform traditional methods, particularly on noisy and small datasets where training examples vary significantly in text length and genre. However, psycholinguistic features and encoder-based models can achieve performance comparable to language models when trained on texts from individuals with clinically confirmed depression, highlighting their potential effectiveness in targeted clinical applications.
Mental Disorders Detection in the Era of Large Language Models
Gleb Kuzmin2,4 Petr Strepetov3 Maksim Stankevich4 Ivan Smirnov4 Artem Shelmanov1 1MBZUAI 2AIRI 3MIPT 4FRC CSC RAS kuzmin@airi.net strepetov.pa@phystech.edu stankevich@isa.ru ivs@isa.ru artem.shelmanov@mbzuai.ac.ae
1 Introduction
The problem of detecting mental disorders and patient emotions through text analysis and machine learning has been of increasing interest to researchers over the past decade (Graham et al., 2019; Zhang et al., 2022; Calixto et al., 2022; Mayer et al., 2024). In fact, advances in data science and natural language processing methods offer promising opportunities for screening, monitoring, early detection, and prevention of negative outcomes of mental disorders. Although there are studies that work with interviews (Morales and Levitan, 2016; Ringeval et al., 2017) and offline texts (Lynn et al., 2018; Stankevich et al., 2019), in most cases the material for such research comes from social media (Guntuku et al., 2017; Garg, 2023). These studies tend to focus on but are not limited to, conditions such as depression, anxiety, stress, suicidality, post-traumatic stress disorder, and anorexia. Unsurprisingly, the methods considered for predicting mental state from text fell into traditional machine learning, using hand-crafted linguistic features, and various forms of deep learning (Zhang et al., 2022). The deep learning approach is often more accurate, especially when there are enough data samples, while traditional machine learning produces more interpretative results.
In this paper, we compare the performance of linguistic features, encoder-based models, and large language models (LLMs) on the task of identifying mental disorders. We consider two types of mental states, depression and anxiety, and several datasets in Russian that differ in text format and in the way a pathology is detected.
We are particularly interested in evaluating the performance on the essay dataset because it consists of texts written by patients with clinically diagnosed depression and texts written by volunteers who were considered healthy (Stankevich et al., 2019). It is worth pointing out the problem of dividing data into texts written by people with a certain mental pathology. In most cases, social media-based studies use self-report, affiliation, or questionnaire data to determine mental health status. However, there is a lack of work using clinically validated data, which can vary widely (Chancellor and De Choudhury, 2020; Ernala et al., 2019). We believe that verified training data from essays can significantly contribute to building models for the social media domain.
This paper addresses the following research questions:
RQ1: What is the most prominent technique for predicting depression and anxiety: traditional machine learning, encoder-based models, or recent LLMs?
RQ2: Do models trained on a dataset of essays in which depression was defined by a clinical diagnosis generalize to social media, where the depression status is defined by a questionnaire?
Our main contributions are the following:
We outperformed the existing state-of-the-art depression detection method on one dataset and established classification baselines on three previously unexamined anxiety datasets.
We conducted a thorough comparison of various groups of models on the depression and anxiety detection tasks in Russian, which could be used for the practitioners in this field for future experiments.
We investigated the transferability of models from the task with clinical diagnosis as a target to the task with a target obtained as a questionnaire result to alleviate the lack of clinically validated data in the mental disorder detection task.
2 Related Work
2.1 Traditional and Advanced ML Methods
Researchers have employed diverse methodologies and techniques in a comprehensive exploration of depression and anxiety detection across various social media platforms. Tadesse et al. (2019) consider the Reddit users’ dataset, comparing single and combined feature learning for depression detection. N-gram features, Linguistic Inquiry and Word Count (LIWC) dictionary features, and topics from Latent Dirichlet Allocation are considered, showcasing the effectiveness of combined features on the multi-layer perceptron with a remarkable 91% accuracy and 0.93 F1 score.
In (Shah et al., 2020), the authors leverage NLP methods for depression detection of Reddit users based on their posts. GloVe, Word2Vec, and FastText embeddings, as well as handcrafted statistical metadata features and features from the LIWC dictionary, are used for text representation. The two-headed model, combining BiLSTM for embedding transformation and a fully connected layer for meta-features, demonstrates superior results with Word2Vec embeddings and meta-features achieving a noteworthy 0.81 F1 score. Additionally, the authors use Early Risk Detection Error and Latency metrics to take the time of classification into account.
The study (Owen et al., 2020) considers depression and anxiety detection for Twitter posts. SVM on TF-IDF vectors and GloVe embeddings, along with BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2019), are explored, showcasing the nuanced performance of BERT on a balanced dataset and better performance of SVM on an unbalanced one.
In a comprehensive survey (Babu and Kanaga, 2021), the importance of emoticons in texts in sentiment analysis for depression detection is underscored. The study covers 101 publications, emphasizing the effectiveness of combining deep learning algorithms, with CNNLSTM yielding the highest precision. Moreover, multi-class classification, categorizing sentiments into subclasses based on polarities, provides more precise results than binary and ternary classification.
In (Tejaswini et al., 2022), the authors introduce the FCL (FasttextCNNLSTM) model, comparing it with LSTM and CNN models on Glove and Word2vec embeddings. They also use datasets containing Reddit and Twitter posts. FCL outperforms other models in identifying depression in both small and large datasets.
Assessing the anxious Twitter posts caused by the COVID-19 pandemic, (Jeong et al., 2023) delves into anxiety detection using BERT trained on the Korean language, achieving a 0.67 macro F1-score and 0.91 accuracy and establishing a correlation between the anxiety index and COVID-19 waves.
The study (Ansari et al., 2023) focuses on identifying depression in social media datasets (CLpsych, Reddit, eRisk) through various text classification methods, combining sentiment lexicons with deep learning pipelines. Authors utilize sentiment lexicons with logistic regression and LSTM with attention on Glove embeddings, comparing them and combining them into ensembles. Ensemble models outperform hybrid lexicon- and DL-based models on all datasets, achieving 75% accuracy and 0.77 F1 score on the Reddit dataset.
2.2 Large Language Models
According to the systematic review (Omar and Levkovich, 2024), most studies about depression detection focus on BERT-based models, indicating the field’s early stages in adopting newer technologies like GPT-4 and Google’s Gemini. However, LLMs are demonstrating significant potential in enhancing the capabilities of depression detection systems.
The Chat-Diagnose approach (Qin et al., 2023) integrates diagnostic criteria from the Diagnostic and Statistical Manual of Mental Disorders (DSM-5) into prompts and uses the Chain of Thoughts technique to deliver explainable diagnoses via an LLMs-augmented system based on ChatGPT/GPT-3. This method demonstrates state-of-the-art results on Twitter and Weibo depression datasets by employing zero-shot and few-shot learning, using prompts with selected social media content alongside predictions from a Time2VecTransformer model to enhance diagnostic accuracy.
Another study (Hadzic et al., 2024) compares the effectiveness of fine-tuned BERT with GPT-3.5 and GPT-4 in the depression detection task. The authors use Patient Health Questionnaire-8 scores for classifying transcribed audio data from the Distress Analysis Interview Corpus, KID, and a simulated dataset. With scores separated into depressive and non-depressive groups, the zero-shot method for GPT-4 outperforms GPT-3.5 and BERT across all datasets, highlighting the potential of LLMs in depression detection.
Additionally, research (Wang et al., 2024) on eRisk 2023 and eRisk 2021 datasets explores detecting depression symptoms and classifying depression severity using LLMs. Utilizing Beck’s Depression Inventory to form queries related to depression symptoms and the Universal Sentence Encoder for text embeddings, the study creates two datasets containing top-1 and top-5 ranked texts for each query. LLMs fine-tuned with QLoRA, including Llama-2-13b-chat, SUS-Chat-34B, and Neural-chat-7b-v3, are used for classification into four levels of depression severity. The top-five dataset performs better overall, with the combined model neural-chat+SUS-Chat_top5 excelling in Average Hit Rate and Average Closeness Rate, while neural-chat_top5 and SUS-Chat_top5 achieve the highest Average DODL and Depression Category Hit Rate among single models.
The DORIS (Lan et al., 2024) system addresses the challenges of detecting depression through social media posts from the Sina Weibo Depression Dataset. The authors use GPT3.5-Turbo-1103 for annotating high-risk texts according to the DSM-5 depression scale, also LLM is used to summarize critical information from users’ historical mood records (mood courses). The final model based on XGBoost is learned on features from annotations and gte-small-zh model vector representations of post histories and mood courses. It shows an improvement in the AUPRC metric by 0.036 compared to baseline methods. This study underscores the system’s potential for early detection and intervention, providing a valuable tool for mental health assessment on social media platforms.
We are the first to examine and compare three generations of the discussed models for depression and anxiety detection tasks in Russian. Unlike other works, we used different models from each group and carefully compared the results of the models between the groups on five datasets, aiming for a general recommendation on the best models to use in practice.
3 Data
Classes in all datasets were represented in the binary format: a healthy class and a pathology class. The general description of the datasets used in our study is shown in Table 1.
| Label | DE | DSM | AL | AD | AC | ||
| Condition | Depression | Depression | Anxiety | Anxiety | Anxiety | ||
| Text format | Essay | Social media messages | Letter | Description of the picture | Short comments | ||
| Class criteria | Clinical diagnosis | BDI | HADS | HADS | SCL-90-R | ||
| # healthy samples | 447 | 135 | 109 | 101 | 222 | ||
| # pathology samples | 110 | 89 | 93 | 89 | 191 | ||
309 | 122 | 144 | 84 | 7 | ||
| Reference | (Stankevich et al., 2019) | (Ignatiev et al., 2022) | (Litvinova and Ryzhkova, 2018) | (Medvedeva et al., 2021) | |||
3.1 Depression-Essays
To compile this dataset, subjects were asked to write a short essay (up to 5,000 characters) on the topic of “Myself, others, world” (Stankevich et al., 2019). A total of 557 essays were collected, where 110 of them were written by people with clinical depression. The essays written by people with clinically validated depression were provided by the Mental Health Research Center, Moscow, Russia. The label DE (Depression-Essay) is further used to describe this dataset. The best-reported performance on this data reaches 73% F1 for the depression class in the cross-validation evaluation with the random forest model trained on n-grams and psycholinguistic features.
3.2 Depression-Social Media
The other depression dataset under consideration was created from data collected to study the problem of identifying depression in social network users (Ignatiev et al., 2022). The best-reported classification performance using textual data reaches only 65% F1 for the depression class with a logistic regression model trained on psycholinguistic features. This dataset contains text messages from the social network VKontakte and results from the Beck Depression Inventory (BDI) questionnaire (Beck et al., 1987). The healthy class included users with a scale score of 10 or less, and the pathological class between 30 and 63.
We have made some changes to the original dataset. For each subject, all messages were combined for a period of 170 days prior to the date of the questionnaire screening, and the total text was limited to 6,000 characters. Such restrictions were imposed to bring the final texts from this dataset closer to the format of the essay dataset in terms of total text length for each subject and to account for the fact that the relevance of depression screening becomes less over a period of more than six months. The label DSM (Depression-Social Media) is used hereafter to describe this dataset.
3.3 Anxiety-Letter and Picture Description
The RusNeuroPsych corpus (Litvinova and Ryzhkova, 2018) was used as the data for creating the anxiety detection model. This corpus was prepared to study the relationships between a person’s text, personal traits, mental status, and demographic characteristics. To compile it, subjects were asked to write an informal letter to a friend and give a textual description of a picture, and then complete a series of psychological questionnaires. Among the questionnaires, the Hospital Anxiety and Depression Scale (HADS) (Bjelland et al., 2002) was used. Thus, the data were divided using the results of this questionnaire, where the healthy class included subjects with a score of 7 and below, and the pathology class included all subjects with a score of 8 and above. Since the corpus contained 2 different types of text, it was divided into 2 separate datasets for the experiments. It is worth noting that the results of the HADS questionnaire also allow us to divide subjects into groups based on the presence and absence of depression, but a number of subjects with high depression scale scores was less than 20, which prevented us from conducting experiments to train classification models. The labels AL (Anxiety-Letter) for letters and AD (Anxiety-Description) for picture description are further used to describe these datasets. To the best of our knowledge, no classification experiments have been performed on this data before.
3.4 Anxiety-COVID Comments
In addition, we used a corpus of subjects’ comments on the COVID-19 pandemic situation (Medvedeva et al., 2021). Subjects were asked to complete a series of questionnaires and write a free-form commentary describing their attitudes towards the world situation around the pandemic and self-isolation. Among the questionnaires used was the SCL-90-R symptom questionnaire (Derogatis, 1983), the anxiety scale from which was used to form 2 groups: the healthy group, with an anxiety scale score below the 33rd percentile, and the pathology group, with an anxiety scale score above the 66th percentile. The label AC (Anxiety-Covid) is used hereafter to describe this dataset. To the best of our knowledge, no classification experiments have been performed on this data before.
4 Method
4.1 Linguistic Features
4.1.1 Psycholinguistic Features
The linguistic features used in our research were calculated using the tool described in (Smirnov et al., 2021). This tool allows the calculation of morphological, syntactic, and vocabulary parameters of texts, including various psycholinguistic coefficients. A total of 113 features were used. A detailed description of the features used and the features obtained with this tool on our data is given in the Huggingface repository111https://huggingface.co/datasets/anonymizedauthor/paper_data.
4.1.2 Classification Setup
As a classification baseline, we use AutoML system auto-sklearn (Feurer et al., 2015) and learn it on psycholinguistic features and on n-grams. Auto-sklearn classifier employs a Bayesian optimizer which considers 15 classification algorithms, 14 feature preprocessing methods, and 4 data preprocessing methods. Also, the classifier utilizes a meta-learning approach and automated ensemble construction to speed up optimization.
For psycholinguistic features, we consider several methods of feature selection: filter method, wrapper method (forward selection and backward elimination), and embedded method. Selected features are examined for linear correlation and ones with absolute Pearson coefficient values more than 0.95 are deleted.
We also train models on TF-IDF vectors on unigrams and unigrams with bigrams. In addition to full vectors, we consider different subsets of features for vectors on unigrams: from 20% to 100% of the features are selected with 20% increment. For each subset, correlated features are deleted in the same way as with psycholinguistic features.
On each set of features we launch the classifier 6 times with different seeds and then calculate the mean and standard deviation values of the metrics based on the results of the launches.
4.2 Encoder Models
We employ classification with encoder models to compare methods, based on psycholinguistic features, with more modern solutions. As we target classification on Russian corpora, we consider models pretrained on multilingual or Russian datasets. As a simple baseline, we used a base version of multilingual BERT, a model with 110 million parameters, first introduced in (Devlin et al., 2019). Another considered baseline is RuBERT (Kuratov and Arkhipov, 2019), a pretrained Russian version of BERT-base with the same amount of parameters. We also finetuned several more recent models, such as RuBioRoBERTa (Yalunin et al., 2022) which is a RoBERTa, pretrained on Russian language biomedical texts, and RuRoberta-large - a bigger version of RoBERTa, also pretrained on Russian language datasets. For all of the four models, we optimized hyperparameters using Bayesian optimization from HuggingFace (Wolf et al., 2019) framework.
The used hyperparameter grid and optimal parameters, alongside with used checkpoint of models, are presented in Appendix A.
4.3 Large Language Models
We conducted experiments with LLMs in various settings. First of all, we evaluated models by 0-shot and 5-shot prompting, considering only normalized probabilities for tokens “0” or “1” in the first generated token, as it was done in MMLU (Hendrycks et al., 2021). We will refer to these settings as “0-shot MMLU” and “5-shot MMLU” correspondingly. Secondly, we employed 0-shot and 5-shot prompting with bigger generation lengths and matched the generated answer to one of the possible classes. Finally, we conducted the finetuning of the models using LoRA (Hu et al., 2022).
We selected a set of relatively small (less than 9B parameters) self-hosted open-source models, either tuned for the Russian language or showing good multilingual capabilities. To the first group belongs SaigaLlama3 8B, a version of Llama 3 8B Instruct (Dubey et al., 2024) tuned on several Russian datasets, as well as models from Vikhr family (Nikolich et al., 2024). We used Vikhr 7B Instruct 0.4, Vikhr 7B Instruct 5.4, and Vikhr Gemma 2B Instruct. The former two are based on Mistral 7B (Jiang et al., 2023) with vocabulary adaptation for the Russian language followed by additional pre-training and instruction tuning. The latter one is based on Gemma2 2B Instruct (Team et al., 2024), additionally trained on Russian data. We also used multilingual models, such as Gemma2 2B Instruct and Gemma2 9B Instruct, as well as Qwen2 7B Instruct (Yang et al., 2024). As we are conducting experiments on sensitive data and with private datasets, we did not consider remotely-hosted models (such as GPT4, Claude) due to the possibility of data leaking.
The full information about used prompts, training hyperparameters, and versions of used models, are presented in Appendix A.
5 Results
5.1 Classification Results on Depression and Anxiety Datasets
The data were divided into training and test samples in an 80% by 20% ratio, with stratification by target variable reduced to binary form. All classification reports in this study show results on the test data. The classification report for the best models from each group is presented in Table 2.
| Corpus | Mode | Model |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| DE | SFT | Linguistic features | 94.00.8 | 95.21.0 | 94.60.8 | 79.34.0 | 75.03.5 | 77.03.3 | 85.82.1 | ||||||||
| SFT | TF-IDF | 91.62.1 | 94.42.2 | 93.01.2 | 74.87.0 | 64.410.3 | 68.56.6 | 80.73.8 | |||||||||
| 5-shot | Vikhr 7B IT 5.4 | 87.060.0 | 82.220.0 | 84.570.0 | 40.740.0 | 50.00.0 | 44.90.0 | 64.730.0 | |||||||||
| 5-shot MMLU | Gemma2 9B IT | 92.960.0 | 73.330.0 | 81.990.0 | 41.460.0 | 77.270.0 | 53.970.0 | 67.980.0 | |||||||||
| LoRA | VikhrGemma 2B IT | 94.740.88 | 96.481.62 | 95.590.66 | 84.965.72 | 78.034.08 | 81.132.42 | 88.361.52 | |||||||||
| SFT | RuBERT | 94.451.16 | 91.111.7 | 92.741.04 | 68.414.04 | 78.034.85 | 72.83.43 | 82.772.21 | |||||||||
| 0-shot | Gemma2 9B IT | 93.330.0 | 62.220.0 | 74.670.0 | 34.620.0 | 81.820.0 | 48.650.0 | 61.660.0 | |||||||||
| DSM | SFT | Linguistic features | 62.81.8 | 59.37.4 | 60.74.0 | 43.72.7 | 47.27.7 | 45.13.8 | 52.92.0 | ||||||||
| SFT | TF-IDF | 62.22.6 | 51.213.9 | 55.38.5 | 42.93.6 | 53.711.4 | 46.94.7 | 51.13.5 | |||||||||
| 5-shot | SaigaLlama3 8B | 100.00.0 | 3.70.0 | 7.140.0 | 40.910.0 | 100.00.0 | 58.060.0 | 32.60.0 | |||||||||
| 5-shot | Vikhr 7B IT 0.4 | 70.970.0 | 81.480.0 | 75.860.0 | 64.290.0 | 50.00.0 | 56.250.0 | 66.060.0 | |||||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00.0 | 3.70.0 | 7.140.0 | 40.910.0 | 100.00.0 | 58.060.0 | 32.60.0 | |||||||||
| 5-shot MMLU | Vikhr 7B IT 5.4 | 68.750.0 | 81.480.0 | 74.580.0 | 61.540.0 | 44.440.0 | 51.610.0 | 63.090.0 | |||||||||
| LoRA | Qwen2 7B IT | 68.876.39 | 72.228.21 | 70.46.75 | 55.3311.08 | 50.9310.84 | 52.8110.26 | 61.618.32 | |||||||||
| SFT | RuBioRoBERTa | 62.12.98 | 62.966.05 | 62.464.3 | 43.664.43 | 42.594.14 | 43.013.7 | 52.743.67 | |||||||||
| AL | SFT | Linguistic features | 47.421.4 | 36.417.8 | 40.919.2 | 48.52.7 | 68.414.6 | 56.13.7 | 48.58.2 | ||||||||
| SFT | TF-IDF | 61.55.0 | 46.23.1 | 52.62.3 | 51.22.5 | 65.87.9 | 57.54.5 | 55.002.9 | |||||||||
| 5-shot | Gemma2 9B IT | 75.00.0 | 27.270.0 | 40.00.0 | 51.520.0 | 89.470.0 | 65.380.0 | 52.690.0 | |||||||||
| LoRA | Gemma2 2B IT | 60.184.51 | 60.6113.55 | 59.668.17 | 54.877.3 | 53.5110.71 | 53.456.06 | 56.565.4 | |||||||||
| SFT | RuRoBERTa | 56.825.65 | 75.07.76 | 64.495.59 | 52.5410.76 | 33.3312.77 | 40.1712.48 | 52.338.31 | |||||||||
| 0-shot | Gemma2 9B IT | 66.670.0 | 63.640.0 | 65.120.0 | 60.00.0 | 63.160.0 | 61.540.0 | 63.330.0 | |||||||||
| 0-shot MMLU | Gemma2 9B IT | 64.710.0 | 50.00.0 | 56.410.0 | 54.170.0 | 68.420.0 | 60.470.0 | 58.440.0 | |||||||||
| 0-shot MMLU | Gemma2 2B IT | 0.00.0 | 0.00.0 | 0.00.0 | 46.340.0 | 100.00.0 | 63.330.0 | 31.670.0 | |||||||||
| AD | SFT | Linguistic features | 50.83.3 | 43.312.5 | 45.38.1 | 44.72.6 | 51.915.9 | 47.37.4 | 46.31.8 | ||||||||
| SFT | TF-IDF | 54.915.1 | 37.54.8 | 43.22.5 | 44.67.8 | 59.321.0 | 50.412.7 | 46.87.1 | |||||||||
| LoRA | Vikhr 7B IT 0.4 | 59.814.42 | 51.6712.13 | 54.667.59 | 53.474.23 | 61.1111.56 | 56.436.19 | 55.554.55 | |||||||||
| SFT | RuRoBERTa | 57.632.52 | 56.674.71 | 57.073.19 | 52.82.82 | 53.74.14 | 53.172.84 | 55.122.58 | |||||||||
| 0-shot | Vikhr 7B IT 5.4 | 83.330.0 | 25.00.0 | 38.460.0 | 53.120.0 | 94.440.0 | 68.00.0 | 53.230.0 | |||||||||
| 0-shot | Qwen2 7B IT | 69.570.0 | 80.00.0 | 74.420.0 | 73.330.0 | 61.110.0 | 66.670.0 | 70.540.0 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 73.680.0 | 70.00.0 | 71.790.0 | 68.420.0 | 72.220.0 | 70.270.0 | 71.030.0 | |||||||||
| AC | SFT | Linguistic features | 64.816.0 | 45.621.0 | 47.019.6 | 48.73.6 | 60.520.1 | 52.67.3 | 49.87.8 | ||||||||
| SFT | TF-IDF | 56.32.1 | 63.34.2 | 59.52.5 | 48.93.3 | 41.75.6 | 44.94.3 | 52.22.7 | |||||||||
| 5-shot | Gemma2 9B IT | 100.00.0 | 15.560.0 | 26.920.0 | 50.00.0 | 100.00.0 | 66.670.0 | 46.790.0 | |||||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00.0 | 13.330.0 | 23.530.0 | 49.350.0 | 100.00.0 | 66.090.0 | 44.810.0 | |||||||||
| LoRA | Vikhr 7B IT 5.4 | 61.872.07 | 62.968.18 | 62.134.3 | 55.53.03 | 53.957.25 | 54.333.58 | 58.232.38 | |||||||||
| SFT | BERT | 57.015.15 | 62.9616.81 | 58.446.53 | 46.078.77 | 41.6722.14 | 41.2817.15 | 49.866.84 | |||||||||
| 0-shot | SaigaLlama3 8B | 67.740.0 | 46.670.0 | 55.260.0 | 53.850.0 | 73.680.0 | 62.220.0 | 58.740.0 | |||||||||
| 0-shot MMLU | SaigaLlama3 8B | 70.370.0 | 42.220.0 | 52.780.0 | 53.570.0 | 78.950.0 | 63.830.0 | 58.30.0 |
5.1.1 DE Dataset
The best scores of the F1 for the pathology class and F1-macro in this experiment are achieved on the essay dataset (DE) by the finetuned LLM model with 88.4% F1-macro and 81.1% F1 score for the pathology class. The model trained on the linguistic features shows similar results with 85.8% F1-macro and 77.0% F1-pathology.
In general, the three best results on the DE dataset were achieved by finetuned models, which can be linked to the bigger dataset size and to the longer texts in the dataset, which, in turn, are crucial for supervised finetuning. For all other datasets methods without finetuning significantly outperform SFT and LoRA.
| Corpus | Mode | Model |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| D-all | 5-shot MMLU | Gemma2 9B IT | 89.330.0 | 97.550.0 | 93.260.0 | 69.230.0 | 32.140.0 | 43.90.0 | 68.580.0 | ||||||||
| LoRA | Gemma2 2B IT | 92.330.94 | 97.141.04 | 94.670.72 | 76.446.36 | 52.986.33 | 62.335.6 | 78.53.13 | |||||||||
| SFT | RuBERT | 92.260.52 | 96.221.56 | 94.190.68 | 72.1210.1 | 52.983.81 | 60.62.87 | 77.391.72 | |||||||||
| 0-shot | Gemma2 9B IT | 93.280.0 | 76.690.0 | 84.180.0 | 33.330.0 | 67.860.0 | 44.710.0 | 64.440.0 | |||||||||
| SFT | Linguistic features | 88.10.6 | 88.13.4 | 88.01.6 | 51.65.8 | 50.74.1 | 50.82.6 | 69.41.9 | |||||||||
| SFT | TF-IDF | 89.51.0 | 84.21.0 | 86.80.6 | 47.51.8 | 59.14.2 | 52.62.5 | 69.71.5 | |||||||||
| A-all | 5-shot | Gemma2 2B IT | 72.00.0 | 20.690.0 | 32.140.0 | 49.640.0 | 90.670.0 | 64.150.0 | 48.150.0 | ||||||||
| LoRA | SaigaLlama3 8B | 58.561.32 | 56.324.83 | 57.332.79 | 51.592.03 | 53.783.74 | 52.561.83 | 54.951.57 | |||||||||
| SFT | RuRoBERTa | 61.533.69 | 60.1513.8 | 59.684.47 | 54.390.97 | 54.4418.08 | 52.2711.87 | 55.973.88 | |||||||||
| SFT | BERT | 59.441.74 | 46.933.66 | 52.332.1 | 50.430.98 | 62.674.81 | 55.822.21 | 54.081.08 | |||||||||
| 0-shot | Qwen2 7B IT | 60.820.0 | 67.820.0 | 64.130.0 | 56.920.0 | 49.330.0 | 52.860.0 | 58.490.0 | |||||||||
| 0-shot MMLU | SaigaLlama3 8B | 69.770.0 | 34.480.0 | 46.150.0 | 52.10.0 | 82.670.0 | 63.920.0 | 55.040.0 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 62.790.0 | 62.070.0 | 62.430.0 | 56.580.0 | 57.330.0 | 56.950.0 | 59.690.0 | |||||||||
| SFT | Linguistic features | 53.42.9 | 39.719.3 | 42.017.9 | 46.12.4 | 60.220.4 | 51.07.8 | 46.56.2 | |||||||||
| SFT | TF-IDF | 51.80.7 | 40.015.6 | 43.310.8 | 44.32.1 | 56.417.8 | 48.68.5 | 46.02.4 |
| Transfer | Mode | Model |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| DE to DSM test | LoRA | SaigaLlama3 8B | 58.493.68 | 83.9513.97 | 68.737.39 | 40.3413.85 | 12.044.99 | 17.215.85 | 42.974.2 | ||||||||
| LoRA | Vikhr 7B IT 0.4 | 60.335.19 | 61.7320.91 | 58.8110.42 | 37.585.47 | 37.0424.57 | 33.9815.06 | 46.394.16 | |||||||||
| LoRA | Gemma2 2B IT | 64.143.85 | 71.621.45 | 65.8910.13 | 51.078.34 | 38.8922.91 | 38.6716.04 | 52.286.44 | |||||||||
| LoRA | Gemma2 9B IT | 64.465.24 | 66.6720.73 | 64.1111.57 | 51.248.98 | 45.3718.54 | 44.3611.95 | 54.237.12 | |||||||||
| LoRA | Vikhr 7B IT 5.4 | 58.245.42 | 67.2821.75 | 60.2411.51 | 31.337.8 | 25.9324.36 | 25.0915.4 | 42.664.75 | |||||||||
| LoRA | Qwen2 7B IT | 61.143.36 | 61.1118.61 | 59.7511.42 | 43.628.81 | 42.5915.93 | 41.258.0 | 50.55.32 | |||||||||
| LoRA | VikhrGemma 2B IT | 60.842.46 | 71.616.93 | 64.98.82 | 42.547.31 | 31.4815.27 | 33.9712.07 | 49.445.25 | |||||||||
| SFT | RuRoBERTa | 60.967.01 | 41.9826.59 | 46.0714.36 | 33.7515.28 | 59.2627.34 | 42.9519.5 | 44.515.07 | |||||||||
| SFT | RuBioRoBERTa | 61.652.51 | 50.023.88 | 52.5412.11 | 34.5415.55 | 52.7824.59 | 41.5718.71 | 47.055.29 | |||||||||
| SFT | RuBERT | 68.010.29 | 27.789.01 | 38.6810.16 | 42.662.8 | 80.569.49 | 55.644.19 | 47.165.8 | |||||||||
| SFT | BERT | 76.855.46 | 29.638.28 | 42.219.33 | 45.342.29 | 87.044.14 | 59.542.03 | 50.885.43 | |||||||||
| SFT | Linguistic features | 60.02.5 | 57.47.0 | 58.54.5 | 40.03.5 | 42.66.9 | 41.14.5 | 49.83.1 | |||||||||
| SFT | TF-IDF | 59.93.4 | 30.99.2 | 40.08.5 | 40.21.5 | 69.48.3 | 50.72.7 | 45.43.8 |
5.1.2 DSM Dataset
The best result achieved by Vikhr 7B IT 0.4 model, evaluated in 5-shot, with 66.1% F1-macro. The traditional machine learning methods, as well as encoder models, performed poorly with nearly 53% F1-macro on linguistic features.
Comparing the results between the DE and DSM datasets, a significant difference in classification quality can be observed. This may be due to several factors. First of all, we can note the sample size, which was significantly smaller in the DSM dataset. Anyway, almost all machine learning models, regardless of the selected features, can show low accuracy on a small amount of data. On the other hand, if we refer to the study (Ignatiev et al., 2022), where less stringent restrictions on text volume and temporal proximity to questionnaire screening dates were applied to the same raw data, the results were still not very high: about 60% F1-macro for psycholinguistic features and TF-IDF features.
Secondly, the problem may lie in the format of the texts from the social network. The texts in DSM are concatenated together with a collection of posts from users’ personal pages, and they mostly lack coherent logic and cohesion in the resulting text. Even with the 6,000 character limit, the standard deviation in word count in this data is on the order of 300 words, while in DE the similar value is around 100. Thus, DSM is a much noisier dataset, which can strongly affect the quality of classification with psycholinguistic markers, where the values of some markers can be affected by the volume of text analyzed, even considering their normalization with respect to the text volume.
Finally, the way in which the target class of pathology is defined can be of great importance. Although a widely used and validated psychological questionnaire was used for the DSM dataset, it is not possible to directly compare its results with a clinically confirmed diagnosis. The same findings can be outlined in work that criticizes the way in which social media users are labeled for mental illness by indirectly affiliating or self-identifying mental ill-health (Ernala et al., 2019). In favor of the significance of this factor is also the fact that TF-IDF-based features performed significantly better on DE data than on DSM, although, unlike psycholinguistic markers, they do not have an initial specialization for detecting signs of mental ill-health.
To take into account all of the aforementioned considerations, we applied LLMs to this task and showed, that the bigger model is able to partially overcome these issues. While RuBERT and the model trained on linguistic features both show comparatively low results, the evaluated in 5-shot format model performs significantly better. The same outcomes are observed in the experiments with anxiety datasets.
5.1.3 AL, AD, and AC Datasets
Turning to the anxiety detection task, the difference between the AutoML-based and encoder-based models is only noticeable on the picture description (AD) dataset, where the accuracy of the RuRoBERTa model was 55% F1-macro. Overall, it can be said that none of the non-LLM models performed well in this task. However, LLMs performed significantly better on these tasks. Models with LoRA in general perform better than encoder-based and AutoML-based models, but the best performance is achieved by 0-shot and 5-shot prompting. Again, this can be linked to the complexity of the domain and the small amount of training data, which makes it hard to finetune a model.
In general, the LLMs outperform other models on all datasets, but on the DE and DSM, the gap between various types of models is smaller, leaving the usage of non-LLM models reasonable for some specific cases.
5.2 Classification on D-all and A-all Datasets
We combined the depression and anxiety datasets into pooled datasets D-all (DE and DSM) and A-all (AL, AD, and AC). Classification performance of the best models on D-all and A-all combined data presented in Table 3.
The LLMs performed better than other models on both D-all and A-all. It is also noticeable that methods with model tuning work better on D-all, while prompting shows better results on A-all. It can be explained with significant genre and length variations on A-all datasets, so the general-purpose models perform better than finetuned ones due to the complexity of the data.
5.3 Classification of DSM Data with Best Models from DE
The classification results of models that demonstrated the best performance on the DE dataset applied to DSM test data are shown in Table 4.
This experiment shows that models trained on essays from depressed subjects cannot be used to detect depression from social media texts from users who have taken a depression questionnaire and received a high score. Similar findings were shown in (Ernala et al., 2019), but with a reverse logic of the experiments.
Such results may also be just due to the fact that essay texts and collections of social network posts are very different in genre. A sample of social network users with clinically diagnosed depression would be needed to clarify this issue.
6 Conclusion
In this paper, we have investigated the effectiveness of traditional machine learning methods and LLMs on the task of detecting depression and anxiety. The results obtained in our work establish the new state-of-the-art on five Russian-language datasets.
Our investigation shows that psycholinguistic features can produce results at the level of encoder-based models when texts from individuals with clinically diagnosed depression are used for training. BERT models, in turn, perform better on noisy text data, where examples from the training sample may vary widely in text length or genre. LLM-based models performed best on all five different mental health datasets. Even without fine-tuning, LLMs usually demonstrate relatively high performance. In response to RQ1, the experimental results indicate that LLM-based models have high potential for detecting mental disorders from texts.
In response to RQ2, the findings reveal that models trained on essays from depressed individuals are not effective for detecting depression in social media texts from users who have completed a depression questionnaire and scored high.
7 Limitations
The main limitation of this paper is that it is impossible to share all the raw texts from used datasets, as they are all distributed under different terms.
The amount of sample data used for predicting anxiety is small, which does not allow us to adequately judge the possibility of predicting anxiety in Russian text using machine learning methods. Although the results on all the anxiety datasets used show poor accuracy, perhaps a very different scale of data is needed for this task. The paper does not discuss the differences between the results inside one group of the models in detail as this was not the aim of the paper.
We conducted the experiments only for the Russian due to the poor availability of related datasets for other languages. However, the used methods in general are language agnostic so the results could be extended to other languages. The overall result about LLMs as the best-performing models matches similar studies for English on closed data.
8 Ethical Considerations
The problem discussed in this paper is the sensitive issue of mental health. To avoid any possible harm, we did not fully open-source used datasets. All shared data does not contain any information that names or uniquely identifies individuals, nor does it contain offensive content. The examined models for the detection of mental disorders do not aim to replace a professional physician, on the contrary, these models are intended to support a human expert.
Appendix A Hyperparameters
The used net and optimal parameters, alongside with used checkpoint of encoder models, are presented in Table 6. The hyperparameters for LLMs, tuned with LoRA are presented in Table 7. The LoRA was applied to the following target layers: q_proj, up_proj, o_proj, k_proj, down_proj, gate_proj, v_proj. For LLMs 0-shot and 5-shot evaluation, we used the following generation parameters: do_sample=False, temperature=1.0, top_p=1.0, top_k=50, repetition_penalty=1.0. For MMLU-style evaluation generation length was limited to one token while for regular evaluation it was set to 64 tokens. Prompt examples are presented in Table 5.
| Task | Mode | Prompt example (system & user) |
| Depression | 0-shot |
System: You play the role of a psychologist’s assistant who helps diagnose the presence or absence of a depressive disorder. You will be given a text written by a person. Determine the author’s depression level from the text, where 0 is no depression, 1 is depression, and then write why you chose this answer.
User: Text: {input_text} Answer (0 or 1): |
| Depression | 5-shot |
System: You play the role of a psychologist’s assistant who helps diagnose the presence or absence of a depressive disorder. You will be given a text written by a person. Determine the author’s depression level from the text, where 0 is no depression, 1 is depression, and then write why you chose this answer.
User: Text: {example_1} Answer (0 or 1): 0 Text: {example_2} Answer (0 or 1): 0 Text: {example_3} Answer (0 or 1): 0 Text: {example_4} Answer (0 or 1): 1 Text: {example_5} Answer (0 or 1): 1 Text: {input_text} Answer (0 or 1): |
| Anxiety | 0-shot |
System: You play the role of a psychologist’s assistant who helps diagnose the presence or absence of an anxiety disorder. You will be given a text written by a person. Determine the level of anxiety of the author of the text, where 0 is no anxiety, 1 is anxiety, and then write why you chose this answer.
User: Text: {input_text} Answer (0 or 1): |
| Anxiety | 5-shot |
System: You play the role of a psychologist’s assistant who helps diagnose the presence or absence of an anxiety disorder. You will be given a text written by a person. Determine the level of anxiety of the author of the text, where 0 is no anxiety, 1 is anxiety, and then write why you chose this answer.
User: Text: {example_1} Answer (0 or 1): 0 Text: {example_2} Answer (0 or 1): 0 Text: {example_3} Answer (0 or 1): 0 Text: {example_4} Answer (0 or 1): 1 Text: {example_5} Answer (0 or 1): 1 Text: {input_text} Answer (0 or 1): |
| Model name | Corpus |
| RuBERT | A-all | 15 | 5e-05 | 16 | 0.1 | ||||||||
| D-all | 15 | 5e-05 | 16 | 0.1 | |||||||||
| DE | 10 | 7e-05 | 32 | 0.1 | |||||||||
| AD | 2 | 6e-06 | 4 | 0.1 | |||||||||
| AL | 2 | 9e-06 | 4 | 0.1 | |||||||||
| AC | 2 | 6e-06 | 4 | 0.1 | |||||||||
| DSM | 13 | 1e-04 | 4 | 0.0 | |||||||||
| RuRoBERTa | A-all | 13 | 9e-06 | 16 | 0.01 | ||||||||
| D-all | 11 | 6e-06 | 8 | 0.0 | |||||||||
| DE | 9 | 9e-06 | 4 | 0.0 | |||||||||
| AD | 7 | 2e-05 | 4 | 0.0 | |||||||||
| AL | 11 | 6e-06 | 8 | 0.0 | |||||||||
| AC | 6 | 1e-04 | 32 | 0.01 | |||||||||
| DSM | 13 | 5e-06 | 16 | 0.1 | |||||||||
| RuBioRoBERTa | A-all | 15 | 3e-05 | 16 | 0.01 | ||||||||
| D-all | 11 | 7e-06 | 32 | 0.0 | |||||||||
| DE | 8 | 1e-05 | 4 | 0.0 | |||||||||
| AD | 2 | 7e-06 | 8 | 0.1 | |||||||||
| AL | 9 | 2e-05 | 16 | 0.01 | |||||||||
| AC | 8 | 1e-05 | 4 | 0.0 | |||||||||
| DSM | 15 | 9e-06 | 16 | 0.01 | |||||||||
| BERT | A-all | 14 | 7e-06 | 16 | 0.01 | ||||||||
| D-all | 15 | 5e-05 | 16 | 0.1 | |||||||||
| DE | 15 | 5e-05 | 16 | 0.1 | |||||||||
| AD | 2 | 6e-06 | 4 | 0.1 | |||||||||
| AL | 2 | 6e-06 | 4 | 0.1 | |||||||||
| AC | 4 | 1e-04 | 32 | 0.01 | |||||||||
| DSM | 5 | 1e-04 | 32 | 0.01 |
| Model name | Corpus |
LoRA |
| SaigaLlama3 8B | DE | 13 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 8 | ||||||||||||
| DSM | 3 | 1e-05 | 16 | 1e-02 | 32 | 5e-02 | 16 | |||||||||||||
| AL | 15 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 16 | |||||||||||||
| AD | 15 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 16 | |||||||||||||
| AC | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| K-all | 14 | 3e-05 | 16 | 0e+00 | 16 | 5e-02 | 16 | |||||||||||||
| D-all | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| Vikhr 7B IT 0.4 | DE | 2 | 5e-05 | 4 | 1e-01 | 32 | 5e-02 | 8 | ||||||||||||
| DSM | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| AL | 4 | 9e-06 | 8 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| AD | 11 | 1e-04 | 4 | 0e+00 | 32 | 1e-01 | 16 | |||||||||||||
| AC | 3 | 1e-04 | 16 | 0e+00 | 16 | 5e-02 | 16 | |||||||||||||
| K-all | 11 | 2e-05 | 16 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| D-all | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| Vikhr 7B IT 5.4 | DE | 15 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 16 | ||||||||||||
| DSM | 4 | 9e-06 | 8 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| AL | 7 | 6e-06 | 8 | 0e+00 | 16 | 1e-01 | 8 | |||||||||||||
| AD | 8 | 5e-06 | 4 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| AC | 15 | 1e-04 | 4 | 0e+00 | 16 | 1e-01 | 8 | |||||||||||||
| K-all | 12 | 3e-05 | 16 | 0e+00 | 16 | 1e-01 | 8 | |||||||||||||
| D-all | 8 | 5e-06 | 4 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| VikhrGemma 2B IT | DE | 15 | 1e-04 | 4 | 0e+00 | 16 | 1e-01 | 8 | ||||||||||||
| DSM | 2 | 6e-06 | 8 | 0e+00 | 16 | 1e-01 | 16 | |||||||||||||
| AL | 5 | 5e-05 | 4 | 1e-02 | 16 | 5e-02 | 16 | |||||||||||||
| AD | 8 | 2e-05 | 4 | 0e+00 | 32 | 5e-02 | 16 | |||||||||||||
| AC | 2 | 5e-05 | 4 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| K-all | 11 | 2e-05 | 16 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| D-all | 15 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 16 | |||||||||||||
| Gemma2 2B IT | DE | 15 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 16 | ||||||||||||
| DSM | 8 | 7e-05 | 16 | 0e+00 | 32 | 5e-02 | 8 | |||||||||||||
| AL | 13 | 1e-04 | 4 | 0e+00 | 16 | 1e-01 | 8 | |||||||||||||
| AD | 12 | 7e-05 | 16 | 0e+00 | 32 | 5e-02 | 8 | |||||||||||||
| AC | 14 | 1e-04 | 4 | 0e+00 | 32 | 1e-01 | 16 | |||||||||||||
| K-all | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| D-all | 15 | 1e-04 | 4 | 1e-01 | 16 | 1e-01 | 16 | |||||||||||||
| Gemma2 9B IT | DE | 9 | 1e-04 | 8 | 0e+00 | 16 | 5e-02 | 8 | ||||||||||||
| DSM | 13 | 5e-06 | 8 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| AL | 6 | 1e-04 | 8 | 0e+00 | 16 | 1e-01 | 16 | |||||||||||||
| AD | 11 | 2e-05 | 16 | 1e-02 | 16 | 1e-01 | 16 | |||||||||||||
| AC | 15 | 1e-04 | 4 | 0e+00 | 16 | 1e-01 | 8 | |||||||||||||
| K-all | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| D-all | 15 | 1e-04 | 4 | 0e+00 | 16 | 1e-01 | 8 | |||||||||||||
| Qwen2 7B IT | DE | 5 | 5e-05 | 4 | 1e-02 | 16 | 5e-02 | 16 | ||||||||||||
| DSM | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| AL | 4 | 5e-05 | 8 | 0e+00 | 16 | 5e-02 | 8 | |||||||||||||
| AD | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| AC | 9 | 3e-05 | 4 | 0e+00 | 32 | 1e-01 | 8 | |||||||||||||
| K-all | 11 | 7e-05 | 16 | 1e-01 | 32 | 5e-02 | 8 | |||||||||||||
| D-all | 13 | 1e-04 | 8 | 1e-01 | 32 | 5e-02 | 8 |
Appendix B Full Results for Models
Table 8 shows results for TF-IDF, models on linguistic features, and encoder-based transformers. Tables 9, 11 and 10 contain results for LLMs with PEFT, 0-shot and 5-shot evaluation, and 0-shot and 5-shot MMLU-style evaluation correspondingly.
| Corpus | Model |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| DE | RuRoBERTa | 90.444.62 | 95.932.28 | 92.981.92 | 64.729.29 | 56.8225.81 | 60.3727.21 | 76.6714.52 | ||||||||
| RuBioRoBERTa | 91.925.21 | 95.193.05 | 93.362.13 | 64.2529.33 | 63.6428.63 | 63.7328.62 | 78.5415.3 | |||||||||
| RuBERT | 94.451.16 | 91.111.7 | 92.741.04 | 68.414.04 | 78.034.85 | 72.83.43 | 82.772.21 | |||||||||
| BERT | 92.970.8 | 92.961.23 | 92.960.85 | 71.343.96 | 71.213.39 | 71.233.14 | 82.091.98 | |||||||||
| Linguistic features | 94.00.8 | 95.21.0 | 94.60.8 | 79.34.0 | 75.03.5 | 77.03.3 | 85.82.1 | |||||||||
| TF-IDF | 91.62.1 | 94.42.2 | 93.01.2 | 74.87.0 | 64.410.3 | 68.56.6 | 80.73.8 | |||||||||
| DSM | RuRoBERTa | 61.314.38 | 73.4612.73 | 66.587.66 | 46.7610.29 | 31.486.93 | 36.686.2 | 51.635.8 | ||||||||
| RuBioRoBERTa | 62.12.98 | 62.966.05 | 62.464.3 | 43.664.43 | 42.594.14 | 43.013.7 | 52.743.67 | |||||||||
| RuBERT | 60.781.75 | 96.916.9 | 74.521.07 | 9.0920.33 | 5.5612.42 | 6.915.42 | 40.717.18 | |||||||||
| BERT | 61.091.9 | 69.1414.61 | 64.135.24 | 34.5915.71 | 33.3316.97 | 33.7116.01 | 48.925.62 | |||||||||
| Linguistic features | 62.81.8 | 59.37.4 | 60.74.0 | 43.72.7 | 47.27.7 | 45.13.8 | 52.92.0 | |||||||||
| TF-IDF | 62.22.6 | 51.213.9 | 55.38.5 | 42.93.6 | 53.711.4 | 46.94.7 | 51.13.5 | |||||||||
| AL | RuRoBERTa | 56.825.65 | 75.07.76 | 64.495.59 | 52.5410.76 | 33.3312.77 | 40.1712.48 | 52.338.31 | ||||||||
| RuBioRoBERTa | 53.621.09 | 78.7915.67 | 63.225.04 | 30.6621.79 | 21.0515.79 | 24.7618.06 | 43.996.71 | |||||||||
| RuBERT | 53.633.2 | 77.2719.99 | 62.087.52 | 45.8329.56 | 21.9320.48 | 24.6517.57 | 43.365.79 | |||||||||
| BERT | 52.711.02 | 93.948.16 | 67.442.96 | 6.2513.98 | 2.635.88 | 3.78.28 | 35.572.99 | |||||||||
| Linguistic features | 47.421.4 | 36.417.8 | 40.919.2 | 48.52.7 | 68.414.6 | 56.13.7 | 48.58.2 | |||||||||
| TF-IDF | 61.55.0 | 46.23.1 | 52.62.3 | 51.22.5 | 65.87.9 | 57.54.5 | 55.02.9 | |||||||||
| AD | RuRoBERTa | 57.632.52 | 56.674.71 | 57.073.19 | 52.82.82 | 53.74.14 | 53.172.84 | 55.122.58 | ||||||||
| RuBioRoBERTa | 45.065.52 | 43.3319.72 | 41.4913.8 | 38.298.58 | 40.7423.28 | 37.7212.46 | 39.65.38 | |||||||||
| RuBERT | 48.637.83 | 80.029.58 | 59.3415.36 | 12.5918.14 | 12.9618.33 | 12.718.04 | 36.024.94 | |||||||||
| BERT | 52.031.0 | 93.337.99 | 66.742.88 | 20.5621.12 | 4.634.99 | 7.347.71 | 37.042.63 | |||||||||
| Linguistic features | 50.83.3 | 43.312.5 | 45.38.1 | 44.72.6 | 51.915.9 | 47.37.4 | 46.31.8 | |||||||||
| TF-IDF | 54.915.1 | 37.54.8 | 43.22.5 | 44.67.8 | 59.321.0 | 50.412.7 | 46.87.1 | |||||||||
| AC | RuRoBERTa | 45.720.48 | 77.0435.48 | 57.1925.66 | 24.5724.96 | 24.5635.49 | 21.2223.82 | 39.27.55 | ||||||||
| RuBioRoBERTa | 52.921.41 | 76.314.15 | 62.095.54 | 34.8815.72 | 20.1811.64 | 24.8412.88 | 43.474.21 | |||||||||
| RuBERT | 54.291.57 | 87.0419.35 | 66.08.24 | 43.5922.45 | 14.0417.1 | 17.1515.7 | 41.584.88 | |||||||||
| BERT | 57.015.15 | 62.9616.81 | 58.446.53 | 46.078.77 | 41.6722.14 | 41.2817.15 | 49.866.84 | |||||||||
| Linguistic features | 64.816.0 | 45.621.0 | 47.019.6 | 48.73.6 | 60.520.1 | 52.67.3 | 49.87.8 | |||||||||
| TF-IDF | 56.32.1 | 63.34.2 | 59.52.5 | 48.93.3 | 41.75.6 | 44.94.3 | 52.22.7 |
| Corpus | Model |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| DE | Gemma2 2B IT | 97.221.4 | 92.961.39 | 95.020.48 | 69.76.78 | 86.634.79 | 76.843.12 | 85.931.75 | ||||||||
| Gemma2 9B IT | 96.481.0 | 91.90.9 | 94.130.58 | 65.154.29 | 82.113.76 | 72.532.91 | 83.331.72 | |||||||||
| Qwen2 7B IT | 95.192.1 | 90.180.34 | 92.60.96 | 57.582.14 | 75.467.33 | 65.052.19 | 78.831.56 | |||||||||
| SaigaLlama3 8B | 98.71.19 | 91.590.79 | 95.010.41 | 62.884.08 | 92.885.97 | 74.742.12 | 84.871.22 | |||||||||
| Vikhr 7B IT 0.4 | 94.072.77 | 90.261.15 | 92.11.11 | 58.336.11 | 72.8212.43 | 63.923.46 | 78.012.17 | |||||||||
| Vikhr 7B IT 5.4 | 95.742.52 | 90.82.36 | 93.151.15 | 59.8512.14 | 79.079.93 | 66.918.35 | 80.034.65 | |||||||||
| VikhrGemma 2B IT | 96.481.62 | 94.740.88 | 95.590.66 | 78.034.08 | 84.965.72 | 81.132.42 | 88.361.52 | |||||||||
| DSM | Gemma2 2B IT | 58.649.18 | 63.294.98 | 60.626.33 | 49.079.31 | 44.425.41 | 46.246.49 | 53.435.33 | ||||||||
| Gemma2 9B IT | 56.796.65 | 64.23.12 | 60.164.97 | 52.784.24 | 45.124.14 | 48.543.57 | 54.353.9 | |||||||||
| Qwen2 7B IT | 72.228.21 | 68.876.39 | 70.46.75 | 50.9310.84 | 55.3311.08 | 52.8110.26 | 61.618.32 | |||||||||
| SaigaLlama3 8B | 56.1715.48 | 57.467.16 | 56.1310.59 | 38.8913.98 | 37.439.05 | 37.2110.11 | 46.677.7 | |||||||||
| Vikhr 7B IT 0.4 | 70.9911.4 | 64.122.52 | 67.065.9 | 40.747.64 | 50.7511.53 | 44.074.72 | 55.574.06 | |||||||||
| Vikhr 7B IT 5.4 | 54.328.73 | 55.55.73 | 54.796.93 | 35.197.64 | 34.37.65 | 34.587.06 | 44.696.59 | |||||||||
| VikhrGemma 2B IT | 58.6425.16 | 61.416.24 | 57.6718.04 | 49.0716.79 | 48.6612.71 | 45.686.56 | 51.688.61 | |||||||||
| AL | Gemma2 2B IT | 60.6113.55 | 60.184.51 | 59.668.17 | 53.5110.71 | 54.877.3 | 53.456.06 | 56.565.4 | ||||||||
| Gemma2 9B IT | 46.214.08 | 47.525.48 | 46.794.42 | 40.358.95 | 39.06.04 | 39.67.34 | 43.25.65 | |||||||||
| Qwen2 7B IT | 50.013.38 | 53.2510.65 | 51.411.9 | 50.011.67 | 46.7411.99 | 48.1611.5 | 49.7811.33 | |||||||||
| SaigaLlama3 8B | 53.036.25 | 56.453.35 | 54.323.1 | 51.7511.13 | 48.334.02 | 49.677.13 | 51.993.55 | |||||||||
| Vikhr 7B IT 0.4 | 50.06.94 | 54.875.68 | 52.044.93 | 51.7510.27 | 47.014.37 | 49.016.6 | 50.534.71 | |||||||||
| Vikhr 7B IT 5.4 | 56.068.16 | 57.333.3 | 56.474.96 | 51.756.39 | 50.764.71 | 50.993.89 | 53.733.56 | |||||||||
| VikhrGemma 2B IT | 56.828.6 | 59.32.53 | 57.594.7 | 54.399.92 | 52.13.11 | 52.815.51 | 55.22.95 | |||||||||
| AD | Gemma2 2B IT | 55.8310.17 | 59.215.93 | 56.626.27 | 55.5615.38 | 52.585.69 | 53.369.71 | 54.995.86 | ||||||||
| Gemma2 9B IT | 46.675.53 | 52.034.74 | 49.054.48 | 51.858.28 | 46.54.47 | 48.915.93 | 48.984.5 | |||||||||
| Qwen2 7B IT | 49.1712.72 | 48.12.19 | 47.917.44 | 41.6713.13 | 41.984.13 | 41.087.91 | 44.52.76 | |||||||||
| SaigaLlama3 8B | 53.334.71 | 57.846.15 | 55.223.94 | 55.5612.42 | 51.135.98 | 53.018.89 | 54.125.88 | |||||||||
| Vikhr 7B IT 0.4 | 51.6712.13 | 59.814.42 | 54.667.59 | 61.1111.56 | 53.474.23 | 56.436.19 | 55.554.55 | |||||||||
| Vikhr 7B IT 5.4 | 55.012.58 | 51.918.89 | 53.3210.58 | 44.448.49 | 47.6610.86 | 45.889.46 | 49.69.75 | |||||||||
| VikhrGemma 2B IT | 39.176.72 | 45.934.26 | 42.135.5 | 49.075.93 | 42.093.57 | 45.224.01 | 43.683.91 | |||||||||
| AC | Gemma2 2B IT | 63.74.19 | 58.733.32 | 61.083.43 | 46.935.35 | 52.174.36 | 49.354.71 | 55.223.84 | ||||||||
| Gemma2 9B IT | 61.485.97 | 56.841.85 | 58.852.17 | 44.38.78 | 49.072.0 | 46.215.61 | 52.532.19 | |||||||||
| Qwen2 7B IT | 57.045.97 | 57.363.77 | 57.154.76 | 50.04.02 | 49.734.74 | 49.814.08 | 53.484.25 | |||||||||
| SaigaLlama3 8B | 65.565.7 | 59.023.26 | 62.033.81 | 46.055.84 | 53.164.52 | 49.194.61 | 55.613.74 | |||||||||
| Vikhr 7B IT 0.4 | 60.378.36 | 59.623.49 | 59.785.6 | 51.756.56 | 52.724.82 | 52.04.61 | 55.894.23 | |||||||||
| Vikhr 7B IT 5.4 | 62.968.18 | 61.872.07 | 62.134.3 | 53.957.25 | 55.53.03 | 54.333.58 | 58.232.38 | |||||||||
| VikhrGemma 2B IT | 64.816.21 | 57.772.9 | 60.983.77 | 43.866.39 | 51.44.47 | 47.114.91 | 54.053.57 |
| Corpus | Model | Mode |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| DE | Gemma2 2B IT | 5-shot MMLU | 13.33 | 80.0 | 22.86 | 86.36 | 19.59 | 31.93 | 27.39 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 13.33 | 100.0 | 23.53 | 100.0 | 22.0 | 36.07 | 29.8 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 73.33 | 92.96 | 81.99 | 77.27 | 41.46 | 53.97 | 67.98 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 53.33 | 92.31 | 67.61 | 81.82 | 30.0 | 43.9 | 55.75 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 25.56 | 88.46 | 39.66 | 86.36 | 22.09 | 35.19 | 37.42 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 38.89 | 92.11 | 54.69 | 86.36 | 25.68 | 39.58 | 47.14 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 18.89 | 85.0 | 30.91 | 86.36 | 20.65 | 33.33 | 32.12 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 35.56 | 88.89 | 50.79 | 81.82 | 23.68 | 36.73 | 43.76 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 14.44 | 72.22 | 24.07 | 77.27 | 18.09 | 29.31 | 26.69 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 19.64 | 32.84 | 16.42 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 82.22 | 87.06 | 84.57 | 50.0 | 40.74 | 44.9 | 64.73 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 5.56 | 71.43 | 10.31 | 90.91 | 19.05 | 31.5 | 20.9 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 92.22 | 85.57 | 88.77 | 36.36 | 53.33 | 43.24 | 66.01 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 97.78 | 80.0 | 88.0 | 0.0 | 0.0 | 0.0 | 44.0 | |||||||||
| DSM | Gemma2 2B IT | 5-shot MMLU | 25.93 | 70.0 | 37.84 | 83.33 | 42.86 | 56.6 | 47.22 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 44.44 | 60.0 | 51.06 | 55.56 | 40.0 | 46.51 | 48.79 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 85.19 | 65.71 | 74.19 | 33.33 | 60.0 | 42.86 | 58.53 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 59.26 | 64.0 | 61.54 | 50.0 | 45.0 | 47.37 | 54.45 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 44.44 | 66.67 | 53.33 | 66.67 | 44.44 | 53.33 | 53.33 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 62.96 | 62.96 | 62.96 | 44.44 | 44.44 | 44.44 | 53.7 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 3.7 | 100.0 | 7.14 | 100.0 | 40.91 | 58.06 | 32.6 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 51.85 | 63.64 | 57.14 | 55.56 | 43.48 | 48.78 | 52.96 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 100.0 | 60.0 | 75.0 | 0.0 | 0.0 | 0.0 | 37.5 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 83.33 | 35.71 | 50.0 | 25.0 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 81.48 | 68.75 | 74.58 | 44.44 | 61.54 | 51.61 | 63.09 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 51.85 | 60.87 | 56.0 | 50.0 | 40.91 | 45.0 | 50.5 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 92.59 | 64.1 | 75.76 | 22.22 | 66.67 | 33.33 | 54.55 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 60.0 | 75.0 | 0.0 | 0.0 | 0.0 | 37.5 | |||||||||
| AL | Gemma2 2B IT | 5-shot MMLU | 4.55 | 50.0 | 8.33 | 94.74 | 46.15 | 62.07 | 35.2 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 46.34 | 63.33 | 31.67 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 45.45 | 66.67 | 54.05 | 73.68 | 53.85 | 62.22 | 58.14 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 50.0 | 64.71 | 56.41 | 68.42 | 54.17 | 60.47 | 58.44 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 27.27 | 60.0 | 37.5 | 78.95 | 48.39 | 60.0 | 48.75 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 46.34 | 63.33 | 31.67 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 31.82 | 63.64 | 42.42 | 78.95 | 50.0 | 61.22 | 51.82 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 86.36 | 55.88 | 67.86 | 21.05 | 57.14 | 30.77 | 49.31 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 4.55 | 33.33 | 8.0 | 89.47 | 44.74 | 59.65 | 33.82 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 77.27 | 53.12 | 62.96 | 21.05 | 44.44 | 28.57 | 45.77 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 53.66 | 69.84 | 0.0 | 0.0 | 0.0 | 34.92 | |||||||||
| AD | Gemma2 2B IT | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 10.0 | 50.0 | 16.67 | 88.89 | 47.06 | 61.54 | 39.1 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 15.0 | 50.0 | 23.08 | 83.33 | 46.88 | 60.0 | 41.54 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 35.0 | 77.78 | 48.28 | 88.89 | 55.17 | 68.09 | 58.18 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 70.0 | 73.68 | 71.79 | 72.22 | 68.42 | 70.27 | 71.03 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 20.0 | 80.0 | 32.0 | 94.44 | 51.52 | 66.67 | 49.33 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 85.0 | 53.12 | 65.38 | 16.67 | 50.0 | 25.0 | 45.19 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 25.0 | 83.33 | 38.46 | 94.44 | 53.12 | 68.0 | 53.23 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 10.0 | 33.33 | 15.38 | 77.78 | 43.75 | 56.0 | 35.69 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 52.63 | 68.97 | 0.0 | 0.0 | 0.0 | 34.48 | |||||||||
| AC | Gemma2 2B IT | 5-shot MMLU | 17.78 | 72.73 | 28.57 | 92.11 | 48.61 | 63.64 | 46.1 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 17.78 | 72.73 | 28.57 | 92.11 | 48.61 | 63.64 | 46.1 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 88.89 | 57.14 | 69.57 | 21.05 | 61.54 | 31.37 | 50.47 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 35.56 | 61.54 | 45.07 | 73.68 | 49.12 | 58.95 | 52.01 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 31.11 | 66.67 | 42.42 | 81.58 | 50.0 | 62.0 | 52.21 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 48.89 | 59.46 | 53.66 | 60.53 | 50.0 | 54.76 | 54.21 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 13.33 | 100.0 | 23.53 | 100.0 | 49.35 | 66.09 | 44.81 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 42.22 | 70.37 | 52.78 | 78.95 | 53.57 | 63.83 | 58.3 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 100.0 | 54.22 | 70.31 | 0.0 | 0.0 | 0.0 | 35.16 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 40.0 | 64.29 | 49.32 | 73.68 | 50.91 | 60.22 | 54.77 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 17.78 | 53.33 | 26.67 | 81.58 | 45.59 | 58.49 | 42.58 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 51.11 | 57.5 | 54.12 | 55.26 | 48.84 | 51.85 | 52.98 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 97.78 | 53.66 | 69.29 | 0.0 | 0.0 | 0.0 | 34.65 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 97.78 | 53.66 | 69.29 | 0.0 | 0.0 | 0.0 | 34.65 |
| Corpus | Model | Mode |
F1-healthy |
F1-pathology | F1-macro |
| DE | Gemma2 2B IT | 5-shot | 12.22 | 84.62 | 21.36 | 90.91 | 20.41 | 33.33 | 18.23 | 99.11 | ||||||||||
| Gemma2 2B IT | 0-shot | 25.56 | 95.83 | 40.35 | 95.45 | 23.86 | 38.18 | 39.27 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 55.56 | 92.59 | 69.44 | 81.82 | 31.03 | 45.0 | 57.22 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 62.22 | 93.33 | 74.67 | 81.82 | 34.62 | 48.65 | 61.66 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 25.56 | 88.46 | 39.66 | 86.36 | 22.09 | 35.19 | 37.42 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 41.11 | 90.24 | 56.49 | 81.82 | 25.35 | 38.71 | 47.6 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 18.89 | 85.0 | 30.91 | 86.36 | 20.65 | 33.33 | 32.12 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 8.89 | 100.0 | 16.33 | 100.0 | 21.15 | 34.92 | 25.62 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 13.33 | 70.59 | 22.43 | 77.27 | 17.89 | 29.06 | 25.74 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 100.0 | 19.64 | 32.84 | 16.42 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 82.22 | 87.06 | 84.57 | 50.0 | 40.74 | 44.9 | 64.73 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 5.56 | 71.43 | 10.31 | 90.91 | 19.05 | 31.5 | 20.9 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 67.78 | 88.41 | 76.73 | 63.64 | 32.56 | 43.08 | 59.9 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 5.56 | 83.33 | 10.42 | 90.91 | 19.8 | 32.52 | 14.31 | 95.54 | |||||||||||
| DSM | Gemma2 2B IT | 5-shot | 25.93 | 70.0 | 37.84 | 77.78 | 41.18 | 53.85 | 30.56 | 97.78 | ||||||||||
| Gemma2 2B IT | 0-shot | 40.74 | 64.71 | 50.0 | 66.67 | 42.86 | 52.17 | 51.09 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 44.44 | 66.67 | 53.33 | 66.67 | 44.44 | 53.33 | 53.33 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 77.78 | 61.76 | 68.85 | 27.78 | 45.45 | 34.48 | 51.67 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 44.44 | 70.59 | 54.55 | 72.22 | 46.43 | 56.52 | 55.53 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 70.37 | 65.52 | 67.86 | 44.44 | 50.0 | 47.06 | 57.46 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 3.7 | 100.0 | 7.14 | 100.0 | 40.91 | 58.06 | 32.6 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 25.93 | 58.33 | 35.9 | 72.22 | 39.39 | 50.98 | 43.44 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 81.48 | 70.97 | 75.86 | 50.0 | 64.29 | 56.25 | 66.06 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 83.33 | 35.71 | 50.0 | 25.0 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 81.48 | 68.75 | 74.58 | 44.44 | 61.54 | 51.61 | 63.09 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 48.15 | 59.09 | 53.06 | 50.0 | 39.13 | 43.9 | 48.48 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 37.04 | 71.43 | 48.78 | 77.78 | 45.16 | 57.14 | 52.96 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 14.81 | 57.14 | 23.53 | 72.22 | 41.94 | 53.06 | 25.53 | 84.44 | |||||||||||
| AL | Gemma2 2B IT | 5-shot | 18.18 | 57.14 | 27.59 | 84.21 | 47.06 | 60.38 | 43.98 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 31.82 | 63.64 | 42.42 | 78.95 | 50.0 | 61.22 | 51.82 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 27.27 | 75.0 | 40.0 | 89.47 | 51.52 | 65.38 | 52.69 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 63.64 | 66.67 | 65.12 | 63.16 | 60.0 | 61.54 | 63.33 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 27.27 | 66.67 | 38.71 | 84.21 | 50.0 | 62.75 | 50.73 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 95.45 | 60.0 | 73.68 | 26.32 | 83.33 | 40.0 | 56.84 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 4.55 | 100.0 | 8.7 | 100.0 | 47.5 | 64.41 | 36.55 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 9.09 | 50.0 | 15.38 | 89.47 | 45.95 | 60.71 | 38.05 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 45.45 | 50.0 | 47.62 | 47.37 | 42.86 | 45.0 | 46.31 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 4.55 | 33.33 | 8.0 | 89.47 | 44.74 | 59.65 | 33.82 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 27.27 | 66.67 | 38.71 | 84.21 | 50.0 | 62.75 | 50.73 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | 100.0 | |||||||||||
| AD | Gemma2 2B IT | 5-shot | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 45.0 | 64.29 | 52.94 | 72.22 | 54.17 | 61.9 | 57.42 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 70.0 | 60.87 | 65.12 | 50.0 | 60.0 | 54.55 | 59.83 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 80.0 | 69.57 | 74.42 | 61.11 | 73.33 | 66.67 | 70.54 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 30.0 | 54.55 | 38.71 | 72.22 | 48.15 | 57.78 | 48.24 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 30.0 | 54.55 | 38.71 | 72.22 | 48.15 | 57.78 | 48.24 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 25.0 | 83.33 | 38.46 | 94.44 | 53.12 | 68.0 | 53.23 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 20.0 | 66.67 | 30.77 | 88.89 | 50.0 | 64.0 | 47.38 | 100.0 | |||||||||||
| AC | Gemma2 2B IT | 5-shot | 22.22 | 66.67 | 33.33 | 86.84 | 49.25 | 62.86 | 32.06 | 98.8 | ||||||||||
| Gemma2 2B IT | 0-shot | 15.56 | 87.5 | 26.42 | 97.37 | 49.33 | 65.49 | 45.95 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 15.56 | 100.0 | 26.92 | 100.0 | 50.0 | 66.67 | 46.79 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 26.67 | 54.55 | 35.82 | 73.68 | 47.46 | 57.73 | 31.18 | 97.59 | |||||||||||
| Qwen2 7B IT | 5-shot | 15.56 | 77.78 | 25.93 | 94.74 | 48.65 | 64.29 | 45.11 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 48.89 | 56.41 | 52.38 | 55.26 | 47.73 | 51.22 | 51.8 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 13.33 | 100.0 | 23.53 | 100.0 | 49.35 | 66.09 | 44.81 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 46.67 | 67.74 | 55.26 | 73.68 | 53.85 | 62.22 | 58.74 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 88.89 | 54.79 | 67.8 | 13.16 | 50.0 | 20.83 | 44.31 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 37.78 | 65.38 | 47.89 | 76.32 | 50.88 | 61.05 | 54.47 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 17.78 | 53.33 | 26.67 | 81.58 | 45.59 | 58.49 | 42.58 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 51.11 | 57.5 | 54.12 | 55.26 | 48.84 | 51.85 | 52.98 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 88.89 | 54.79 | 67.8 | 13.16 | 50.0 | 20.83 | 44.31 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 22.22 | 62.5 | 32.79 | 73.68 | 50.0 | 59.57 | 30.79 | 86.75 |
Appendix C Full Results on D-all and A-all Datasets
Tables 12, 13 and 14 show results for all models on D-all and A-all datasets. Among models with SFT Gemma2 2B IT is the best on D-all, while on A-all the best are encoder-based models – RuRoBERTa and BERT. While in general prompting performs worse than SFT on D-all, the best model (Gemma 2 9B IT) shows comparable performance with AutoML models. On the contrary, on A-all prompting shows better results than SFT. The best models in various prompting settings are Qwen2 7B IT, SaigaLlama3 8B, and Gemma2 2B IT.
| Corpus | Model |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| D-all | Gemma2 2B IT | 97.141.04 | 92.330.94 | 94.670.72 | 52.986.33 | 76.446.36 | 62.335.6 | 78.53.13 | ||||||||
| Gemma2 9B IT | 97.650.82 | 91.581.24 | 94.510.68 | 47.628.42 | 77.866.23 | 58.696.91 | 76.63.77 | |||||||||
| Qwen2 7B IT | 97.030.96 | 91.090.64 | 93.960.17 | 44.644.94 | 72.975.26 | 54.982.77 | 74.471.36 | |||||||||
| SaigaLlama3 8B | 96.731.49 | 90.881.0 | 93.711.01 | 43.456.65 | 70.1610.04 | 53.417.35 | 73.564.13 | |||||||||
| Vikhr 7B IT 0.4 | 96.831.2 | 90.461.23 | 93.531.01 | 40.488.16 | 68.8611.28 | 50.738.61 | 72.134.79 | |||||||||
| Vikhr 7B IT 5.4 | 92.021.28 | 90.181.16 | 91.091.16 | 41.677.04 | 47.327.9 | 44.287.35 | 67.694.24 | |||||||||
| VikhrGemma 2B IT | 97.441.68 | 91.830.8 | 94.540.66 | 49.45.98 | 79.2511.48 | 60.043.8 | 77.292.09 | |||||||||
| RuRoBERTa | 92.480.63 | 95.51.16 | 93.960.42 | 68.195.05 | 54.764.45 | 60.482.31 | 77.221.27 | |||||||||
| RuBioRoBERTa | 90.742.52 | 96.931.66 | 93.691.04 | 58.526.99 | 41.6719.2 | 48.6122.33 | 71.1511.6 | |||||||||
| RuBERT | 92.260.52 | 96.221.56 | 94.190.68 | 72.1210.1 | 52.983.81 | 60.62.87 | 77.391.72 | |||||||||
| BERT | 90.163.31 | 96.631.69 | 93.221.43 | 50.5228.15 | 37.524.38 | 42.526.53 | 67.8613.9 | |||||||||
| Linguistic features | 88.10.6 | 88.13.4 | 88.01.6 | 51.65.8 | 50.74.1 | 50.82.6 | 69.41.9 | |||||||||
| TF-IDF | 89.51.0 | 84.21.0 | 86.80.6 | 47.51.8 | 59.14.2 | 52.62.5 | 69.71.5 | |||||||||
| A-all | Gemma2 2B IT | 57.094.28 | 57.613.21 | 57.273.15 | 51.116.0 | 50.613.62 | 50.774.48 | 54.023.39 | ||||||||
| Gemma2 9B IT | 56.324.04 | 55.441.46 | 55.832.54 | 47.563.14 | 48.471.77 | 47.951.96 | 51.891.63 | |||||||||
| Qwen2 7B IT | 56.133.95 | 56.431.57 | 56.232.59 | 49.782.85 | 49.512.04 | 49.591.86 | 52.911.79 | |||||||||
| SaigaLlama3 8B | 56.324.83 | 58.561.32 | 57.332.79 | 53.783.74 | 51.592.03 | 52.561.83 | 54.951.57 | |||||||||
| Vikhr 7B IT 0.4 | 56.132.43 | 57.083.28 | 56.582.68 | 50.894.95 | 49.923.4 | 50.384.09 | 53.483.32 | |||||||||
| Vikhr 7B IT 5.4 | 55.367.22 | 56.892.4 | 55.954.71 | 51.564.79 | 50.13.07 | 50.652.68 | 53.32.82 | |||||||||
| VikhrGemma 2B IT | 53.262.94 | 54.792.65 | 53.972.34 | 48.894.91 | 47.342.63 | 48.063.59 | 51.012.62 | |||||||||
| RuRoBERTa | 61.533.69 | 60.1513.8 | 59.684.47 | 54.390.97 | 54.4418.08 | 52.2711.87 | 55.973.88 | |||||||||
| RuBioRoBERTa | 55.161.73 | 73.1819.96 | 61.666.68 | 32.9423.36 | 30.2222.11 | 31.3122.37 | 46.488.32 | |||||||||
| RuBERT | 57.181.93 | 51.535.61 | 54.093.67 | 49.691.98 | 55.334.34 | 52.252.28 | 53.172.1 | |||||||||
| BERT | 59.441.74 | 46.933.66 | 52.332.1 | 50.430.98 | 62.674.81 | 55.822.21 | 54.081.08 | |||||||||
| Linguistic features | 53.42.9 | 39.719.3 | 42.017.9 | 46.12.4 | 60.220.4 | 51.07.8 | 46.56.2 | |||||||||
| TF-IDF | 51.80.7 | 40.015.6 | 43.310.8 | 44.32.1 | 56.417.8 | 48.68.5 | 46.02.4 |
| Corpus | Model | Mode |
F1-healthy |
F1-pathology | F1-macro | ||||||||
| D-all | Gemma2 2B IT | 5-shot MMLU | 7.36 | 100.0 | 13.71 | 100.0 | 15.64 | 27.05 | 20.38 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 30.06 | 100.0 | 46.23 | 100.0 | 19.72 | 32.94 | 39.58 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 97.55 | 89.33 | 93.26 | 32.14 | 69.23 | 43.90 | 68.58 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 68.1 | 94.87 | 79.29 | 78.57 | 29.73 | 43.14 | 61.21 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 22.09 | 94.74 | 35.82 | 92.86 | 16.99 | 28.73 | 32.28 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 52.15 | 95.51 | 67.46 | 85.71 | 23.53 | 36.92 | 52.19 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 0.61 | 100.0 | 1.22 | 100.0 | 14.74 | 25.69 | 13.45 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 41.1 | 93.06 | 57.02 | 82.14 | 19.33 | 31.29 | 44.16 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 89.57 | 88.48 | 89.02 | 32.14 | 34.62 | 33.33 | 61.18 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 2.45 | 100.0 | 4.79 | 100.0 | 14.97 | 26.05 | 15.42 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 57.06 | 96.88 | 71.81 | 89.29 | 26.32 | 40.65 | 56.23 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 31.29 | 91.07 | 46.58 | 82.14 | 17.04 | 28.22 | 37.4 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 38.04 | 84.93 | 52.54 | 60.71 | 14.41 | 23.29 | 37.92 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 98.77 | 85.19 | 91.48 | 0.0 | 0.0 | 0.0 | 45.74 | |||||||||
| A-all | Gemma2 2B IT | 5-shot MMLU | 10.34 | 64.29 | 17.82 | 93.33 | 47.3 | 62.78 | 40.3 | ||||||||
| Gemma2 2B IT | 0-shot MMLU | 11.49 | 66.67 | 19.61 | 93.33 | 47.62 | 63.06 | 41.34 | |||||||||
| Gemma2 9B IT | 5-shot MMLU | 73.56 | 57.14 | 64.32 | 36.0 | 54.0 | 43.2 | 53.76 | |||||||||
| Gemma2 9B IT | 0-shot MMLU | 39.08 | 65.38 | 48.92 | 76.0 | 51.82 | 61.62 | 55.27 | |||||||||
| Qwen2 7B IT | 5-shot MMLU | 56.32 | 62.82 | 59.39 | 61.33 | 54.76 | 57.86 | 58.63 | |||||||||
| Qwen2 7B IT | 0-shot MMLU | 62.07 | 62.79 | 62.43 | 57.33 | 56.58 | 56.95 | 59.69 | |||||||||
| SaigaLlama3 8B | 5-shot MMLU | 10.34 | 69.23 | 18.0 | 94.67 | 47.65 | 63.39 | 40.7 | |||||||||
| SaigaLlama3 8B | 0-shot MMLU | 34.48 | 69.77 | 46.15 | 82.67 | 52.1 | 63.92 | 55.04 | |||||||||
| Vikhr 7B IT 0.4 | 5-shot MMLU | 96.55 | 53.85 | 69.14 | 4.0 | 50.0 | 7.41 | 38.27 | |||||||||
| Vikhr 7B IT 0.4 | 0-shot MMLU | 22.99 | 62.5 | 33.61 | 84.0 | 48.46 | 61.46 | 47.54 | |||||||||
| Vikhr 7B IT 5.4 | 5-shot MMLU | 13.79 | 70.59 | 23.08 | 93.33 | 48.28 | 63.64 | 43.36 | |||||||||
| Vikhr 7B IT 5.4 | 0-shot MMLU | 33.33 | 59.18 | 42.65 | 73.33 | 48.67 | 58.51 | 50.58 | |||||||||
| VikhrGemma 2B IT | 5-shot MMLU | 94.25 | 54.67 | 69.2 | 9.33 | 58.33 | 16.09 | 42.65 | |||||||||
| VikhrGemma 2B IT | 0-shot MMLU | 98.85 | 53.42 | 69.35 | 0.0 | 0.0 | 0.0 | 34.68 |
| Corpus | Model | Mode |
F1-healthy |
F1-pathology | F1-macro |
| D-all | Gemma2 2B IT | 5-shot | 19.63 | 94.12 | 32.49 | 92.86 | 16.56 | 28.11 | 30.3 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 42.33 | 97.18 | 58.97 | 92.86 | 21.67 | 35.14 | 47.05 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 31.9 | 98.11 | 48.15 | 96.43 | 19.57 | 32.53 | 40.34 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 76.69 | 93.28 | 84.18 | 67.86 | 33.33 | 44.71 | 64.44 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 21.47 | 94.59 | 35.0 | 92.86 | 16.88 | 28.57 | 31.79 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 54.6 | 94.68 | 69.26 | 82.14 | 23.71 | 36.8 | 53.03 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 5.52 | 100.0 | 10.47 | 100.0 | 15.38 | 26.67 | 18.57 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 15.95 | 96.3 | 27.37 | 96.43 | 16.46 | 28.12 | 27.75 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 39.88 | 92.86 | 55.79 | 82.14 | 19.01 | 30.87 | 43.33 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 2.45 | 100.0 | 4.79 | 100.0 | 14.97 | 26.05 | 15.42 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 57.06 | 96.88 | 71.81 | 89.29 | 26.32 | 40.65 | 56.23 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 31.29 | 91.07 | 46.58 | 82.14 | 17.04 | 28.22 | 37.4 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 18.4 | 93.75 | 30.77 | 92.86 | 16.35 | 27.81 | 29.29 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 17.18 | 96.55 | 29.17 | 92.86 | 16.88 | 28.57 | 19.25 | 95.81 | |||||||||||
| A-all | Gemma2 2B IT | 5-shot | 20.69 | 72.0 | 32.14 | 90.67 | 49.64 | 64.15 | 48.15 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 26.44 | 69.7 | 38.33 | 86.67 | 50.39 | 63.73 | 51.03 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 45.98 | 66.67 | 54.42 | 73.33 | 53.92 | 62.15 | 58.28 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 45.98 | 60.61 | 52.29 | 65.33 | 52.13 | 57.99 | 36.76 | 98.77 | |||||||||||
| Qwen2 7B IT | 5-shot | 37.93 | 63.46 | 47.48 | 74.67 | 50.91 | 60.54 | 54.01 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 67.82 | 60.82 | 64.13 | 49.33 | 56.92 | 52.86 | 58.49 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 14.94 | 68.42 | 24.53 | 92.0 | 48.25 | 63.3 | 43.92 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 33.33 | 63.04 | 43.61 | 77.33 | 50.0 | 60.73 | 52.17 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 78.16 | 54.4 | 64.15 | 24.0 | 48.65 | 32.14 | 48.15 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 21.84 | 63.33 | 32.48 | 85.33 | 48.48 | 61.84 | 47.16 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 13.79 | 70.59 | 23.08 | 93.33 | 48.28 | 63.64 | 43.36 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 33.33 | 59.18 | 42.65 | 73.33 | 48.67 | 58.51 | 50.58 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 42.53 | 56.92 | 48.68 | 62.67 | 48.96 | 54.97 | 34.55 | 99.38 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 16.09 | 60.87 | 25.45 | 82.67 | 48.44 | 61.08 | 28.85 | 93.21 |

### v3

@@
-Mental Disorders Detection in the Era of Large Language Models
+Exploring Large Language Models for Detecting Mental Disorders
 Abstract
 This paper compares the effectiveness of traditional machine learning methods, encoder-based models, and large language models (LLMs) on the task of detecting depression and anxiety.
-Five datasets were considered, each differing in format and the method used to define the target pathology class.
+Five Russian-language datasets were considered, each differing in format and in the method used to define the target pathology class.
 We tested AutoML models based on linguistic features, several variations of encoder-based Transformers such as BERT, and state-of-the-art LLMs as pathology classification models.
 The results demonstrated that LLMs outperform traditional methods, particularly on noisy and small datasets where training examples vary significantly in text length and genre.
-However, psycholinguistic features and encoder-based models can achieve performance comparable to language models when trained on texts from individuals with clinically confirmed depression, highlighting their potential effectiveness in targeted clinical applications.
-Mental Disorders Detection in the Era of Large Language Models
-Gleb Kuzmin2,4 Petr Strepetov3 Maksim Stankevich4 Ivan Smirnov4 Artem Shelmanov1 1MBZUAI 2AIRI 3MIPT 4FRC CSC RAS kuzmin@airi.net strepetov.pa@phystech.edu stankevich@isa.ru ivs@isa.ru artem.shelmanov@mbzuai.ac.ae
+However, psycholinguistic features and encoder-based models can achieve performance comparable to language models when trained on texts from individuals with clinically confirmed depression, highlighting their potential effectiveness in targeted clinical applications.111https://github.com/glkuzi/llm-mental-disorders-detection
+Exploring Large Language Models for Detecting Mental Disorders
+Gleb Kuzmin1,2,3 Petr Strepetov4 Maksim Stankevich3 Natalia Chudova3 Artem Shelmanov6 Ivan Smirnov3,5 1AIRI 2ISP RAS Research Center for Trusted Artificial Intelligence 3FRC CSC RAS 4MIPT 5RUDN University 6MBZUAI kuzmin@airi.net strepetov.pa@phystech.edu stankevich@isa.ru nchudova@gmail.com artem.shelmanov@mbzuai.ac.ae ivs@isa.ru
 1 Introduction
@@
 Although there are studies that work with interviews (Morales and Levitan, 2016; Ringeval et al., 2017) and offline texts (Lynn et al., 2018; Stankevich et al., 2019), in most cases the material for such research comes from social media (Guntuku et al., 2017; Garg, 2023).
-These studies tend to focus on but are not limited to, conditions such as depression, anxiety, stress, suicidality, post-traumatic stress disorder, and anorexia.
+These studies tend to focus on, but are not limited to, conditions such as depression, anxiety, stress, suicidality, post-traumatic stress disorder, and anorexia.
 Unsurprisingly, the methods considered for predicting mental state from text fell into traditional machine learning, using hand-crafted linguistic features, and various forms of deep learning (Zhang et al., 2022).
-The deep learning approach is often more accurate, especially when there are enough data samples, while traditional machine learning produces more interpretative results.
+The deep learning approach is often more accurate, especially when there are enough data samples, while traditional machine learning produces more interpretable results.
 In this paper, we compare the performance of linguistic features, encoder-based models, and large language models (LLMs) on the task of identifying mental disorders.
 We consider two types of mental states, depression and anxiety, and several datasets in Russian that differ in text format and in the way a pathology is detected.
-We are particularly interested in evaluating the performance on the essay dataset because it consists of texts written by patients with clinically diagnosed depression and texts written by volunteers who were considered healthy (Stankevich et al., 2019).
-It is worth pointing out the problem of dividing data into texts written by people with a certain mental pathology.
-In most cases, social media-based studies use self-report, affiliation, or questionnaire data to determine mental health status.
-However, there is a lack of work using clinically validated data, which can vary widely (Chancellor and De Choudhury, 2020; Ernala et al., 2019).
-We believe that verified training data from essays can significantly contribute to building models for the social media domain.
+To the best of our knowledge, this is the first comprehensive study on mental disorder detection in Russian texts, involving a wide range of techniques, including LLMs.
+For our study, we involve a clinically verified dataset of essays, which contains texts written by patients with clinically diagnosed depression as well as texts from healthy volunteers (Stankevich et al., 2019).
+Most studies based on social media rely on self-reports, group affiliations, or questionnaire responses to determine mental health status.
+In contrast, only a few works use clinically validated data, which can differ substantially in quality and reliability (Chancellor and De Choudhury, 2020; Ernala et al., 2019).
 This paper addresses the following research questions:
@@
 RQ2: Do models trained on a dataset of essays in which depression was defined by a clinical diagnosis generalize to social media, where the depression status is defined by a questionnaire?
+RQ3: How do the LLM-generated explanations for detected depression align with the clinician’s perspective?
 Our main contributions are the following:
 We outperformed the existing state-of-the-art depression detection method on one dataset and established classification baselines on three previously unexamined anxiety datasets.
-We conducted a thorough comparison of various groups of models on the depression and anxiety detection tasks in Russian, which could be used for the practitioners in this field for future experiments.
-We investigated the transferability of models from the task with clinical diagnosis as a target to the task with a target obtained as a questionnaire result to alleviate the lack of clinically validated data in the mental disorder detection task.
+We conducted a thorough comparison of various groups of models on the depression and anxiety detection tasks in Russian, which could be used by practitioners in this field for future experiments.
+We explored the transferability of models from tasks using clinical diagnoses as targets to those based on questionnaire-derived labels, aiming to mitigate the scarcity of clinically validated data in mental disorder detection.
+We evaluated and categorized LLM-generated explanations for detected depression from the point of view of clinicians, which could be used for future improvement of LLM-assisted systems in this field.
 2 Related Work
 2.1 Traditional and Advanced ML Methods
-Researchers have employed diverse methodologies and techniques in a comprehensive exploration of depression and anxiety detection across various social media platforms.
+Researchers have employed various methods for detecting depression and anxiety across social media platforms.
 Tadesse et al.
 (2019) consider the Reddit users’ dataset, comparing single and combined feature learning for depression detection.
-N-gram features, Linguistic Inquiry and Word Count (LIWC) dictionary features, and topics from Latent Dirichlet Allocation are considered, showcasing the effectiveness of combined features on the multi-layer perceptron with a remarkable 91% accuracy and 0.93 F1 score.
-In (Shah et al., 2020), the authors leverage NLP methods for depression detection of Reddit users based on their posts.
-GloVe, Word2Vec, and FastText embeddings, as well as handcrafted statistical metadata features and features from the LIWC dictionary, are used for text representation.
-The two-headed model, combining BiLSTM for embedding transformation and a fully connected layer for meta-features, demonstrates superior results with Word2Vec embeddings and meta-features achieving a noteworthy 0.81 F1 score.
+N-gram features, Linguistic Inquiry and Word Count (LIWC) dictionary features, and topics from Latent Dirichlet Allocation are considered, showcasing the effectiveness of combined features on the multi-layer perceptron.
+In (Shah et al., 2020), NLP methods are applied for depression detection of Reddit users based on their posts.
+GloVe, Word2Vec, and FastText embeddings, as well as handcrafted statistical metadata features and LIWC features, are used for text representation.
+The two-headed model, combining BiLSTM for embeddings and a fully connected layer for meta-features, demonstrates superior results with Word2Vec embeddings and meta-features.
 Additionally, the authors use Early Risk Detection Error and Latency metrics to take the time of classification into account.
-The study (Owen et al., 2020) considers depression and anxiety detection for Twitter posts.
-SVM on TF-IDF vectors and GloVe embeddings, along with BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2019), are explored, showcasing the nuanced performance of BERT on a balanced dataset and better performance of SVM on an unbalanced one.
-In a comprehensive survey (Babu and Kanaga, 2021), the importance of emoticons in texts in sentiment analysis for depression detection is underscored.
+Owen et al.
+(2020) consider depression and anxiety detection for tweets, using SVM on TF-IDF vectors and GloVe embeddings, as well as BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2020).
+Results show that BERT is better on a balanced dataset, while SVM excels on an unbalanced one.
+Babu and Kanaga (2022) underscores the importance of emoticons in texts in sentiment analysis for depression detection.
 The study covers 101 publications, emphasizing the effectiveness of combining deep learning algorithms, with CNNLSTM yielding the highest precision.
-Moreover, multi-class classification, categorizing sentiments into subclasses based on polarities, provides more precise results than binary and ternary classification.
-In (Tejaswini et al., 2022), the authors introduce the FCL (FasttextCNNLSTM) model, comparing it with LSTM and CNN models on Glove and Word2vec embeddings.
-They also use datasets containing Reddit and Twitter posts.
-FCL outperforms other models in identifying depression in both small and large datasets.
-Assessing the anxious Twitter posts caused by the COVID-19 pandemic, (Jeong et al., 2023) delves into anxiety detection using BERT trained on the Korean language, achieving a 0.67 macro F1-score and 0.91 accuracy and establishing a correlation between the anxiety index and COVID-19 waves.
+Moreover, multi-class sentiment classification provides more precise results than binary and ternary classifications.
+The FCL (FasttextCNNLSTM) model proposed in (Tejaswini et al., 2024) outperforms LSTM and CNN models, based on GloVe and Word2vec embeddings, on datasets for depression detection with Reddit and Twitter posts.
+Assessing the anxious Twitter posts caused by the COVID-19 pandemic, (Jeong et al., 2023) uses BERT trained on the Korean language, achieving strong accuracy and establishing a correlation between the anxiety index and COVID-19 waves.
 The study (Ansari et al., 2023) focuses on identifying depression in social media datasets (CLPsych, Reddit, eRisk) through various text classification methods, combining sentiment lexicons with deep learning pipelines.
 Authors utilize sentiment lexicons with logistic regression and LSTM with attention on GloVe embeddings, comparing them and combining them into ensembles.
-Ensemble models outperform hybrid lexicon- and DL-based models on all datasets, achieving 75% accuracy and 0.77 F1 score on the Reddit dataset.
 2.2 Large Language Models
-According to the systematic review (Omar and Levkovich, 2024), most studies about depression detection focus on BERT-based models, indicating the field’s early stages in adopting newer technologies like GPT-4 and Google’s Gemini.
-However, LLMs are demonstrating significant potential in enhancing the capabilities of depression detection systems.
+According to the systematic review (Omar and Levkovich, 2025), most studies about depression detection focus on BERT-based models, indicating the field’s early stages in adopting newer technologies like GPT-4 and Google’s Gemini.
+However, LLMs are demonstrating significant potential in advancing depression detection systems.
 The Chat-Diagnose approach (Qin et al., 2023) integrates diagnostic criteria from the Diagnostic and Statistical Manual of Mental Disorders (DSM-5) into prompts and uses the Chain of Thoughts technique to deliver explainable diagnoses via an LLMs-augmented system based on ChatGPT/GPT-3.
-This method demonstrates state-of-the-art results on Twitter and Weibo depression datasets by employing zero-shot and few-shot learning, using prompts with selected social media content alongside predictions from a Time2VecTransformer model to enhance diagnostic accuracy.
+This method demonstrates state-of-the-art results on Twitter and Weibo depression datasets by employing zero-shot and few-shot learning.
 Another study (Hadzic et al., 2024) compares the effectiveness of fine-tuned BERT with GPT-3.5 and GPT-4 in the depression detection task.
@@
 With scores separated into depressive and non-depressive groups, the zero-shot method for GPT-4 outperforms GPT-3.5 and BERT across all datasets, highlighting the potential of LLMs in depression detection.
-Additionally, research (Wang et al., 2024) on eRisk 2023 and eRisk 2021 datasets explores detecting depression symptoms and classifying depression severity using LLMs.
+Additionally, Wang et al.
+(2024) investigates depression symptom detection and severity classification using LLMs on the eRisk 2021 and eRisk 2023 datasets.
 Utilizing Beck’s Depression Inventory to form queries related to depression symptoms and the Universal Sentence Encoder for text embeddings, the study creates two datasets containing top-1 and top-5 ranked texts for each query.
-LLMs fine-tuned with QLoRA, including Llama-2-13b-chat, SUS-Chat-34B, and Neural-chat-7b-v3, are used for classification into four levels of depression severity.
-The top-five dataset performs better overall, with the combined model neural-chat+SUS-Chat_top5 excelling in Average Hit Rate and Average Closeness Rate, while neural-chat_top5 and SUS-Chat_top5 achieve the highest Average DODL and Depression Category Hit Rate among single models.
+LLMs fine-tuned with QLoRA are used for classification into four levels of depression severity.
 The DORIS (Lan et al., 2024) system addresses the challenges of detecting depression through social media posts from the Sina Weibo Depression Dataset.
-The authors use GPT3.5-Turbo-1103 for annotating high-risk texts according to the DSM-5 depression scale, also LLM is used to summarize critical information from users’ historical mood records (mood courses).
-The final model based on XGBoost is learned on features from annotations and gte-small-zh model vector representations of post histories and mood courses.
-It shows an improvement in the AUPRC metric by 0.036 compared to baseline methods.
-This study underscores the system’s potential for early detection and intervention, providing a valuable tool for mental health assessment on social media platforms.
-We are the first to examine and compare three generations of the discussed models for depression and anxiety detection tasks in Russian.
-Unlike other works, we used different models from each group and carefully compared the results of the models between the groups on five datasets, aiming for a general recommendation on the best models to use in practice.
+The authors use GPT3.5-Turbo-1103 for annotating high-risk texts according to the DSM-5 depression scale; also, LLM is used to summarize critical information from users’ historical mood records (mood courses).
+The final model based on XGBoost is learned on features from annotations and gte-small-zh model vector representations of post histories and mood courses and shows an improvement over the baseline.
+We are the first to examine and compare three generations of the discussed models for depression and anxiety detection tasks in Russian, namely, traditional ML models, encoder-decoder models, and LLMs.
+Unlike other works, we used various models from each group and carefully compared the results of the models between the groups on five datasets, aiming for a general recommendation on the best models to use in practice.
 3 Data
-Classes in all datasets were represented in the binary format: a healthy class and a pathology class. The general description of the datasets used in our study is shown in Table 1.
-| Label | DE | DSM | AL | AD | AC | ||
+This paper considers five Russian-language datasets: 2 for depression and 3 for anxiety.
+Classes in all datasets were represented in the binary format: a healthy class (no signs of mental disorders) and a pathology class (depression or anxiety).
+The general description of the datasets used in our study is shown in Table 1.
+| Name | Depression-Essays (DE) | Depression-Social Media (DSM) | Anxiety-Letter (AL) | Anxiety-Description (AD) | Anxiety-COVID Comments (AC) | ||
+| Reference | (Stankevich et al., 2019) | (Ignatiev et al., 2022) | (Litvinova and Ryzhkova, 2018) | (Medvedeva et al., 2021) | |||
 | Condition | Depression | Depression | Anxiety | Anxiety | Anxiety | ||
@@
 309 | 122 | 144 | 84 | 7 | ||
-| Reference | (Stankevich et al., 2019) | (Ignatiev et al., 2022) | (Litvinova and Ryzhkova, 2018) | (Medvedeva et al., 2021) | |||
-3.1 Depression-Essays
-To compile this dataset, subjects were asked to write a short essay (up to 5,000 characters) on the topic of “Myself, others, world” (Stankevich et al., 2019).
-A total of 557 essays were collected, where 110 of them were written by people with clinical depression.
+3.1 Depression-Essays (DE)
+To compile this dataset, subjects were asked to write a short essay (from 1,500 to 5,000 characters with spaces and punctuation) on the topic of “Myself, others, world” (Stankevich et al., 2019).
+In total, 557 essays were collected, including 110 authored by patients with clinical depression.
 The essays written by people with clinically validated depression were provided by the Mental Health Research Center, Moscow, Russia.
-The label DE (Depression-Essay) is further used to describe this dataset.
-The best-reported performance on this data reaches 73% F1 for the depression class in the cross-validation evaluation with the random forest model trained on n-grams and psycholinguistic features.
-3.2 Depression-Social Media
-The other depression dataset under consideration was created from data collected to study the problem of identifying depression in social network users (Ignatiev et al., 2022).
-The best-reported classification performance using textual data reaches only 65% F1 for the depression class with a logistic regression model trained on psycholinguistic features.
-This dataset contains text messages from the social network VKontakte and results from the Beck Depression Inventory (BDI) questionnaire (Beck et al., 1987).
+The collection of essays was done on a voluntary basis, under conditions of anonymity and for research purposes only.
+The best-reported performance on this data reaches 73% F1 for the depression class in the cross-validation evaluation with the random forest model trained on n-grams and psycholinguistic features (Stankevich et al., 2019).
+For this dataset, we provided several anonymized examples in Table˜21 in Appendix˜G.
+3.2 Depression-Social Media (DSM)
+The Depression-Social Media dataset was developed to support research on detecting depression in social network users (Ignatiev et al., 2022).
+It consists of text messages from the VKontakte platform, accompanied by results from the Russian-language adaptation of the 21-item Beck Depression Inventory (BDI) questionnaire (Beck et al., 1961).
 The healthy class included users with a scale score of 10 or less, and the pathological class between 30 and 63.
+The best-reported classification performance using textual data reaches only 65% F1 for the depression class with a logistic regression model trained on psycholinguistic features (Ignatiev et al., 2022).
 We have made some changes to the original dataset.
-For each subject, all messages were combined for a period of 170 days prior to the date of the questionnaire screening, and the total text was limited to 6,000 characters.
+For each subject, all messages were combined for a period of 170 days prior to the date of the questionnaire screening, and the total text was limited to 6,000 characters (symbols), resulting in a median length per sample of 122 words.
 Such restrictions were imposed to bring the final texts from this dataset closer to the format of the essay dataset in terms of total text length for each subject and to account for the fact that the relevance of depression screening becomes less over a period of more than six months.
-The label DSM (Depression-Social Media) is used hereafter to describe this dataset.
-3.3 Anxiety-Letter and Picture Description
-The RusNeuroPsych corpus (Litvinova and Ryzhkova, 2018) was used as the data for creating the anxiety detection model.
+3.3 Anxiety-Letter and Anxiety Picture Description (AL and AD)
+For anxiety detection experiments, we use the RusNeuroPsych corpus (Litvinova and Ryzhkova, 2018).
 This corpus was prepared to study the relationships between a person’s text, personal traits, mental status, and demographic characteristics.
-To compile it, subjects were asked to write an informal letter to a friend and give a textual description of a picture, and then complete a series of psychological questionnaires.
-Among the questionnaires, the Hospital Anxiety and Depression Scale (HADS) (Bjelland et al., 2002) was used.
-Thus, the data were divided using the results of this questionnaire, where the healthy class included subjects with a score of 7 and below, and the pathology class included all subjects with a score of 8 and above.
-Since the corpus contained 2 different types of text, it was divided into 2 separate datasets for the experiments.
-It is worth noting that the results of the HADS questionnaire also allow us to divide subjects into groups based on the presence and absence of depression, but a number of subjects with high depression scale scores was less than 20, which prevented us from conducting experiments to train classification models.
-The labels AL (Anxiety-Letter) for letters and AD (Anxiety-Description) for picture description are further used to describe these datasets.
+To compile the corpus, participants were asked to write an informal letter to a friend, provide a textual description of a picture, and complete a series of psychological questionnaires.
+Among these, the Hospital Anxiety and Depression Scale (HADS) (Bjelland et al., 2002) was employed to assess anxiety levels.
+Based on HADS scores, subjects with a score of 7 or below were assigned to the healthy class, while those with a score of 8 or higher were assigned to the pathology class.
+Because the corpus contains two distinct text types, it was further split into two separate datasets for the experiments.
+We denote them AL (Anxiety-Letter) for letters and AD (Anxiety-Description) for picture descriptions.
 To the best of our knowledge, no classification experiments have been performed on this data before.
-3.4 Anxiety-COVID Comments
+3.4 Anxiety-COVID Comments (AC)
 In addition, we used a corpus of subjects’ comments on the COVID-19 pandemic situation (Medvedeva et al., 2021).
@@
 Among the questionnaires used was the SCL-90-R symptom questionnaire (Derogatis, 1983), the anxiety scale from which was used to form 2 groups: the healthy group, with an anxiety scale score below the 33rd percentile, and the pathology group, with an anxiety scale score above the 66th percentile.
-The label AC (Anxiety-Covid) is used hereafter to describe this dataset.
 To the best of our knowledge, no classification experiments have been performed on this data before.
-4 Method
+4 Methods
 4.1 Linguistic Features
 4.1.1 Psycholinguistic Features
-The linguistic features used in our research were calculated using the tool described in (Smirnov et al., 2021).
-This tool allows the calculation of morphological, syntactic, and vocabulary parameters of texts, including various psycholinguistic coefficients.
+The linguistic features used in our research were extracted using the tool described in (Smirnov et al., 2021).
+This tool extracts morphological, syntactic, and vocabulary parameters of texts, including various psycholinguistic coefficients.
 A total of 113 features were used.
-A detailed description of the features used and the features obtained with this tool on our data is given in the Huggingface repository111https://huggingface.co/datasets/anonymizedauthor/paper_data.
+A detailed description of the features utilized and those extracted with this tool on our data is available in the Hugging Face repository.222https://huggingface.co/datasets/anonymizedauthor/paper_data
 4.1.2 Classification Setup
-As a classification baseline, we use AutoML system auto-sklearn (Feurer et al., 2015) and learn it on psycholinguistic features and on n-grams.
-Auto-sklearn classifier employs a Bayesian optimizer which considers 15 classification algorithms, 14 feature preprocessing methods, and 4 data preprocessing methods.
-Also, the classifier utilizes a meta-learning approach and automated ensemble construction to speed up optimization.
-For psycholinguistic features, we consider several methods of feature selection: filter method, wrapper method (forward selection and backward elimination), and embedded method.
-Selected features are examined for linear correlation and ones with absolute Pearson coefficient values more than 0.95 are deleted.
+As a classification baseline, we use the AutoML system auto-sklearn (Feurer et al., 2015) trained on psycholinguistic features and n-grams.
+The auto-sklearn classifier employs a Bayesian optimizer that considers 15 classification algorithms, 14 feature preprocessing methods, and 4 data preprocessing techniques.
+Additionally, the classifier utilizes a meta-learning approach and automated ensemble construction to speed up optimization.
+For psycholinguistic features, we consider several feature selection methods: filter method, wrapper method (forward selection and backward elimination), and embedding method.
+Selected features are examined for pairwise linear correlation, and those with absolute Pearson coefficient values of more than 0.95 are deleted.
 We also train models on TF-IDF vectors on unigrams and unigrams with bigrams.
-In addition to full vectors, we consider different subsets of features for vectors on unigrams: from 20% to 100% of the features are selected with 20% increment.
+In addition to full vectors, we consider different subsets of features for vectors on unigrams: from 20% to 100% of the features are selected with a 20% increment.
 For each subset, correlated features are deleted in the same way as with psycholinguistic features.
-On each set of features we launch the classifier 6 times with different seeds and then calculate the mean and standard deviation values of the metrics based on the results of the launches.
-4.2 Encoder Models
-We employ classification with encoder models to compare methods, based on psycholinguistic features, with more modern solutions.
+On each set of features, we launch the classifier 6 times with different seeds and then calculate the mean and standard deviation values of the metrics based on the results of the launches.
+The chosen AutoML models are presented in Appendix˜A.
+4.2 Encoder-Based Models
+Encoder-based classification models eliminate the need for tedious feature engineering in favor of a deep multi-layer neural architecture.
 As we target classification on Russian corpora, we consider models pretrained on multilingual or Russian datasets.
@@
 Another considered baseline is RuBERT (Kuratov and Arkhipov, 2019), a pretrained Russian version of BERT-base with the same amount of parameters.
-We also finetuned several more recent models, such as RuBioRoBERTa (Yalunin et al., 2022) which is a RoBERTa, pretrained on Russian language biomedical texts, and RuRoberta-large - a bigger version of RoBERTa, also pretrained on Russian language datasets.
-For all of the four models, we optimized hyperparameters using Bayesian optimization from HuggingFace (Wolf et al., 2019) framework.
-The used hyperparameter grid and optimal parameters, alongside with used checkpoint of models, are presented in Appendix A.
+We also finetuned several more recent models, such as RuBioRoBERTa (Yalunin et al., 2022), which is a RoBERTa pretrained on Russian language biomedical texts, and RuRoBERTa-large – a bigger version of RoBERTa, also pretrained on Russian language datasets.
+For all of the four models, we optimized hyperparameters using Bayesian optimization from the HuggingFace framework (Wolf et al., 2020).
+The used hyperparameter grid and optimal parameters, along with the used checkpoint of models, are presented in Appendix˜A.
 4.3 Large Language Models
@@
 We will refer to these settings as “0-shot MMLU” and “5-shot MMLU” correspondingly.
-Secondly, we employed 0-shot and 5-shot prompting with bigger generation lengths and matched the generated answer to one of the possible classes.
-Finally, we conducted the finetuning of the models using LoRA (Hu et al., 2022).
-We selected a set of relatively small (less than 9B parameters) self-hosted open-source models, either tuned for the Russian language or showing good multilingual capabilities.
-To the first group belongs SaigaLlama3 8B, a version of Llama 3 8B Instruct (Dubey et al., 2024) tuned on several Russian datasets, as well as models from Vikhr family (Nikolich et al., 2024).
+Secondly, we employed 0-shot and 5-shot prompting with bigger generation lengths and matched the generated answer to one of the possible classes with string matching.
+Finally, we conducted the fine-tuning of the models using LoRA (Hu et al., 2022).
+We selected a set of relatively small (less than 9B parameters) self-hosted open-source models, either fine-tuned for the Russian language or showing good multilingual capabilities.
+To the first group belongs SaigaLlama3 8B, a version of Llama 3 8B Instruct (Grattafiori et al., 2024) fine-tuned on several Russian datasets, as well as models from the Vikhr family (Nikolich et al., 2024).
 We used Vikhr 7B Instruct 0.4, Vikhr 7B Instruct 5.4, and Vikhr Gemma 2B Instruct.
-The former two are based on Mistral 7B (Jiang et al., 2023) with vocabulary adaptation for the Russian language followed by additional pre-training and instruction tuning.
+The former two are based on Mistral 7B (Jiang et al., 2023) with vocabulary adaptation for the Russian language, followed by additional pretraining and instruction tuning.
 The latter one is based on Gemma2 2B Instruct (Team et al., 2024), additionally trained on Russian data.
 We also used multilingual models, such as Gemma2 2B Instruct and Gemma2 9B Instruct, as well as Qwen2 7B Instruct (Yang et al., 2024).
-As we are conducting experiments on sensitive data and with private datasets, we did not consider remotely-hosted models (such as GPT4, Claude) due to the possibility of data leaking.
-The full information about used prompts, training hyperparameters, and versions of used models, are presented in Appendix A.
+As we are conducting experiments on sensitive data and with private datasets, we did not consider remotely-hosted models (such as GPT-4, Claude) due to the possibility of data leakage.
+The full information about used prompts, training hyperparameters, and versions of used models is presented in Appendix˜A.
 5 Results
 5.1 Classification Results on Depression and Anxiety Datasets
-The data were divided into training and test samples in an 80% by 20% ratio, with stratification by target variable reduced to binary form.
+The data were divided into training and test samples in an 80% by 20% ratio, with stratification by the target variable reduced to binary form.
 All classification reports in this study show results on the test data.
-The classification report for the best models from each group is presented in Table 2.
+The classification report for the best models from each group is presented in Table˜2.
 | Corpus | Mode | Model |
@@
 F1-pathology | F1-macro | ||||||||
-| DE | SFT | Linguistic features | 94.00.8 | 95.21.0 | 94.60.8 | 79.34.0 | 75.03.5 | 77.03.3 | 85.82.1 | ||||||||
-| SFT | TF-IDF | 91.62.1 | 94.42.2 | 93.01.2 | 74.87.0 | 64.410.3 | 68.56.6 | 80.73.8 | |||||||||
-| 5-shot | Vikhr 7B IT 5.4 | 87.060.0 | 82.220.0 | 84.570.0 | 40.740.0 | 50.00.0 | 44.90.0 | 64.730.0 | |||||||||
-| 5-shot MMLU | Gemma2 9B IT | 92.960.0 | 73.330.0 | 81.990.0 | 41.460.0 | 77.270.0 | 53.970.0 | 67.980.0 | |||||||||
+| DE | SFT | Linguistic features | 94.00 0.80 | 95.20 1.00 | 94.60 0.80 | 79.30 4.00 | 75.00 3.50 | 77.00 3.30 | 85.80 2.10 | ||||||||
+| SFT | TF-IDF | 91.60 2.10 | 94.40 2.20 | 93.00 1.20 | 74.80 7.00 | 64.40 10.30 | 68.50 6.60 | 80.70 3.80 | |||||||||
+| 5-shot | Vikhr 7B IT 5.4 | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | |||||||||
+| 5-shot MMLU | Gemma2 9B IT | 93.11 0.08 | 75.11 0.89 | 83.15 0.58 | 43.16 0.85 | 77.27 0.00 | 55.38 0.71 | 69.26 0.64 | |||||||||
 | LoRA | VikhrGemma 2B IT | 94.74 0.88 | 96.48 1.62 | 95.59 0.66 | 84.96 5.72 | 78.03 4.08 | 81.13 2.42 | 88.36 1.52 | |||||||||
-| SFT | RuBERT | 94.451.16 | 91.111.7 | 92.741.04 | 68.414.04 | 78.034.85 | 72.83.43 | 82.772.21 | |||||||||
-| 0-shot | Gemma2 9B IT | 93.330.0 | 62.220.0 | 74.670.0 | 34.620.0 | 81.820.0 | 48.650.0 | 61.660.0 | |||||||||
-| DSM | SFT | Linguistic features | 62.81.8 | 59.37.4 | 60.74.0 | 43.72.7 | 47.27.7 | 45.13.8 | 52.92.0 | ||||||||
-| SFT | TF-IDF | 62.22.6 | 51.213.9 | 55.38.5 | 42.93.6 | 53.711.4 | 46.94.7 | 51.13.5 | |||||||||
-| 5-shot | SaigaLlama3 8B | 100.00.0 | 3.70.0 | 7.140.0 | 40.910.0 | 100.00.0 | 58.060.0 | 32.60.0 | |||||||||
-| 5-shot | Vikhr 7B IT 0.4 | 70.970.0 | 81.480.0 | 75.860.0 | 64.290.0 | 50.00.0 | 56.250.0 | 66.060.0 | |||||||||
-| 5-shot MMLU | SaigaLlama3 8B | 100.00.0 | 3.70.0 | 7.140.0 | 40.910.0 | 100.00.0 | 58.060.0 | 32.60.0 | |||||||||
-| 5-shot MMLU | Vikhr 7B IT 5.4 | 68.750.0 | 81.480.0 | 74.580.0 | 61.540.0 | 44.440.0 | 51.610.0 | 63.090.0 | |||||||||
-| LoRA | Qwen2 7B IT | 68.876.39 | 72.228.21 | 70.46.75 | 55.3311.08 | 50.9310.84 | 52.8110.26 | 61.618.32 | |||||||||
-| SFT | RuBioRoBERTa | 62.12.98 | 62.966.05 | 62.464.3 | 43.664.43 | 42.594.14 | 43.013.7 | 52.743.67 | |||||||||
-| AL | SFT | Linguistic features | 47.421.4 | 36.417.8 | 40.919.2 | 48.52.7 | 68.414.6 | 56.13.7 | 48.58.2 | ||||||||
-| SFT | TF-IDF | 61.55.0 | 46.23.1 | 52.62.3 | 51.22.5 | 65.87.9 | 57.54.5 | 55.002.9 | |||||||||
-| 5-shot | Gemma2 9B IT | 75.00.0 | 27.270.0 | 40.00.0 | 51.520.0 | 89.470.0 | 65.380.0 | 52.690.0 | |||||||||
-| LoRA | Gemma2 2B IT | 60.184.51 | 60.6113.55 | 59.668.17 | 54.877.3 | 53.5110.71 | 53.456.06 | 56.565.4 | |||||||||
-| SFT | RuRoBERTa | 56.825.65 | 75.07.76 | 64.495.59 | 52.5410.76 | 33.3312.77 | 40.1712.48 | 52.338.31 | |||||||||
-| 0-shot | Gemma2 9B IT | 66.670.0 | 63.640.0 | 65.120.0 | 60.00.0 | 63.160.0 | 61.540.0 | 63.330.0 | |||||||||
-| 0-shot MMLU | Gemma2 9B IT | 64.710.0 | 50.00.0 | 56.410.0 | 54.170.0 | 68.420.0 | 60.470.0 | 58.440.0 | |||||||||
-| 0-shot MMLU | Gemma2 2B IT | 0.00.0 | 0.00.0 | 0.00.0 | 46.340.0 | 100.00.0 | 63.330.0 | 31.670.0 | |||||||||
-| AD | SFT | Linguistic features | 50.83.3 | 43.312.5 | 45.38.1 | 44.72.6 | 51.915.9 | 47.37.4 | 46.31.8 | ||||||||
-| SFT | TF-IDF | 54.915.1 | 37.54.8 | 43.22.5 | 44.67.8 | 59.321.0 | 50.412.7 | 46.87.1 | |||||||||
+| SFT | RuBERT | 94.45 1.16 | 91.11 1.70 | 92.74 1.04 | 68.41 4.04 | 78.03 4.85 | 72.80 3.43 | 82.77 2.21 | |||||||||
+| 0-shot | Gemma2 9B IT | 93.33 0.00 | 62.22 0.00 | 74.67 0.00 | 34.62 0.00 | 81.82 0.00 | 48.65 0.00 | 61.66 0.00 | |||||||||
+| DSM | SFT | Linguistic features | 62.80 1.80 | 59.30 7.40 | 60.70 4.00 | 43.70 2.70 | 47.20 7.70 | 45.10 3.80 | 52.90 2.00 | ||||||||
+| SFT | TF-IDF | 62.20 2.60 | 51.20 13.90 | 55.30 8.50 | 42.90 3.60 | 53.70 11.40 | 46.90 4.70 | 51.10 3.50 | |||||||||
+| 5-shot | SaigaLlama3 8B | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | |||||||||
+| 5-shot | Vikhr 7B IT 0.4 | 69.19 0.89 | 81.48 0.00 | 74.83 0.51 | 62.09 1.10 | 45.56 2.22 | 52.54 1.85 | 63.69 1.18 | |||||||||
+| 5-shot MMLU | SaigaLlama3 8B | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | |||||||||
+| 5-shot MMLU | Vikhr 7B IT 5.4 | 68.75 0.00 | 81.48 0.00 | 74.58 0.00 | 61.54 0.00 | 44.44 0.00 | 51.61 0.00 | 63.09 0.00 | |||||||||
+| LoRA | Qwen2 7B IT | 68.87 6.39 | 72.22 8.21 | 70.40 6.75 | 55.33 11.08 | 50.93 10.84 | 52.81 10.26 | 61.61 8.32 | |||||||||
+| SFT | RuBioRoBERTa | 62.10 2.98 | 62.96 6.05 | 62.46 4.30 | 43.66 4.43 | 42.59 4.14 | 43.01 3.70 | 52.74 3.67 | |||||||||
+| AL | SFT | Linguistic features | 47.40 21.40 | 36.40 17.80 | 40.90 19.20 | 48.50 2.70 | 68.40 14.60 | 56.10 3.70 | 48.50 8.20 | ||||||||
+| SFT | TF-IDF | 61.50 5.00 | 46.20 3.10 | 52.60 2.30 | 51.20 2.50 | 65.80 7.90 | 57.50 4.50 | 55.00 2.90 | |||||||||
+| 5-shot | Gemma2 9B IT | 75.00 0.00 | 27.27 0.00 | 40.00 0.00 | 51.52 0.00 | 89.47 0.00 | 65.38 0.00 | 52.69 0.00 | |||||||||
+| LoRA | Gemma2 2B IT | 60.18 4.51 | 60.61 13.55 | 59.66 8.17 | 54.87 7.30 | 53.51 10.71 | 53.45 6.06 | 56.56 5.40 | |||||||||
+| SFT | RuRoBERTa | 56.82 5.65 | 75.00 7.76 | 64.49 5.59 | 52.54 10.76 | 33.33 12.77 | 40.17 12.48 | 52.33 8.31 | |||||||||
+| 0-shot | Gemma2 9B IT | 66.67 0.00 | 63.64 0.00 | 65.12 0.00 | 60.00 0.00 | 63.16 0.00 | 61.54 0.00 | 63.33 0.00 | |||||||||
+| 0-shot MMLU | Gemma2 2B IT | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 46.34 0.00 | 100.00 0.00 | 63.33 0.00 | 31.67 0.00 | |||||||||
+| 0-shot MMLU | Qwen2 7B IT | 61.03 0.52 | 85.45 1.82 | 71.20 0.99 | 68.73 2.55 | 36.84 0.00 | 47.95 0.64 | 59.58 0.82 | |||||||||
+| AD | SFT | Linguistic features | 50.80 3.30 | 43.30 12.50 | 45.30 8.10 | 44.70 2.60 | 51.90 15.90 | 47.30 7.40 | 46.30 1.80 | ||||||||
+| SFT | TF-IDF | 54.90 15.10 | 37.50 4.80 | 43.20 2.50 | 44.60 7.80 | 59.30 21.00 | 50.40 12.70 | 46.80 7.10 | |||||||||
 | LoRA | Vikhr 7B IT 0.4 | 59.81 4.42 | 51.67 12.13 | 54.66 7.59 | 53.47 4.23 | 61.11 11.56 | 56.43 6.19 | 55.55 4.55 | |||||||||
-| SFT | RuRoBERTa | 57.632.52 | 56.674.71 | 57.073.19 | 52.82.82 | 53.74.14 | 53.172.84 | 55.122.58 | |||||||||
-| 0-shot | Vikhr 7B IT 5.4 | 83.330.0 | 25.00.0 | 38.460.0 | 53.120.0 | 94.440.0 | 68.00.0 | 53.230.0 | |||||||||
-| 0-shot | Qwen2 7B IT | 69.570.0 | 80.00.0 | 74.420.0 | 73.330.0 | 61.110.0 | 66.670.0 | 70.540.0 | |||||||||
-| 0-shot MMLU | Qwen2 7B IT | 73.680.0 | 70.00.0 | 71.790.0 | 68.420.0 | 72.220.0 | 70.270.0 | 71.030.0 | |||||||||
-| AC | SFT | Linguistic features | 64.816.0 | 45.621.0 | 47.019.6 | 48.73.6 | 60.520.1 | 52.67.3 | 49.87.8 | ||||||||
-| SFT | TF-IDF | 56.32.1 | 63.34.2 | 59.52.5 | 48.93.3 | 41.75.6 | 44.94.3 | 52.22.7 | |||||||||
-| 5-shot | Gemma2 9B IT | 100.00.0 | 15.560.0 | 26.920.0 | 50.00.0 | 100.00.0 | 66.670.0 | 46.790.0 | |||||||||
-| 5-shot MMLU | SaigaLlama3 8B | 100.00.0 | 13.330.0 | 23.530.0 | 49.350.0 | 100.00.0 | 66.090.0 | 44.810.0 | |||||||||
-| LoRA | Vikhr 7B IT 5.4 | 61.872.07 | 62.968.18 | 62.134.3 | 55.53.03 | 53.957.25 | 54.333.58 | 58.232.38 | |||||||||
+| SFT | RuRoBERTa | 57.63 2.52 | 56.67 4.71 | 57.07 3.19 | 52.80 2.82 | 53.70 4.14 | 53.17 2.84 | 55.12 2.58 | |||||||||
+| 0-shot | Vikhr 7B IT 5.4 | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | |||||||||
+| 0-shot | Qwen2 7B IT | 67.25 1.16 | 80.00 0.00 | 73.07 0.68 | 71.81 0.76 | 56.67 2.22 | 63.33 1.67 | 68.20 1.17 | |||||||||
+| 0-shot MMLU | Vikhr 7B IT 5.4 | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | |||||||||
+| 0-shot MMLU | Qwen2 7B IT | 70.74 1.47 | 70.00 0.00 | 70.36 0.72 | 67.02 0.70 | 67.78 2.22 | 67.39 1.44 | 68.87 1.08 | |||||||||
+| AC | SFT | Linguistic features | 64.80 16.00 | 45.60 21.00 | 47.00 19.60 | 48.70 3.60 | 60.50 20.10 | 52.60 7.30 | 49.80 7.80 | ||||||||
+| SFT | TF-IDF | 56.30 2.10 | 63.30 4.20 | 59.50 2.50 | 48.90 3.30 | 41.70 5.60 | 44.90 4.30 | 52.20 2.70 | |||||||||
+| 5-shot | Qwen2 7B IT | 84.89 3.56 | 26.22 5.33 | 39.85 6.96 | 52.08 1.72 | 94.74 0.00 | 67.20 1.46 | 53.52 4.21 | |||||||||
+| 5-shot MMLU | SaigaLlama3 8B | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 49.35 0.00 | 100.00 0.00 | 66.09 0.00 | 44.81 0.00 | |||||||||
+| LoRA | Vikhr 7B IT 5.4 | 61.87 2.07 | 62.96 8.18 | 62.13 4.30 | 55.50 3.03 | 53.95 7.25 | 54.33 3.58 | 58.23 2.38 | |||||||||
 | SFT | BERT | 57.01 5.15 | 62.96 16.81 | 58.44 6.53 | 46.07 8.77 | 41.67 22.14 | 41.28 17.15 | 49.86 6.84 | |||||||||
-| 0-shot | SaigaLlama3 8B | 67.740.0 | 46.670.0 | 55.260.0 | 53.850.0 | 73.680.0 | 62.220.0 | 58.740.0 | |||||||||
-| 0-shot MMLU | SaigaLlama3 8B | 70.370.0 | 42.220.0 | 52.780.0 | 53.570.0 | 78.950.0 | 63.830.0 | 58.30.0 |
+| 0-shot | SaigaLlama3 8B | 66.05 0.85 | 46.67 0.00 | 54.69 0.29 | 53.12 0.36 | 71.58 1.05 | 60.98 0.62 | 57.84 0.45 | |||||||||
+| 0-shot MMLU | Vikhr 7B IT 0.4 | 69.15 2.43 | 41.78 0.89 | 52.09 1.39 | 53.04 1.06 | 77.89 2.11 | 63.11 1.45 | 57.60 1.42 |
 5.1.1 DE Dataset
-The best scores of the F1 for the pathology class and F1-macro in this experiment are achieved on the essay dataset (DE) by the finetuned LLM model with 88.4% F1-macro and 81.1% F1 score for the pathology class.
-The model trained on the linguistic features shows similar results with 85.8% F1-macro and 77.0% F1-pathology.
-In general, the three best results on the DE dataset were achieved by finetuned models, which can be linked to the bigger dataset size and to the longer texts in the dataset, which, in turn, are crucial for supervised finetuning.
-For all other datasets methods without finetuning significantly outperform SFT and LoRA.
+The best scores of the F1 for the pathology class and F1-macro in this experiment are achieved on the essay dataset (DE) by the fine-tuned LLM model with 88.4% F1-macro and 81.1% F1 score for the pathology class.
+The model trained on the linguistic features shows comparable results with 85.8% F1-macro and 77.0% F1-pathology.
+In general, the three best results on the DE dataset were achieved by fine-tuned models, which can be linked to the bigger dataset size and to the longer texts in the dataset, which, in turn, are crucial for supervised fine-tuning.
+For all other datasets, various prompting methods without finetuning significantly outperform SFT and LoRA.
 | Corpus | Mode | Model |
@@
 F1-pathology | F1-macro | ||||||||
-| D-all | 5-shot MMLU | Gemma2 9B IT | 89.330.0 | 97.550.0 | 93.260.0 | 69.230.0 | 32.140.0 | 43.90.0 | 68.580.0 | ||||||||
-| LoRA | Gemma2 2B IT | 92.330.94 | 97.141.04 | 94.670.72 | 76.446.36 | 52.986.33 | 62.335.6 | 78.53.13 | |||||||||
-| SFT | RuBERT | 92.260.52 | 96.221.56 | 94.190.68 | 72.1210.1 | 52.983.81 | 60.62.87 | 77.391.72 | |||||||||
-| 0-shot | Gemma2 9B IT | 93.280.0 | 76.690.0 | 84.180.0 | 33.330.0 | 67.860.0 | 44.710.0 | 64.440.0 | |||||||||
-| SFT | Linguistic features | 88.10.6 | 88.13.4 | 88.01.6 | 51.65.8 | 50.74.1 | 50.82.6 | 69.41.9 | |||||||||
-| SFT | TF-IDF | 89.51.0 | 84.21.0 | 86.80.6 | 47.51.8 | 59.14.2 | 52.62.5 | 69.71.5 | |||||||||
-| A-all | 5-shot | Gemma2 2B IT | 72.00.0 | 20.690.0 | 32.140.0 | 49.640.0 | 90.670.0 | 64.150.0 | 48.150.0 | ||||||||
+| D-all | 5-shot MMLU | Gemma2 9B IT | 88.83 0.25 | 96.56 0.49 | 92.53 0.36 | 59.56 4.84 | 29.29 1.43 | 39.26 2.32 | 65.90 1.34 | ||||||||
+| LoRA | Gemma2 2B IT | 92.33 0.94 | 97.14 1.04 | 94.67 0.72 | 76.44 6.36 | 52.98 6.33 | 62.33 5.60 | 78.50 3.13 | |||||||||
+| SFT | RuBERT | 92.26 0.52 | 96.22 1.56 | 94.19 0.68 | 72.12 10.10 | 52.98 3.81 | 60.60 2.87 | 77.39 1.72 | |||||||||
+| 0-shot | Gemma2 9B IT | 93.28 0.00 | 76.69 0.00 | 84.18 0.00 | 33.33 0.00 | 67.86 0.00 | 44.71 0.00 | 64.44 0.00 | |||||||||
+| 0-shot MMLU | Gemma2 9B IT | 94.87 0.00 | 68.10 0.00 | 79.29 0.00 | 29.73 0.00 | 78.57 0.00 | 43.14 0.00 | 61.21 0.00 | |||||||||
+| SFT | Linguistic features | 88.10 0.60 | 88.10 3.40 | 88.00 1.60 | 51.60 5.80 | 50.70 4.10 | 50.80 2.60 | 69.40 1.90 | |||||||||
+| SFT | TF-IDF | 89.50 1.00 | 84.20 1.00 | 86.80 0.60 | 47.50 1.80 | 59.10 4.20 | 52.60 2.50 | 69.70 1.50 | |||||||||
+| A-all | 5-shot MMLU | SaigaLlama3 8B | 80.51 5.64 | 11.26 0.46 | 19.76 0.88 | 48.46 0.41 | 96.80 1.07 | 64.59 0.60 | 42.18 0.74 | ||||||||
 | LoRA | SaigaLlama3 8B | 58.56 1.32 | 56.32 4.83 | 57.33 2.79 | 51.59 2.03 | 53.78 3.74 | 52.56 1.83 | 54.95 1.57 | |||||||||
-| SFT | RuRoBERTa | 61.533.69 | 60.1513.8 | 59.684.47 | 54.390.97 | 54.4418.08 | 52.2711.87 | 55.973.88 | |||||||||
-| SFT | BERT | 59.441.74 | 46.933.66 | 52.332.1 | 50.430.98 | 62.674.81 | 55.822.21 | 54.081.08 | |||||||||
-| 0-shot | Qwen2 7B IT | 60.820.0 | 67.820.0 | 64.130.0 | 56.920.0 | 49.330.0 | 52.860.0 | 58.490.0 | |||||||||
-| 0-shot MMLU | SaigaLlama3 8B | 69.770.0 | 34.480.0 | 46.150.0 | 52.10.0 | 82.670.0 | 63.920.0 | 55.040.0 | |||||||||
-| 0-shot MMLU | Qwen2 7B IT | 62.790.0 | 62.070.0 | 62.430.0 | 56.580.0 | 57.330.0 | 56.950.0 | 59.690.0 | |||||||||
-| SFT | Linguistic features | 53.42.9 | 39.719.3 | 42.017.9 | 46.12.4 | 60.220.4 | 51.07.8 | 46.56.2 | |||||||||
-| SFT | TF-IDF | 51.80.7 | 40.015.6 | 43.310.8 | 44.32.1 | 56.417.8 | 48.68.5 | 46.02.4 |
+| SFT | RuRoBERTa | 61.53 3.69 | 60.15 13.80 | 59.68 4.47 | 54.39 0.97 | 54.44 18.08 | 52.27 11.87 | 55.97 3.88 | |||||||||
+| SFT | BERT | 59.44 1.74 | 46.93 3.66 | 52.33 2.10 | 50.43 0.98 | 62.67 4.81 | 55.82 2.21 | 54.08 1.08 | |||||||||
+| 0-shot | Qwen2 7B IT | 60.01 0.41 | 69.66 0.92 | 64.47 0.17 | 56.72 0.10 | 46.13 1.60 | 50.87 0.99 | 57.67 0.41 | |||||||||
+| 0-shot | Gemma2 2B IT | 71.44 0.87 | 26.44 0.00 | 38.59 0.13 | 50.69 0.15 | 87.73 0.53 | 64.26 0.27 | 51.42 0.20 | |||||||||
+| 0-shot MMLU | Qwen2 7B IT | 62.67 0.06 | 64.83 1.38 | 63.72 0.65 | 57.51 0.47 | 55.20 1.07 | 56.32 0.32 | 60.02 0.17 | |||||||||
+| SFT | Linguistic features | 53.40 2.90 | 39.70 19.30 | 42.00 17.90 | 46.10 2.40 | 60.20 20.40 | 51.00 7.80 | 46.50 6.20 | |||||||||
+| SFT | TF-IDF | 51.80 0.70 | 40.00 15.60 | 43.30 10.80 | 44.30 2.10 | 56.40 17.80 | 48.60 8.50 | 46.00 2.40 |
 | Transfer | Mode | Model |
@@
 F1-pathology | F1-macro | ||||||||
-| DE to DSM test | LoRA | SaigaLlama3 8B | 58.493.68 | 83.9513.97 | 68.737.39 | 40.3413.85 | 12.044.99 | 17.215.85 | 42.974.2 | ||||||||
+| DE to DSM test | LoRA | SaigaLlama3 8B | 58.49 3.68 | 83.95 13.97 | 68.73 7.39 | 40.34 13.85 | 12.04 4.99 | 17.21 5.85 | 42.97 4.20 | ||||||||
 | LoRA | Vikhr 7B IT 0.4 | 60.33 5.19 | 61.73 20.91 | 58.81 10.42 | 37.58 5.47 | 37.04 24.57 | 33.98 15.06 | 46.39 4.16 | |||||||||
-| LoRA | Gemma2 2B IT | 64.143.85 | 71.621.45 | 65.8910.13 | 51.078.34 | 38.8922.91 | 38.6716.04 | 52.286.44 | |||||||||
+| LoRA | Gemma2 2B IT | 64.14 3.85 | 71.60 21.45 | 65.89 10.13 | 51.07 8.34 | 38.89 22.91 | 38.67 16.04 | 52.28 6.44 | |||||||||
 | LoRA | Gemma2 9B IT | 64.46 5.24 | 66.67 20.73 | 64.11 11.57 | 51.24 8.98 | 45.37 18.54 | 44.36 11.95 | 54.23 7.12 | |||||||||
-| LoRA | Vikhr 7B IT 5.4 | 58.245.42 | 67.2821.75 | 60.2411.51 | 31.337.8 | 25.9324.36 | 25.0915.4 | 42.664.75 | |||||||||
-| LoRA | Qwen2 7B IT | 61.143.36 | 61.1118.61 | 59.7511.42 | 43.628.81 | 42.5915.93 | 41.258.0 | 50.55.32 | |||||||||
-| LoRA | VikhrGemma 2B IT | 60.842.46 | 71.616.93 | 64.98.82 | 42.547.31 | 31.4815.27 | 33.9712.07 | 49.445.25 | |||||||||
-| SFT | RuRoBERTa | 60.967.01 | 41.9826.59 | 46.0714.36 | 33.7515.28 | 59.2627.34 | 42.9519.5 | 44.515.07 | |||||||||
-| SFT | RuBioRoBERTa | 61.652.51 | 50.023.88 | 52.5412.11 | 34.5415.55 | 52.7824.59 | 41.5718.71 | 47.055.29 | |||||||||
-| SFT | RuBERT | 68.010.29 | 27.789.01 | 38.6810.16 | 42.662.8 | 80.569.49 | 55.644.19 | 47.165.8 | |||||||||
+| LoRA | Vikhr 7B IT 5.4 | 58.24 5.42 | 67.28 21.75 | 60.24 11.51 | 31.33 7.80 | 25.93 24.36 | 25.09 15.40 | 42.66 4.75 | |||||||||
+| LoRA | Qwen2 7B IT | 61.14 3.36 | 61.11 18.61 | 59.75 11.42 | 43.62 8.81 | 42.59 15.93 | 41.25 8.00 | 50.50 5.32 | |||||||||
+| LoRA | VikhrGemma 2B IT | 60.84 2.46 | 71.60 16.93 | 64.90 8.82 | 42.54 7.31 | 31.48 15.27 | 33.97 12.07 | 49.44 5.25 | |||||||||
+| SFT | RuRoBERTa | 60.96 7.01 | 41.98 26.59 | 46.07 14.36 | 33.75 15.28 | 59.26 27.34 | 42.95 19.50 | 44.51 5.07 | |||||||||
+| SFT | RuBioRoBERTa | 61.65 2.51 | 50.00 23.88 | 52.54 12.11 | 34.54 15.55 | 52.78 24.59 | 41.57 18.71 | 47.05 5.29 | |||||||||
+| SFT | RuBERT | 68.00 10.29 | 27.78 9.01 | 38.68 10.16 | 42.66 2.80 | 80.56 9.49 | 55.64 4.19 | 47.16 5.80 | |||||||||
 | SFT | BERT | 76.85 5.46 | 29.63 8.28 | 42.21 9.33 | 45.34 2.29 | 87.04 4.14 | 59.54 2.03 | 50.88 5.43 | |||||||||
-| SFT | Linguistic features | 60.02.5 | 57.47.0 | 58.54.5 | 40.03.5 | 42.66.9 | 41.14.5 | 49.83.1 | |||||||||
-| SFT | TF-IDF | 59.93.4 | 30.99.2 | 40.08.5 | 40.21.5 | 69.48.3 | 50.72.7 | 45.43.8 |
+| SFT | Linguistic features | 60.00 2.50 | 57.40 7.00 | 58.50 4.50 | 40.00 3.50 | 42.60 6.90 | 41.10 4.50 | 49.80 3.10 | |||||||||
+| SFT | TF-IDF | 59.90 3.40 | 30.90 9.20 | 40.00 8.50 | 40.20 1.50 | 69.40 8.30 | 50.70 2.70 | 45.40 3.80 |
 5.1.2 DSM Dataset
-The best result achieved by Vikhr 7B IT 0.4 model, evaluated in 5-shot, with 66.1% F1-macro.
+The best result achieved by the Vikhr 7B IT 0.4 model, evaluated in the 5-shot regime, with 63.69% F1-macro.
 The traditional machine learning methods, as well as encoder models, performed poorly with nearly 53% F1-macro on linguistic features.
@@
 This may be due to several factors.
-First of all, we can note the sample size, which was significantly smaller in the DSM dataset.
-Anyway, almost all machine learning models, regardless of the selected features, can show low accuracy on a small amount of data.
+First of all, the sample size is significantly smaller in the DSM dataset, as machine learning models in general can show low accuracy on limited data.
 On the other hand, if we refer to the study (Ignatiev et al., 2022), where less stringent restrictions on text volume and temporal proximity to questionnaire screening dates were applied to the same raw data, the results were still not very high: about 60% F1-macro for psycholinguistic features and TF-IDF features.
-Secondly, the problem may lie in the format of the texts from the social network.
+Secondly, the text format in the DSM introduces some noise.
 The texts in DSM are concatenated together with a collection of posts from users’ personal pages, and they mostly lack coherent logic and cohesion in the resulting text.
-Even with the 6,000 character limit, the standard deviation in word count in this data is on the order of 300 words, while in DE the similar value is around 100.
+Even with the 6,000-character limit, the standard deviation in word count is approximately 300 words, compared with 100 in DE.
 Thus, DSM is a much noisier dataset, which can strongly affect the quality of classification with psycholinguistic markers, where the values of some markers can be affected by the volume of text analyzed, even considering their normalization with respect to the text volume.
 Finally, the way in which the target class of pathology is defined can be of great importance.
-Although a widely used and validated psychological questionnaire was used for the DSM dataset, it is not possible to directly compare its results with a clinically confirmed diagnosis.
+Although a widely used and validated psychological questionnaire was used for the DSM dataset, its results cannot be compared with a clinically confirmed diagnosis.
 The same findings can be outlined in work that criticizes the way in which social media users are labeled for mental illness by indirectly affiliating or self-identifying mental ill-health (Ernala et al., 2019).
 In favor of the significance of this factor is also the fact that TF-IDF-based features performed significantly better on DE data than on DSM, although, unlike psycholinguistic markers, they do not have an initial specialization for detecting signs of mental ill-health.
-To take into account all of the aforementioned considerations, we applied LLMs to this task and showed, that the bigger model is able to partially overcome these issues.
-While RuBERT and the model trained on linguistic features both show comparatively low results, the evaluated in 5-shot format model performs significantly better.
+To take into account these considerations, we applied LLMs to this task and showed that the bigger model is able to partially overcome these issues.
+While RuBERT and the model trained on linguistic features both show comparatively low results, a 5-shot LLM performs significantly better.
 The same outcomes are observed in the experiments with anxiety datasets.
+For a better understanding of per-class performance of models, we also provided results for the best models on the DSM dataset for various splits within a class based on the initial BDI scores.
+The results are provided in Appendix˜F.
 5.1.3 AL, AD, and AC Datasets
@@
 However, LLMs performed significantly better on these tasks.
-Models with LoRA in general perform better than encoder-based and AutoML-based models, but the best performance is achieved by 0-shot and 5-shot prompting.
-Again, this can be linked to the complexity of the domain and the small amount of training data, which makes it hard to finetune a model.
+Models with LoRA, in general, perform better than encoder-based and AutoML-based models, but the best performance is achieved by 0-shot and 5-shot prompting.
+Again, this can be linked to the complexity of the domain and the small amount of training data, which makes it hard to fine-tune a model.
 In general, the LLMs outperform other models on all datasets, but on the DE and DSM, the gap between various types of models is smaller, leaving the usage of non-LLM models reasonable for some specific cases.
@@
 We combined the depression and anxiety datasets into pooled datasets D-all (DE and DSM) and A-all (AL, AD, and AC).
-Classification performance of the best models on D-all and A-all combined data presented in Table 3.
+Classification performance of the best models on D-all and A-all combined data is presented in Table˜3.
 The LLMs performed better than other models on both D-all and A-all.
 It is also noticeable that methods with model tuning work better on D-all, while prompting shows better results on A-all.
-It can be explained with significant genre and length variations on A-all datasets, so the general-purpose models perform better than finetuned ones due to the complexity of the data.
+It can be explained with significant genre and length variations in A-all datasets, so the general-purpose models perform better than finetuned ones due to the complexity of the data.
 5.3 Classification of DSM Data with Best Models from DE
-The classification results of models that demonstrated the best performance on the DE dataset applied to DSM test data are shown in Table 4.
+The classification results of models that demonstrated the best performance on the DE dataset applied to DSM test data are shown in Table˜4.
 This experiment shows that models trained on essays from depressed subjects cannot be used to detect depression from social media texts from users who have taken a depression questionnaire and received a high score.
@@
 A sample of social network users with clinically diagnosed depression would be needed to clarify this issue.
+5.4 LLM Results Interpretation
+As shown in Section˜5.3, the overall best results were obtained on the DE dataset with LLMs.
+To further analyze the quality of these results from the clinical perspective, we conducted an additional experiment on LLM results interpretation.
+We chose the best LLM in a 5-shot setting on the DE dataset (Vikhr 7B IT 5.4) and asked two psychologists with relevant experience to rate the LLM generation on a scale from 1 to 5.
+The LLM-generated answer in this setting consists of a predicted class label and a detailed explanation of why this label was chosen.
+Details of the rating scale and statistics are provided in Appendix˜D.
+The psychologists assessed only texts from patients with clinically diagnosed depression that the LLM correctly labeled.
+The average score from expert psychologists for LLM explanations was assigned to 2.84 out of 5.
+Moreover, the psychologists also noted each explanation, which can be marked as true to some extent (e.g., with some true claims in the explanation) – there are approximately 66% of such explanations.
+These results show that the explanations for depression detection from texts contain both correct and erroneous parts.
+Overall, the LLM (in the described settings) does not generate enough explanations that meet the requirements of clinicians.
+To further extend the error analysis in the LLM generations in the psychological domain, we asked psychologists to describe the most common errors in the generated explanations.
+To do so, psychologists annotated each explanation with the list of errors and categorized these mistakes into the following groups: (1) tautology, (2) groundless generalization, (3) false conclusion, (4) confabulation, (5) distortion of medical understanding of depression, (6) incompleteness of selected signs of depression.
+Description and examples of these error types are located in Table˜17 in the Appendix˜D.
+It should also be noted that from the general NLP perspective, most of these types of errors can be defined as hallucinations; however, we suppose that a more precise definition is needed in this specific domain.
+On average, each explanation contains two or more errors.
+The most common types of errors are groundless generalization, false conclusion, and confabulation, which occur in 56%, 56%, and 50% of samples, respectively.
+However, these errors appear several times in each explanation.
+The incompleteness of selected signs of depression appears in 44%, while distortion of medical understanding of depression and tautology are the rare errors, which appear in 19% and 13%.
+These results show that there exist various types of errors, specific to the mental health disorders domain.
+The additional results with detailed analysis of feature importance for the pathology and healthy class prediction are located in Appendix˜E.
 6 Conclusion
@@
 In response to RQ2, the findings reveal that models trained on essays from depressed individuals are not effective for detecting depression in social media texts from users who have completed a depression questionnaire and scored high.
-7 Limitations
-The main limitation of this paper is that it is impossible to share all the raw texts from used datasets, as they are all distributed under different terms.
+Finally, we evaluated LLM-generated explanations and showed that in the current state, these explanations do not meet the clinicians’ requirements with an overall score of 2.84 out of 5, which answers RQ3.
+We also constructed a detailed classification of common errors in LLM explanations from the clinicians’ perspective, to guide further improvements of LLMs in this domain.
+Limitations
+The main limitation of this paper is that it is impossible to share all the raw texts from the used datasets, as they are all distributed under different terms.
+However, we present several anonymized textual examples and extracted psycholinguistic features.
 The amount of sample data used for predicting anxiety is small, which does not allow us to adequately judge the possibility of predicting anxiety in Russian text using machine learning methods.
 Although the results on all the anxiety datasets used show poor accuracy, perhaps a very different scale of data is needed for this task.
-The paper does not discuss the differences between the results inside one group of the models in detail as this was not the aim of the paper.
-We conducted the experiments only for the Russian due to the poor availability of related datasets for other languages.
-However, the used methods in general are language agnostic so the results could be extended to other languages.
+The paper does not discuss the differences between the results within one group of the models in detail, as this was not the aim of the paper.
+We conducted the experiments only for Russian due to the poor availability of related datasets for other languages.
+However, the used methods in general are language agnostic, so the results could be extended to other languages.
 The overall result about LLMs as the best-performing models matches similar studies for English on closed data.
-8 Ethical Considerations
+The final limitation is the temporal validity of the results.
+Due to the fast growth of LLM-based solutions, future LLMs could outperform the obtained results.
+However, we suppose that the current state of the field already presents an interest and is therefore investigated in the paper.
+Ethical Considerations
 The problem discussed in this paper is the sensitive issue of mental health.
-To avoid any possible harm, we did not fully open-source used datasets.
+To avoid any possible harm, we did not fully open-source the used datasets.
 All shared data does not contain any information that names or uniquely identifies individuals, nor does it contain offensive content.
-The examined models for the detection of mental disorders do not aim to replace a professional physician, on the contrary, these models are intended to support a human expert.
+The examined models for the detection of mental disorders do not aim to replace a professional physician; on the contrary, these models are intended to support a human expert.
+Acknowledgements
+The work of Gleb Kuzmin was supported by a grant, provided by the MED of Russia (agreement identifier 000000C313925P4G0002) and the agreement with the ISP RAS dated June 20, 2025 No. 139-15-2025-011.
+The research was carried out using the infrastructure of the shared research facilities «High Performance Computing and Big Data » of FRC CSC RAS (CKP «Informatics»)
 Appendix A Hyperparameters
-The used net and optimal parameters, alongside with used checkpoint of encoder models, are presented in Table 6.
-The hyperparameters for LLMs, tuned with LoRA are presented in Table 7.
+The used net and optimal parameters, alongside the used checkpoint of encoder models, are presented in Table˜7.
+For the models trained on the psycholinguistic features and TF-IDF, the best model was selected with the AutoML pipeline, with the most models chosen as a Random Forest ensemble of up to four models.
+The hyperparameters for LLMs, tuned with LoRA, are presented in Table˜8.
 The LoRA was applied to the following target layers: q_proj, up_proj, o_proj, k_proj, down_proj, gate_proj, v_proj.
 For LLMs 0-shot and 5-shot evaluation, we used the following generation parameters: do_sample=False, temperature=1.0, top_p=1.0, top_k=50, repetition_penalty=1.0.
-For MMLU-style evaluation generation length was limited to one token while for regular evaluation it was set to 64 tokens.
-Prompt examples are presented in Table 5.
+The predicted classes were extracted using the string matching algorithm.
+For MMLU-style evaluation, the generation length was limited to one token, while for regular evaluation, it was set to 64 tokens.
+Prompt examples are presented in Table˜5.
+For the 0-shot and 5-shot evaluation, we averaged the results on several prompts with various system parts to take into account the sensitivity of the generation with respect to the prompt.
+These prompt variations are presented in Table˜6.
+In all 0-shot and 5-shot experiments, we did not conduct best-of-N aggregation of answers and used one extracted answer per generation, which was averaged over several prompts to ensure generalizability of results.
+The used hardware, as well as GPU hours, carbon footprint, and memory requirements are presented in Table˜9.
 | Task | Mode | Prompt example (system & user) |
@@
 User: Text: {example_1} Answer (0 or 1): 0 Text: {example_2} Answer (0 or 1): 0 Text: {example_3} Answer (0 or 1): 0 Text: {example_4} Answer (0 or 1): 1 Text: {example_5} Answer (0 or 1): 1 Text: {input_text} Answer (0 or 1): |
+| Task | Prompt number | Prompt example (only system) |
+| Depression | 1 | System: Read the provided text and determine whether the author has signs of depression. Use the scale: 0 - no depression, 1 - depression. Then explain why you chose this option. |
+| Depression | 2 | System: Evaluate the text for the author’s depressive state. Scale: 0 - no depression, 1 - depression is present. Justify your choice after indicating the answer. |
+| Depression | 3 | System: Analyze this text and identify the presence of depressive manifestations in its author.
+Use a binary assessment: 0 - no depression, 1 - presence of depression.
+Provide a rationale for your decision.
+|
+| Depression | 4 | System: Assess the psychological state of the author of the text in the context of depression.
+Use the gradation: 0 - no signs of depression, 1 - there are signs of depression.
+Detail the reasons for making your decision.
+|
+| Anxiety | 1 | System: Read the provided text and determine whether the author has signs of anxiety. Use the scale: 0 - no anxiety, 1 - anxiety. Then explain why you chose this option. |
+| Anxiety | 2 | System: Evaluate the text for the author’s anxiety. Scale: 0 - no anxiety, 1 - anxiety is present. Justify your choice after indicating the answer. |
+| Anxiety | 3 | System: Analyze this text and identify the presence of anxiety in its author. Use a binary assessment: 0 - no anxiety, 1 - anxiety present. Justify your decision. |
+| Anxiety | 4 | System: Assess the psychological state of the author of the text in the context of anxiety.
+Use the gradation: 0 - no signs of anxiety, 1 - signs of anxiety present.
+Detail the reasons for making your decision.
+|
 | Model name | Corpus |
-| RuBERT | A-all | 15 | 5e-05 | 16 | 0.1 | ||||||||
-| D-all | 15 | 5e-05 | 16 | 0.1 | |||||||||
-| DE | 10 | 7e-05 | 32 | 0.1 | |||||||||
-| AD | 2 | 6e-06 | 4 | 0.1 | |||||||||
-| AL | 2 | 9e-06 | 4 | 0.1 | |||||||||
-| AC | 2 | 6e-06 | 4 | 0.1 | |||||||||
-| DSM | 13 | 1e-04 | 4 | 0.0 | |||||||||
+| RuBERT | A-all | 15 | 5e-05 | 16 | 0.10 | ||||||||
+| D-all | 15 | 5e-05 | 16 | 0.10 | |||||||||
+| DE | 10 | 7e-05 | 32 | 0.10 | |||||||||
+| AD | 2 | 6e-06 | 4 | 0.10 | |||||||||
+| AL | 2 | 9e-06 | 4 | 0.10 | |||||||||
+| AC | 2 | 6e-06 | 4 | 0.10 | |||||||||
+| DSM | 13 | 1e-04 | 4 | 0.00 | |||||||||
 | RuRoBERTa | A-all | 13 | 9e-06 | 16 | 0.01 | ||||||||
-| D-all | 11 | 6e-06 | 8 | 0.0 | |||||||||
-| DE | 9 | 9e-06 | 4 | 0.0 | |||||||||
-| AD | 7 | 2e-05 | 4 | 0.0 | |||||||||
-| AL | 11 | 6e-06 | 8 | 0.0 | |||||||||
+| D-all | 11 | 6e-06 | 8 | 0.00 | |||||||||
+| DE | 9 | 9e-06 | 4 | 0.00 | |||||||||
+| AD | 7 | 2e-05 | 4 | 0.00 | |||||||||
+| AL | 11 | 6e-06 | 8 | 0.00 | |||||||||
 | AC | 6 | 1e-04 | 32 | 0.01 | |||||||||
-| DSM | 13 | 5e-06 | 16 | 0.1 | |||||||||
+| DSM | 13 | 5e-06 | 16 | 0.10 | |||||||||
 | RuBioRoBERTa | A-all | 15 | 3e-05 | 16 | 0.01 | ||||||||
-| D-all | 11 | 7e-06 | 32 | 0.0 | |||||||||
-| DE | 8 | 1e-05 | 4 | 0.0 | |||||||||
-| AD | 2 | 7e-06 | 8 | 0.1 | |||||||||
+| D-all | 11 | 7e-06 | 32 | 0.00 | |||||||||
+| DE | 8 | 1e-05 | 4 | 0.00 | |||||||||
+| AD | 2 | 7e-06 | 8 | 0.10 | |||||||||
 | AL | 9 | 2e-05 | 16 | 0.01 | |||||||||
-| AC | 8 | 1e-05 | 4 | 0.0 | |||||||||
+| AC | 8 | 1e-05 | 4 | 0.00 | |||||||||
 | DSM | 15 | 9e-06 | 16 | 0.01 | |||||||||
 | BERT | A-all | 14 | 7e-06 | 16 | 0.01 | ||||||||
-| D-all | 15 | 5e-05 | 16 | 0.1 | |||||||||
-| DE | 15 | 5e-05 | 16 | 0.1 | |||||||||
-| AD | 2 | 6e-06 | 4 | 0.1 | |||||||||
-| AL | 2 | 6e-06 | 4 | 0.1 | |||||||||
+| D-all | 15 | 5e-05 | 16 | 0.10 | |||||||||
+| DE | 15 | 5e-05 | 16 | 0.10 | |||||||||
+| AD | 2 | 6e-06 | 4 | 0.10 | |||||||||
+| AL | 2 | 6e-06 | 4 | 0.10 | |||||||||
 | AC | 4 | 1e-04 | 32 | 0.01 | |||||||||
@@
 | D-all | 13 | 1e-04 | 8 | 1e-01 | 32 | 5e-02 | 8 |
+| GPU type | NVIDIA V100 32GB | NVIDIA A100 80 GB | NVIDIA H100 80 GB | Total |
+| GPU Hours | 119 | 371 | 43 | 533 |
+| Carbon footprint, kg CO2 | 11.50 | 29.87 | 3.46 | 44.83 |
 Appendix B Full Results for Models
-Table 8 shows results for TF-IDF, models on linguistic features, and encoder-based transformers.
-Tables 9, 11 and 10 contain results for LLMs with PEFT, 0-shot and 5-shot evaluation, and 0-shot and 5-shot MMLU-style evaluation correspondingly.
+Table˜10 shows results for TF-IDF, models on linguistic features, and encoder-based transformers.
+Tables˜11, 13 and 12 contain results for LLMs with PEFT, 0-shot and 5-shot evaluation, and 0-shot and 5-shot MMLU-style evaluation correspondingly.
 | Corpus | Model |
@@
 F1-pathology | F1-macro | ||||||||
-| DE | RuRoBERTa | 90.444.62 | 95.932.28 | 92.981.92 | 64.729.29 | 56.8225.81 | 60.3727.21 | 76.6714.52 | ||||||||
-| RuBioRoBERTa | 91.925.21 | 95.193.05 | 93.362.13 | 64.2529.33 | 63.6428.63 | 63.7328.62 | 78.5415.3 | |||||||||
-| RuBERT | 94.451.16 | 91.111.7 | 92.741.04 | 68.414.04 | 78.034.85 | 72.83.43 | 82.772.21 | |||||||||
-| BERT | 92.970.8 | 92.961.23 | 92.960.85 | 71.343.96 | 71.213.39 | 71.233.14 | 82.091.98 | |||||||||
-| Linguistic features | 94.00.8 | 95.21.0 | 94.60.8 | 79.34.0 | 75.03.5 | 77.03.3 | 85.82.1 | |||||||||
-| TF-IDF | 91.62.1 | 94.42.2 | 93.01.2 | 74.87.0 | 64.410.3 | 68.56.6 | 80.73.8 | |||||||||
-| DSM | RuRoBERTa | 61.314.38 | 73.4612.73 | 66.587.66 | 46.7610.29 | 31.486.93 | 36.686.2 | 51.635.8 | ||||||||
-| RuBioRoBERTa | 62.12.98 | 62.966.05 | 62.464.3 | 43.664.43 | 42.594.14 | 43.013.7 | 52.743.67 | |||||||||
-| RuBERT | 60.781.75 | 96.916.9 | 74.521.07 | 9.0920.33 | 5.5612.42 | 6.915.42 | 40.717.18 | |||||||||
-| BERT | 61.091.9 | 69.1414.61 | 64.135.24 | 34.5915.71 | 33.3316.97 | 33.7116.01 | 48.925.62 | |||||||||
-| Linguistic features | 62.81.8 | 59.37.4 | 60.74.0 | 43.72.7 | 47.27.7 | 45.13.8 | 52.92.0 | |||||||||
-| TF-IDF | 62.22.6 | 51.213.9 | 55.38.5 | 42.93.6 | 53.711.4 | 46.94.7 | 51.13.5 | |||||||||
-| AL | RuRoBERTa | 56.825.65 | 75.07.76 | 64.495.59 | 52.5410.76 | 33.3312.77 | 40.1712.48 | 52.338.31 | ||||||||
+| DE | RuRoBERTa | 90.44 4.62 | 95.93 2.28 | 92.98 1.92 | 64.70 29.29 | 56.82 25.81 | 60.37 27.21 | 76.67 14.52 | ||||||||
+| RuBioRoBERTa | 91.92 5.21 | 95.19 3.05 | 93.36 2.13 | 64.25 29.33 | 63.64 28.63 | 63.73 28.62 | 78.54 15.30 | |||||||||
+| RuBERT | 94.45 1.16 | 91.11 1.70 | 92.74 1.04 | 68.41 4.04 | 78.03 4.85 | 72.80 3.43 | 82.77 2.21 | |||||||||
+| BERT | 92.97 0.80 | 92.96 1.23 | 92.96 0.85 | 71.34 3.96 | 71.21 3.39 | 71.23 3.14 | 82.09 1.98 | |||||||||
+| Linguistic features | 94.00 0.80 | 95.20 1.00 | 94.60 0.80 | 79.30 4.00 | 75.00 3.50 | 77.00 3.30 | 85.80 2.10 | |||||||||
+| TF-IDF | 91.60 2.10 | 94.40 2.20 | 93.00 1.20 | 74.80 7.00 | 64.40 10.30 | 68.50 6.60 | 80.70 3.80 | |||||||||
+| DSM | RuRoBERTa | 61.31 4.38 | 73.46 12.73 | 66.58 7.66 | 46.76 10.29 | 31.48 6.93 | 36.68 6.20 | 51.63 5.80 | ||||||||
+| RuBioRoBERTa | 62.10 2.98 | 62.96 6.05 | 62.46 4.30 | 43.66 4.43 | 42.59 4.14 | 43.01 3.70 | 52.74 3.67 | |||||||||
+| RuBERT | 60.78 1.75 | 96.91 6.90 | 74.52 1.07 | 9.09 20.33 | 5.56 12.42 | 6.90 15.42 | 40.71 7.18 | |||||||||
+| BERT | 61.09 1.90 | 69.14 14.61 | 64.13 5.24 | 34.59 15.71 | 33.33 16.97 | 33.71 16.01 | 48.92 5.62 | |||||||||
+| Linguistic features | 62.80 1.80 | 59.30 7.40 | 60.70 4.00 | 43.70 2.70 | 47.20 7.70 | 45.10 3.80 | 52.90 2.00 | |||||||||
+| TF-IDF | 62.20 2.60 | 51.20 13.90 | 55.30 8.50 | 42.90 3.60 | 53.70 11.40 | 46.90 4.70 | 51.10 3.50 | |||||||||
+| AL | RuRoBERTa | 56.82 5.65 | 75.00 7.76 | 64.49 5.59 | 52.54 10.76 | 33.33 12.77 | 40.17 12.48 | 52.33 8.31 | ||||||||
 | RuBioRoBERTa | 53.62 1.09 | 78.79 15.67 | 63.22 5.04 | 30.66 21.79 | 21.05 15.79 | 24.76 18.06 | 43.99 6.71 | |||||||||
-| RuBERT | 53.633.2 | 77.2719.99 | 62.087.52 | 45.8329.56 | 21.9320.48 | 24.6517.57 | 43.365.79 | |||||||||
-| BERT | 52.711.02 | 93.948.16 | 67.442.96 | 6.2513.98 | 2.635.88 | 3.78.28 | 35.572.99 | |||||||||
-| Linguistic features | 47.421.4 | 36.417.8 | 40.919.2 | 48.52.7 | 68.414.6 | 56.13.7 | 48.58.2 | |||||||||
-| TF-IDF | 61.55.0 | 46.23.1 | 52.62.3 | 51.22.5 | 65.87.9 | 57.54.5 | 55.02.9 | |||||||||
-| AD | RuRoBERTa | 57.632.52 | 56.674.71 | 57.073.19 | 52.82.82 | 53.74.14 | 53.172.84 | 55.122.58 | ||||||||
-| RuBioRoBERTa | 45.065.52 | 43.3319.72 | 41.4913.8 | 38.298.58 | 40.7423.28 | 37.7212.46 | 39.65.38 | |||||||||
-| RuBERT | 48.637.83 | 80.029.58 | 59.3415.36 | 12.5918.14 | 12.9618.33 | 12.718.04 | 36.024.94 | |||||||||
-| BERT | 52.031.0 | 93.337.99 | 66.742.88 | 20.5621.12 | 4.634.99 | 7.347.71 | 37.042.63 | |||||||||
-| Linguistic features | 50.83.3 | 43.312.5 | 45.38.1 | 44.72.6 | 51.915.9 | 47.37.4 | 46.31.8 | |||||||||
-| TF-IDF | 54.915.1 | 37.54.8 | 43.22.5 | 44.67.8 | 59.321.0 | 50.412.7 | 46.87.1 | |||||||||
-| AC | RuRoBERTa | 45.720.48 | 77.0435.48 | 57.1925.66 | 24.5724.96 | 24.5635.49 | 21.2223.82 | 39.27.55 | ||||||||
-| RuBioRoBERTa | 52.921.41 | 76.314.15 | 62.095.54 | 34.8815.72 | 20.1811.64 | 24.8412.88 | 43.474.21 | |||||||||
-| RuBERT | 54.291.57 | 87.0419.35 | 66.08.24 | 43.5922.45 | 14.0417.1 | 17.1515.7 | 41.584.88 | |||||||||
+| RuBERT | 53.63 3.20 | 77.27 19.99 | 62.08 7.52 | 45.83 29.56 | 21.93 20.48 | 24.65 17.57 | 43.36 5.79 | |||||||||
+| BERT | 52.71 1.02 | 93.94 8.16 | 67.44 2.96 | 6.25 13.98 | 2.63 5.88 | 3.70 8.28 | 35.57 2.99 | |||||||||
+| Linguistic features | 47.40 21.40 | 36.40 17.80 | 40.90 19.20 | 48.50 2.70 | 68.40 14.60 | 56.10 3.70 | 48.50 8.20 | |||||||||
+| TF-IDF | 61.50 5.00 | 46.20 3.10 | 52.60 2.30 | 51.20 2.50 | 65.80 7.90 | 57.50 4.50 | 55.00 2.90 | |||||||||
+| AD | RuRoBERTa | 57.63 2.52 | 56.67 4.71 | 57.07 3.19 | 52.80 2.82 | 53.70 4.14 | 53.17 2.84 | 55.12 2.58 | ||||||||
+| RuBioRoBERTa | 45.06 5.52 | 43.33 19.72 | 41.49 13.80 | 38.29 8.58 | 40.74 23.28 | 37.72 12.46 | 39.60 5.38 | |||||||||
+| RuBERT | 48.63 7.83 | 80.00 29.58 | 59.34 15.36 | 12.59 18.14 | 12.96 18.33 | 12.70 18.04 | 36.02 4.94 | |||||||||
+| BERT | 52.03 1.00 | 93.33 7.99 | 66.74 2.88 | 20.56 21.12 | 4.63 4.99 | 7.34 7.71 | 37.04 2.63 | |||||||||
+| Linguistic features | 50.80 3.30 | 43.30 12.50 | 45.30 8.10 | 44.70 2.60 | 51.90 15.90 | 47.30 7.40 | 46.30 1.80 | |||||||||
+| TF-IDF | 54.90 15.10 | 37.50 4.80 | 43.20 2.50 | 44.60 7.80 | 59.30 21.00 | 50.40 12.70 | 46.80 7.10 | |||||||||
+| AC | RuRoBERTa | 45.70 20.48 | 77.04 35.48 | 57.19 25.66 | 24.57 24.96 | 24.56 35.49 | 21.22 23.82 | 39.20 7.55 | ||||||||
+| RuBioRoBERTa | 52.92 1.41 | 76.30 14.15 | 62.09 5.54 | 34.88 15.72 | 20.18 11.64 | 24.84 12.88 | 43.47 4.21 | |||||||||
+| RuBERT | 54.29 1.57 | 87.04 19.35 | 66.00 8.24 | 43.59 22.45 | 14.04 17.10 | 17.15 15.70 | 41.58 4.88 | |||||||||
 | BERT | 57.01 5.15 | 62.96 16.81 | 58.44 6.53 | 46.07 8.77 | 41.67 22.14 | 41.28 17.15 | 49.86 6.84 | |||||||||
-| Linguistic features | 64.816.0 | 45.621.0 | 47.019.6 | 48.73.6 | 60.520.1 | 52.67.3 | 49.87.8 | |||||||||
-| TF-IDF | 56.32.1 | 63.34.2 | 59.52.5 | 48.93.3 | 41.75.6 | 44.94.3 | 52.22.7 |
+| Linguistic features | 64.80 16.00 | 45.60 21.00 | 47.00 19.60 | 48.70 3.60 | 60.50 20.10 | 52.60 7.30 | 49.80 7.80 | |||||||||
+| TF-IDF | 56.30 2.10 | 63.30 4.20 | 59.50 2.50 | 48.90 3.30 | 41.70 5.60 | 44.90 4.30 | 52.20 2.70 |
 | Corpus | Model |
@@
 F1-pathology | F1-macro | ||||||||
-| DE | Gemma2 2B IT | 97.221.4 | 92.961.39 | 95.020.48 | 69.76.78 | 86.634.79 | 76.843.12 | 85.931.75 | ||||||||
-| Gemma2 9B IT | 96.481.0 | 91.90.9 | 94.130.58 | 65.154.29 | 82.113.76 | 72.532.91 | 83.331.72 | |||||||||
-| Qwen2 7B IT | 95.192.1 | 90.180.34 | 92.60.96 | 57.582.14 | 75.467.33 | 65.052.19 | 78.831.56 | |||||||||
-| SaigaLlama3 8B | 98.71.19 | 91.590.79 | 95.010.41 | 62.884.08 | 92.885.97 | 74.742.12 | 84.871.22 | |||||||||
-| Vikhr 7B IT 0.4 | 94.072.77 | 90.261.15 | 92.11.11 | 58.336.11 | 72.8212.43 | 63.923.46 | 78.012.17 | |||||||||
-| Vikhr 7B IT 5.4 | 95.742.52 | 90.82.36 | 93.151.15 | 59.8512.14 | 79.079.93 | 66.918.35 | 80.034.65 | |||||||||
-| VikhrGemma 2B IT | 96.481.62 | 94.740.88 | 95.590.66 | 78.034.08 | 84.965.72 | 81.132.42 | 88.361.52 | |||||||||
-| DSM | Gemma2 2B IT | 58.649.18 | 63.294.98 | 60.626.33 | 49.079.31 | 44.425.41 | 46.246.49 | 53.435.33 | ||||||||
-| Gemma2 9B IT | 56.796.65 | 64.23.12 | 60.164.97 | 52.784.24 | 45.124.14 | 48.543.57 | 54.353.9 | |||||||||
-| Qwen2 7B IT | 72.228.21 | 68.876.39 | 70.46.75 | 50.9310.84 | 55.3311.08 | 52.8110.26 | 61.618.32 | |||||||||
-| SaigaLlama3 8B | 56.1715.48 | 57.467.16 | 56.1310.59 | 38.8913.98 | 37.439.05 | 37.2110.11 | 46.677.7 | |||||||||
-| Vikhr 7B IT 0.4 | 70.9911.4 | 64.122.52 | 67.065.9 | 40.747.64 | 50.7511.53 | 44.074.72 | 55.574.06 | |||||||||
-| Vikhr 7B IT 5.4 | 54.328.73 | 55.55.73 | 54.796.93 | 35.197.64 | 34.37.65 | 34.587.06 | 44.696.59 | |||||||||
-| VikhrGemma 2B IT | 58.6425.16 | 61.416.24 | 57.6718.04 | 49.0716.79 | 48.6612.71 | 45.686.56 | 51.688.61 | |||||||||
-| AL | Gemma2 2B IT | 60.6113.55 | 60.184.51 | 59.668.17 | 53.5110.71 | 54.877.3 | 53.456.06 | 56.565.4 | ||||||||
-| Gemma2 9B IT | 46.214.08 | 47.525.48 | 46.794.42 | 40.358.95 | 39.06.04 | 39.67.34 | 43.25.65 | |||||||||
-| Qwen2 7B IT | 50.013.38 | 53.2510.65 | 51.411.9 | 50.011.67 | 46.7411.99 | 48.1611.5 | 49.7811.33 | |||||||||
-| SaigaLlama3 8B | 53.036.25 | 56.453.35 | 54.323.1 | 51.7511.13 | 48.334.02 | 49.677.13 | 51.993.55 | |||||||||
-| Vikhr 7B IT 0.4 | 50.06.94 | 54.875.68 | 52.044.93 | 51.7510.27 | 47.014.37 | 49.016.6 | 50.534.71 | |||||||||
-| Vikhr 7B IT 5.4 | 56.068.16 | 57.333.3 | 56.474.96 | 51.756.39 | 50.764.71 | 50.993.89 | 53.733.56 | |||||||||
-| VikhrGemma 2B IT | 56.828.6 | 59.32.53 | 57.594.7 | 54.399.92 | 52.13.11 | 52.815.51 | 55.22.95 | |||||||||
-| AD | Gemma2 2B IT | 55.8310.17 | 59.215.93 | 56.626.27 | 55.5615.38 | 52.585.69 | 53.369.71 | 54.995.86 | ||||||||
-| Gemma2 9B IT | 46.675.53 | 52.034.74 | 49.054.48 | 51.858.28 | 46.54.47 | 48.915.93 | 48.984.5 | |||||||||
-| Qwen2 7B IT | 49.1712.72 | 48.12.19 | 47.917.44 | 41.6713.13 | 41.984.13 | 41.087.91 | 44.52.76 | |||||||||
-| SaigaLlama3 8B | 53.334.71 | 57.846.15 | 55.223.94 | 55.5612.42 | 51.135.98 | 53.018.89 | 54.125.88 | |||||||||
-| Vikhr 7B IT 0.4 | 51.6712.13 | 59.814.42 | 54.667.59 | 61.1111.56 | 53.474.23 | 56.436.19 | 55.554.55 | |||||||||
-| Vikhr 7B IT 5.4 | 55.012.58 | 51.918.89 | 53.3210.58 | 44.448.49 | 47.6610.86 | 45.889.46 | 49.69.75 | |||||||||
-| VikhrGemma 2B IT | 39.176.72 | 45.934.26 | 42.135.5 | 49.075.93 | 42.093.57 | 45.224.01 | 43.683.91 | |||||||||
-| AC | Gemma2 2B IT | 63.74.19 | 58.733.32 | 61.083.43 | 46.935.35 | 52.174.36 | 49.354.71 | 55.223.84 | ||||||||
-| Gemma2 9B IT | 61.485.97 | 56.841.85 | 58.852.17 | 44.38.78 | 49.072.0 | 46.215.61 | 52.532.19 | |||||||||
-| Qwen2 7B IT | 57.045.97 | 57.363.77 | 57.154.76 | 50.04.02 | 49.734.74 | 49.814.08 | 53.484.25 | |||||||||
-| SaigaLlama3 8B | 65.565.7 | 59.023.26 | 62.033.81 | 46.055.84 | 53.164.52 | 49.194.61 | 55.613.74 | |||||||||
-| Vikhr 7B IT 0.4 | 60.378.36 | 59.623.49 | 59.785.6 | 51.756.56 | 52.724.82 | 52.04.61 | 55.894.23 | |||||||||
-| Vikhr 7B IT 5.4 | 62.968.18 | 61.872.07 | 62.134.3 | 53.957.25 | 55.53.03 | 54.333.58 | 58.232.38 | |||||||||
-| VikhrGemma 2B IT | 64.816.21 | 57.772.9 | 60.983.77 | 43.866.39 | 51.44.47 | 47.114.91 | 54.053.57 |
+| DE | Gemma2 2B IT | 92.96 1.39 | 97.22 1.40 | 95.02 0.48 | 86.63 4.79 | 69.70 6.78 | 76.84 3.12 | 85.93 1.75 | ||||||||
+| Gemma2 9B IT | 91.90 0.90 | 96.48 1.00 | 94.13 0.58 | 82.11 3.76 | 65.15 4.29 | 72.53 2.91 | 83.33 1.72 | |||||||||
+| Qwen2 7B IT | 90.18 0.34 | 95.19 2.10 | 92.60 0.96 | 75.46 7.33 | 57.58 2.14 | 65.05 2.19 | 78.83 1.56 | |||||||||
+| SaigaLlama3 8B | 91.59 0.79 | 98.70 1.19 | 95.01 0.41 | 92.88 5.97 | 62.88 4.08 | 74.74 2.12 | 84.87 1.22 | |||||||||
+| Vikhr 7B IT 0.4 | 90.26 1.15 | 94.07 2.77 | 92.10 1.11 | 72.82 12.43 | 58.33 6.11 | 63.92 3.46 | 78.01 2.17 | |||||||||
+| Vikhr 7B IT 5.4 | 90.80 2.36 | 95.74 2.52 | 93.15 1.15 | 79.07 9.93 | 59.85 12.14 | 66.91 8.35 | 80.03 4.65 | |||||||||
+| VikhrGemma 2B IT | 94.74 0.88 | 96.48 1.62 | 95.59 0.66 | 84.96 5.72 | 78.03 4.08 | 81.13 2.42 | 88.36 1.52 | |||||||||
+| DSM | Gemma2 2B IT | 63.29 4.98 | 58.64 9.18 | 60.62 6.33 | 44.42 5.41 | 49.07 9.31 | 46.24 6.49 | 53.43 5.33 | ||||||||
+| Gemma2 9B IT | 64.20 3.12 | 56.79 6.65 | 60.16 4.97 | 45.12 4.14 | 52.78 4.24 | 48.54 3.57 | 54.35 3.90 | |||||||||
+| Qwen2 7B IT | 68.87 6.39 | 72.22 8.21 | 70.40 6.75 | 55.33 11.08 | 50.93 10.84 | 52.81 10.26 | 61.61 8.32 | |||||||||
+| SaigaLlama3 8B | 57.46 7.16 | 56.17 15.48 | 56.13 10.59 | 37.43 9.05 | 38.89 13.98 | 37.21 10.11 | 46.67 7.70 | |||||||||
+| Vikhr 7B IT 0.4 | 64.12 2.52 | 70.99 11.40 | 67.06 5.90 | 50.75 11.53 | 40.74 7.64 | 44.07 4.72 | 55.57 4.06 | |||||||||
+| Vikhr 7B IT 5.4 | 55.50 5.73 | 54.32 8.73 | 54.79 6.93 | 34.30 7.65 | 35.19 7.64 | 34.58 7.06 | 44.69 6.59 | |||||||||
+| VikhrGemma 2B IT | 61.41 6.24 | 58.64 25.16 | 57.67 18.04 | 48.66 12.71 | 49.07 16.79 | 45.68 6.56 | 51.68 8.61 | |||||||||
+| AL | Gemma2 2B IT | 60.18 4.51 | 60.61 13.55 | 59.66 8.17 | 54.87 7.30 | 53.51 10.71 | 53.45 6.06 | 56.56 5.40 | ||||||||
+| Gemma2 9B IT | 47.52 5.48 | 46.21 4.08 | 46.79 4.42 | 39.00 6.04 | 40.35 8.95 | 39.60 7.34 | 43.20 5.65 | |||||||||
+| Qwen2 7B IT | 53.25 10.65 | 50.00 13.38 | 51.40 11.90 | 46.74 11.99 | 50.00 11.67 | 48.16 11.50 | 49.78 11.33 | |||||||||
+| SaigaLlama3 8B | 56.45 3.35 | 53.03 6.25 | 54.32 3.10 | 48.33 4.02 | 51.75 11.13 | 49.67 7.13 | 51.99 3.55 | |||||||||
+| Vikhr 7B IT 0.4 | 54.87 5.68 | 50.00 6.94 | 52.04 4.93 | 47.01 4.37 | 51.75 10.27 | 49.01 6.60 | 50.53 4.71 | |||||||||
+| Vikhr 7B IT 5.4 | 57.33 3.30 | 56.06 8.16 | 56.47 4.96 | 50.76 4.71 | 51.75 6.39 | 50.99 3.89 | 53.73 3.56 | |||||||||
+| VikhrGemma 2B IT | 59.30 2.53 | 56.82 8.60 | 57.59 4.70 | 52.10 3.11 | 54.39 9.92 | 52.81 5.51 | 55.20 2.95 | |||||||||
+| AD | Gemma2 2B IT | 59.21 5.93 | 55.83 10.17 | 56.62 6.27 | 52.58 5.69 | 55.56 15.38 | 53.36 9.71 | 54.99 5.86 | ||||||||
+| Gemma2 9B IT | 52.03 4.74 | 46.67 5.53 | 49.05 4.48 | 46.50 4.47 | 51.85 8.28 | 48.91 5.93 | 48.98 4.50 | |||||||||
+| Qwen2 7B IT | 48.10 2.19 | 49.17 12.72 | 47.91 7.44 | 41.98 4.13 | 41.67 13.13 | 41.08 7.91 | 44.50 2.76 | |||||||||
+| SaigaLlama3 8B | 57.84 6.15 | 53.33 4.71 | 55.22 3.94 | 51.13 5.98 | 55.56 12.42 | 53.01 8.89 | 54.12 5.88 | |||||||||
+| Vikhr 7B IT 0.4 | 59.81 4.42 | 51.67 12.13 | 54.66 7.59 | 53.47 4.23 | 61.11 11.56 | 56.43 6.19 | 55.55 4.55 | |||||||||
+| Vikhr 7B IT 5.4 | 51.91 8.89 | 55.00 12.58 | 53.32 10.58 | 47.66 10.86 | 44.44 8.49 | 45.88 9.46 | 49.60 9.75 | |||||||||
+| VikhrGemma 2B IT | 45.93 4.26 | 39.17 6.72 | 42.13 5.50 | 42.09 3.57 | 49.07 5.93 | 45.22 4.01 | 43.68 3.91 | |||||||||
+| AC | Gemma2 2B IT | 58.73 3.32 | 63.70 4.19 | 61.08 3.43 | 52.17 4.36 | 46.93 5.35 | 49.35 4.71 | 55.22 3.84 | ||||||||
+| Gemma2 9B IT | 56.84 1.85 | 61.48 5.97 | 58.85 2.17 | 49.07 2.00 | 44.30 8.78 | 46.21 5.61 | 52.53 2.19 | |||||||||
+| Qwen2 7B IT | 57.36 3.77 | 57.04 5.97 | 57.15 4.76 | 49.73 4.74 | 50.00 4.02 | 49.81 4.08 | 53.48 4.25 | |||||||||
+| SaigaLlama3 8B | 59.02 3.26 | 65.56 5.70 | 62.03 3.81 | 53.16 4.52 | 46.05 5.84 | 49.19 4.61 | 55.61 3.74 | |||||||||
+| Vikhr 7B IT 0.4 | 59.62 3.49 | 60.37 8.36 | 59.78 5.60 | 52.72 4.82 | 51.75 6.56 | 52.00 4.61 | 55.89 4.23 | |||||||||
+| Vikhr 7B IT 5.4 | 61.87 2.07 | 62.96 8.18 | 62.13 4.30 | 55.50 3.03 | 53.95 7.25 | 54.33 3.58 | 58.23 2.38 | |||||||||
+| VikhrGemma 2B IT | 57.77 2.90 | 64.81 6.21 | 60.98 3.77 | 51.40 4.47 | 43.86 6.39 | 47.11 4.91 | 54.05 3.57 |
 | Corpus | Model | Mode |
@@
 F1-pathology | F1-macro | ||||||||
-| DE | Gemma2 2B IT | 5-shot MMLU | 13.33 | 80.0 | 22.86 | 86.36 | 19.59 | 31.93 | 27.39 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 13.33 | 100.0 | 23.53 | 100.0 | 22.0 | 36.07 | 29.8 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 73.33 | 92.96 | 81.99 | 77.27 | 41.46 | 53.97 | 67.98 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 53.33 | 92.31 | 67.61 | 81.82 | 30.0 | 43.9 | 55.75 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 25.56 | 88.46 | 39.66 | 86.36 | 22.09 | 35.19 | 37.42 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 38.89 | 92.11 | 54.69 | 86.36 | 25.68 | 39.58 | 47.14 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 18.89 | 85.0 | 30.91 | 86.36 | 20.65 | 33.33 | 32.12 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 35.56 | 88.89 | 50.79 | 81.82 | 23.68 | 36.73 | 43.76 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 14.44 | 72.22 | 24.07 | 77.27 | 18.09 | 29.31 | 26.69 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 19.64 | 32.84 | 16.42 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 82.22 | 87.06 | 84.57 | 50.0 | 40.74 | 44.9 | 64.73 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 5.56 | 71.43 | 10.31 | 90.91 | 19.05 | 31.5 | 20.9 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 92.22 | 85.57 | 88.77 | 36.36 | 53.33 | 43.24 | 66.01 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 97.78 | 80.0 | 88.0 | 0.0 | 0.0 | 0.0 | 44.0 | |||||||||
-| DSM | Gemma2 2B IT | 5-shot MMLU | 25.93 | 70.0 | 37.84 | 83.33 | 42.86 | 56.6 | 47.22 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 44.44 | 60.0 | 51.06 | 55.56 | 40.0 | 46.51 | 48.79 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 85.19 | 65.71 | 74.19 | 33.33 | 60.0 | 42.86 | 58.53 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 59.26 | 64.0 | 61.54 | 50.0 | 45.0 | 47.37 | 54.45 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 44.44 | 66.67 | 53.33 | 66.67 | 44.44 | 53.33 | 53.33 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 62.96 | 62.96 | 62.96 | 44.44 | 44.44 | 44.44 | 53.7 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 3.7 | 100.0 | 7.14 | 100.0 | 40.91 | 58.06 | 32.6 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 51.85 | 63.64 | 57.14 | 55.56 | 43.48 | 48.78 | 52.96 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 100.0 | 60.0 | 75.0 | 0.0 | 0.0 | 0.0 | 37.5 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 83.33 | 35.71 | 50.0 | 25.0 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 81.48 | 68.75 | 74.58 | 44.44 | 61.54 | 51.61 | 63.09 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 51.85 | 60.87 | 56.0 | 50.0 | 40.91 | 45.0 | 50.5 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 92.59 | 64.1 | 75.76 | 22.22 | 66.67 | 33.33 | 54.55 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 60.0 | 75.0 | 0.0 | 0.0 | 0.0 | 37.5 | |||||||||
-| AL | Gemma2 2B IT | 5-shot MMLU | 4.55 | 50.0 | 8.33 | 94.74 | 46.15 | 62.07 | 35.2 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 46.34 | 63.33 | 31.67 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 45.45 | 66.67 | 54.05 | 73.68 | 53.85 | 62.22 | 58.14 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 50.0 | 64.71 | 56.41 | 68.42 | 54.17 | 60.47 | 58.44 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 27.27 | 60.0 | 37.5 | 78.95 | 48.39 | 60.0 | 48.75 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 46.34 | 63.33 | 31.67 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 31.82 | 63.64 | 42.42 | 78.95 | 50.0 | 61.22 | 51.82 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 86.36 | 55.88 | 67.86 | 21.05 | 57.14 | 30.77 | 49.31 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 4.55 | 33.33 | 8.0 | 89.47 | 44.74 | 59.65 | 33.82 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 77.27 | 53.12 | 62.96 | 21.05 | 44.44 | 28.57 | 45.77 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 53.66 | 69.84 | 0.0 | 0.0 | 0.0 | 34.92 | |||||||||
-| AD | Gemma2 2B IT | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 10.0 | 50.0 | 16.67 | 88.89 | 47.06 | 61.54 | 39.1 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 15.0 | 50.0 | 23.08 | 83.33 | 46.88 | 60.0 | 41.54 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 35.0 | 77.78 | 48.28 | 88.89 | 55.17 | 68.09 | 58.18 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 70.0 | 73.68 | 71.79 | 72.22 | 68.42 | 70.27 | 71.03 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 20.0 | 80.0 | 32.0 | 94.44 | 51.52 | 66.67 | 49.33 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 85.0 | 53.12 | 65.38 | 16.67 | 50.0 | 25.0 | 45.19 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 25.0 | 83.33 | 38.46 | 94.44 | 53.12 | 68.0 | 53.23 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 10.0 | 33.33 | 15.38 | 77.78 | 43.75 | 56.0 | 35.69 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 100.0 | 52.63 | 68.97 | 0.0 | 0.0 | 0.0 | 34.48 | |||||||||
-| AC | Gemma2 2B IT | 5-shot MMLU | 17.78 | 72.73 | 28.57 | 92.11 | 48.61 | 63.64 | 46.1 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 17.78 | 72.73 | 28.57 | 92.11 | 48.61 | 63.64 | 46.1 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 88.89 | 57.14 | 69.57 | 21.05 | 61.54 | 31.37 | 50.47 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 35.56 | 61.54 | 45.07 | 73.68 | 49.12 | 58.95 | 52.01 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 31.11 | 66.67 | 42.42 | 81.58 | 50.0 | 62.0 | 52.21 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 48.89 | 59.46 | 53.66 | 60.53 | 50.0 | 54.76 | 54.21 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 13.33 | 100.0 | 23.53 | 100.0 | 49.35 | 66.09 | 44.81 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 42.22 | 70.37 | 52.78 | 78.95 | 53.57 | 63.83 | 58.3 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 100.0 | 54.22 | 70.31 | 0.0 | 0.0 | 0.0 | 35.16 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 40.0 | 64.29 | 49.32 | 73.68 | 50.91 | 60.22 | 54.77 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 17.78 | 53.33 | 26.67 | 81.58 | 45.59 | 58.49 | 42.58 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 51.11 | 57.5 | 54.12 | 55.26 | 48.84 | 51.85 | 52.98 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 97.78 | 53.66 | 69.29 | 0.0 | 0.0 | 0.0 | 34.65 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 97.78 | 53.66 | 69.29 | 0.0 | 0.0 | 0.0 | 34.65 |
+| DE | Gemma2 2B IT | 5-shot MMLU | 80.00 0.00 | 13.33 0.00 | 22.86 0.00 | 19.59 0.00 | 86.36 0.00 | 31.93 0.00 | 27.39 0.00 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 22.00 0.00 | 100.00 0.00 | 36.07 0.00 | 29.80 0.00 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 93.11 0.08 | 75.11 0.89 | 83.15 0.58 | 43.16 0.85 | 77.27 0.00 | 55.38 0.71 | 69.26 0.64 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 92.31 0.00 | 53.33 0.00 | 67.61 0.00 | 30.00 0.00 | 81.82 0.00 | 43.90 0.00 | 55.75 0.00 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 88.80 0.17 | 26.44 0.44 | 40.75 0.55 | 22.30 0.10 | 86.36 0.00 | 35.45 0.13 | 38.10 0.34 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 92.42 0.16 | 40.67 0.89 | 56.48 0.89 | 26.25 0.29 | 86.36 0.00 | 40.26 0.34 | 48.37 0.62 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 85.00 0.00 | 18.89 0.00 | 30.91 0.00 | 20.65 0.00 | 86.36 0.00 | 33.33 0.00 | 32.12 0.00 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 88.37 0.26 | 33.78 0.89 | 48.87 0.96 | 23.20 0.24 | 81.82 0.00 | 36.15 0.29 | 42.51 0.63 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 74.44 1.11 | 16.22 0.89 | 26.63 1.28 | 18.40 0.16 | 77.27 0.00 | 29.72 0.21 | 28.18 0.74 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 19.64 0.00 | 100.00 0.00 | 32.84 0.00 | 16.42 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 71.43 0.00 | 5.56 0.00 | 10.31 0.00 | 19.05 0.00 | 90.91 0.00 | 31.50 0.00 | 20.90 0.00 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 85.45 0.06 | 91.33 0.44 | 88.29 0.24 | 50.67 1.33 | 36.36 0.00 | 42.33 0.46 | 65.31 0.35 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 80.14 0.07 | 98.67 0.44 | 88.45 0.22 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 44.22 0.11 | |||||||||
+| DSM | Gemma2 2B IT | 5-shot MMLU | 71.14 0.57 | 20.00 2.96 | 31.10 3.37 | 42.26 0.30 | 87.78 2.22 | 57.04 0.22 | 44.07 1.58 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 60.00 0.00 | 44.44 0.00 | 51.06 0.00 | 40.00 0.00 | 55.56 0.00 | 46.51 0.00 | 48.79 0.00 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 64.91 0.40 | 82.22 1.48 | 72.54 0.82 | 55.64 2.18 | 33.33 0.00 | 41.67 0.59 | 57.11 0.71 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 64.00 0.00 | 59.26 0.00 | 61.54 0.00 | 45.00 0.00 | 50.00 0.00 | 47.37 0.00 | 54.45 0.00 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 66.67 0.00 | 44.44 0.00 | 53.33 0.00 | 44.44 0.00 | 66.67 0.00 | 53.33 0.00 | 53.33 0.00 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 62.96 0.00 | 62.96 0.00 | 62.96 0.00 | 44.44 0.00 | 44.44 0.00 | 44.44 0.00 | 53.70 0.00 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 63.64 0.00 | 51.85 0.00 | 57.14 0.00 | 43.48 0.00 | 55.56 0.00 | 48.78 0.00 | 52.96 0.00 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 60.00 0.00 | 100.00 0.00 | 75.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 37.50 0.00 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 35.71 0.00 | 83.33 0.00 | 50.00 0.00 | 25.00 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 68.75 0.00 | 81.48 0.00 | 74.58 0.00 | 61.54 0.00 | 44.44 0.00 | 51.61 0.00 | 63.09 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 60.87 0.00 | 51.85 0.00 | 56.00 0.00 | 40.91 0.00 | 50.00 0.00 | 45.00 0.00 | 50.50 0.00 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 62.34 0.88 | 95.56 1.48 | 75.44 0.16 | 66.67 0.00 | 13.33 4.44 | 21.90 5.71 | 48.67 2.94 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 60.00 0.00 | 100.00 0.00 | 75.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 37.50 0.00 | |||||||||
+| AL | Gemma2 2B IT | 5-shot MMLU | 50.00 0.00 | 4.55 0.00 | 8.33 0.00 | 46.15 0.00 | 94.74 0.00 | 62.07 0.00 | 35.20 0.00 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 46.34 0.00 | 100.00 0.00 | 63.33 0.00 | 31.67 0.00 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 66.67 0.00 | 45.45 0.00 | 54.05 0.00 | 53.85 0.00 | 73.68 0.00 | 62.22 0.00 | 58.14 0.00 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 64.71 0.00 | 50.00 0.00 | 56.41 0.00 | 54.17 0.00 | 68.42 0.00 | 60.47 0.00 | 58.44 0.00 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 60.00 0.00 | 27.27 0.00 | 37.50 0.00 | 48.39 0.00 | 78.95 0.00 | 60.00 0.00 | 48.75 0.00 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 61.03 0.52 | 85.45 1.82 | 71.20 0.99 | 68.73 2.55 | 36.84 0.00 | 47.95 0.64 | 59.58 0.82 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 46.34 0.00 | 100.00 0.00 | 63.33 0.00 | 31.67 0.00 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 63.64 0.00 | 31.82 0.00 | 42.42 0.00 | 50.00 0.00 | 78.95 0.00 | 61.22 0.00 | 51.82 0.00 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 55.88 0.00 | 86.36 0.00 | 67.86 0.00 | 57.14 0.00 | 21.05 0.00 | 30.77 0.00 | 49.31 0.00 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 45.00 0.00 | 94.74 0.00 | 61.02 0.00 | 30.51 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 60.00 0.00 | 81.82 0.00 | 69.23 0.00 | 63.64 0.00 | 36.84 0.00 | 46.67 0.00 | 57.95 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 33.33 0.00 | 4.55 0.00 | 8.00 0.00 | 44.74 0.00 | 89.47 0.00 | 59.65 0.00 | 33.82 0.00 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 51.92 0.60 | 73.64 1.82 | 60.89 1.03 | 40.89 1.78 | 21.05 0.00 | 27.78 0.39 | 44.34 0.71 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 53.66 0.00 | 100.00 0.00 | 69.84 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.92 0.00 | |||||||||
+| AD | Gemma2 2B IT | 5-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 47.37 0.00 | 100.00 0.00 | 64.29 0.00 | 32.14 0.00 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 50.00 0.00 | 10.00 0.00 | 16.67 0.00 | 47.06 0.00 | 88.89 0.00 | 61.54 0.00 | 39.10 0.00 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 55.71 2.86 | 19.00 2.00 | 28.32 2.62 | 48.08 0.60 | 83.33 0.00 | 60.98 0.49 | 44.65 1.56 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 77.78 0.00 | 35.00 0.00 | 48.28 0.00 | 55.17 0.00 | 88.89 0.00 | 68.09 0.00 | 58.18 0.00 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 70.74 1.47 | 70.00 0.00 | 70.36 0.72 | 67.02 0.70 | 67.78 2.22 | 67.39 1.44 | 68.87 1.08 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 47.37 0.00 | 100.00 0.00 | 64.29 0.00 | 32.14 0.00 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 80.00 0.00 | 20.00 0.00 | 32.00 0.00 | 51.52 0.00 | 94.44 0.00 | 66.67 0.00 | 49.33 0.00 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 55.62 1.25 | 89.00 2.00 | 68.46 1.54 | 63.33 6.67 | 21.11 2.22 | 31.67 3.33 | 50.06 2.44 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 13.33 26.67 | 2.00 4.00 | 3.48 6.96 | 46.47 1.05 | 94.44 0.00 | 62.28 0.93 | 32.88 3.94 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 66.67 0.00 | 10.00 0.00 | 17.39 0.00 | 48.57 0.00 | 94.44 0.00 | 64.15 0.00 | 40.77 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 38.67 2.67 | 10.00 0.00 | 15.88 0.25 | 45.11 0.68 | 82.22 2.22 | 58.26 1.13 | 37.07 0.69 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 52.63 0.00 | 100.00 0.00 | 68.97 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.48 0.00 | |||||||||
+| AC | Gemma2 2B IT | 5-shot MMLU | 74.55 0.91 | 19.56 0.89 | 30.98 1.20 | 49.16 0.27 | 92.11 0.00 | 64.10 0.23 | 47.54 0.72 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 72.73 0.00 | 17.78 0.00 | 28.57 0.00 | 48.61 0.00 | 92.11 0.00 | 63.64 0.00 | 46.10 0.00 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 60.46 1.66 | 51.56 18.67 | 53.91 7.83 | 52.31 4.62 | 58.95 18.95 | 52.50 10.56 | 53.20 1.37 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 58.46 1.54 | 33.78 0.89 | 42.82 1.13 | 47.72 0.70 | 71.58 1.05 | 57.26 0.84 | 50.04 0.98 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 71.52 2.42 | 34.67 1.78 | 46.69 2.13 | 51.97 0.98 | 83.68 1.05 | 64.12 1.06 | 55.41 1.60 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 59.89 0.22 | 52.44 1.78 | 55.91 1.12 | 50.93 0.47 | 58.42 1.05 | 54.41 0.18 | 55.16 0.47 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 49.35 0.00 | 100.00 0.00 | 66.09 0.00 | 44.81 0.00 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 67.41 1.48 | 40.44 0.89 | 50.56 1.11 | 52.14 0.71 | 76.84 1.05 | 62.13 0.85 | 56.34 0.98 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 54.22 0.00 | 100.00 0.00 | 70.31 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 35.16 0.00 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 69.15 2.43 | 41.78 0.89 | 52.09 1.39 | 53.04 1.06 | 77.89 2.11 | 63.11 1.45 | 57.60 1.42 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 90.67 18.67 | 8.89 4.44 | 15.33 5.67 | 47.12 0.76 | 96.32 7.37 | 63.22 2.37 | 39.28 1.65 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 57.50 0.00 | 51.11 0.00 | 54.12 0.00 | 48.84 0.00 | 55.26 0.00 | 51.85 0.00 | 52.98 0.00 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 55.29 0.82 | 97.78 0.00 | 70.63 0.67 | 60.00 30.00 | 6.32 3.16 | 11.43 5.71 | 41.03 3.19 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 53.66 0.00 | 97.78 0.00 | 69.29 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.65 0.00 |
 | Corpus | Model | Mode |
@@
 F1-pathology | F1-macro |
-| DE | Gemma2 2B IT | 5-shot | 12.22 | 84.62 | 21.36 | 90.91 | 20.41 | 33.33 | 18.23 | 99.11 | ||||||||||
-| Gemma2 2B IT | 0-shot | 25.56 | 95.83 | 40.35 | 95.45 | 23.86 | 38.18 | 39.27 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 55.56 | 92.59 | 69.44 | 81.82 | 31.03 | 45.0 | 57.22 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 62.22 | 93.33 | 74.67 | 81.82 | 34.62 | 48.65 | 61.66 | 100.0 | |||||||||||
-| Qwen2 7B IT | 5-shot | 25.56 | 88.46 | 39.66 | 86.36 | 22.09 | 35.19 | 37.42 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 41.11 | 90.24 | 56.49 | 81.82 | 25.35 | 38.71 | 47.6 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 18.89 | 85.0 | 30.91 | 86.36 | 20.65 | 33.33 | 32.12 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 8.89 | 100.0 | 16.33 | 100.0 | 21.15 | 34.92 | 25.62 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 13.33 | 70.59 | 22.43 | 77.27 | 17.89 | 29.06 | 25.74 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 100.0 | 19.64 | 32.84 | 16.42 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 82.22 | 87.06 | 84.57 | 50.0 | 40.74 | 44.9 | 64.73 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 5.56 | 71.43 | 10.31 | 90.91 | 19.05 | 31.5 | 20.9 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 67.78 | 88.41 | 76.73 | 63.64 | 32.56 | 43.08 | 59.9 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 5.56 | 83.33 | 10.42 | 90.91 | 19.8 | 32.52 | 14.31 | 95.54 | |||||||||||
-| DSM | Gemma2 2B IT | 5-shot | 25.93 | 70.0 | 37.84 | 77.78 | 41.18 | 53.85 | 30.56 | 97.78 | ||||||||||
-| Gemma2 2B IT | 0-shot | 40.74 | 64.71 | 50.0 | 66.67 | 42.86 | 52.17 | 51.09 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 44.44 | 66.67 | 53.33 | 66.67 | 44.44 | 53.33 | 53.33 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 77.78 | 61.76 | 68.85 | 27.78 | 45.45 | 34.48 | 51.67 | 100.0 | |||||||||||
-| Qwen2 7B IT | 5-shot | 44.44 | 70.59 | 54.55 | 72.22 | 46.43 | 56.52 | 55.53 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 70.37 | 65.52 | 67.86 | 44.44 | 50.0 | 47.06 | 57.46 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 3.7 | 100.0 | 7.14 | 100.0 | 40.91 | 58.06 | 32.6 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 25.93 | 58.33 | 35.9 | 72.22 | 39.39 | 50.98 | 43.44 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 81.48 | 70.97 | 75.86 | 50.0 | 64.29 | 56.25 | 66.06 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 83.33 | 35.71 | 50.0 | 25.0 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 81.48 | 68.75 | 74.58 | 44.44 | 61.54 | 51.61 | 63.09 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 48.15 | 59.09 | 53.06 | 50.0 | 39.13 | 43.9 | 48.48 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 37.04 | 71.43 | 48.78 | 77.78 | 45.16 | 57.14 | 52.96 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 14.81 | 57.14 | 23.53 | 72.22 | 41.94 | 53.06 | 25.53 | 84.44 | |||||||||||
-| AL | Gemma2 2B IT | 5-shot | 18.18 | 57.14 | 27.59 | 84.21 | 47.06 | 60.38 | 43.98 | 100.0 | ||||||||||
-| Gemma2 2B IT | 0-shot | 31.82 | 63.64 | 42.42 | 78.95 | 50.0 | 61.22 | 51.82 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 27.27 | 75.0 | 40.0 | 89.47 | 51.52 | 65.38 | 52.69 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 63.64 | 66.67 | 65.12 | 63.16 | 60.0 | 61.54 | 63.33 | 100.0 | |||||||||||
-| Qwen2 7B IT | 5-shot | 27.27 | 66.67 | 38.71 | 84.21 | 50.0 | 62.75 | 50.73 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 95.45 | 60.0 | 73.68 | 26.32 | 83.33 | 40.0 | 56.84 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 4.55 | 100.0 | 8.7 | 100.0 | 47.5 | 64.41 | 36.55 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 9.09 | 50.0 | 15.38 | 89.47 | 45.95 | 60.71 | 38.05 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 45.45 | 50.0 | 47.62 | 47.37 | 42.86 | 45.0 | 46.31 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 4.55 | 33.33 | 8.0 | 89.47 | 44.74 | 59.65 | 33.82 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 27.27 | 66.67 | 38.71 | 84.21 | 50.0 | 62.75 | 50.73 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | 100.0 | |||||||||||
-| AD | Gemma2 2B IT | 5-shot | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | 100.0 | ||||||||||
-| Gemma2 2B IT | 0-shot | 45.0 | 64.29 | 52.94 | 72.22 | 54.17 | 61.9 | 57.42 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 70.0 | 60.87 | 65.12 | 50.0 | 60.0 | 54.55 | 59.83 | 100.0 | |||||||||||
-| Qwen2 7B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 80.0 | 69.57 | 74.42 | 61.11 | 73.33 | 66.67 | 70.54 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 30.0 | 54.55 | 38.71 | 72.22 | 48.15 | 57.78 | 48.24 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 30.0 | 54.55 | 38.71 | 72.22 | 48.15 | 57.78 | 48.24 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 25.0 | 83.33 | 38.46 | 94.44 | 53.12 | 68.0 | 53.23 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 20.0 | 66.67 | 30.77 | 88.89 | 50.0 | 64.0 | 47.38 | 100.0 | |||||||||||
-| AC | Gemma2 2B IT | 5-shot | 22.22 | 66.67 | 33.33 | 86.84 | 49.25 | 62.86 | 32.06 | 98.8 | ||||||||||
-| Gemma2 2B IT | 0-shot | 15.56 | 87.5 | 26.42 | 97.37 | 49.33 | 65.49 | 45.95 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 15.56 | 100.0 | 26.92 | 100.0 | 50.0 | 66.67 | 46.79 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 26.67 | 54.55 | 35.82 | 73.68 | 47.46 | 57.73 | 31.18 | 97.59 | |||||||||||
-| Qwen2 7B IT | 5-shot | 15.56 | 77.78 | 25.93 | 94.74 | 48.65 | 64.29 | 45.11 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 48.89 | 56.41 | 52.38 | 55.26 | 47.73 | 51.22 | 51.8 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 13.33 | 100.0 | 23.53 | 100.0 | 49.35 | 66.09 | 44.81 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 46.67 | 67.74 | 55.26 | 73.68 | 53.85 | 62.22 | 58.74 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 88.89 | 54.79 | 67.8 | 13.16 | 50.0 | 20.83 | 44.31 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 37.78 | 65.38 | 47.89 | 76.32 | 50.88 | 61.05 | 54.47 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 17.78 | 53.33 | 26.67 | 81.58 | 45.59 | 58.49 | 42.58 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 51.11 | 57.5 | 54.12 | 55.26 | 48.84 | 51.85 | 52.98 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 88.89 | 54.79 | 67.8 | 13.16 | 50.0 | 20.83 | 44.31 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 22.22 | 62.5 | 32.79 | 73.68 | 50.0 | 59.57 | 30.79 | 86.75 |
+| DE | Gemma2 2B IT | 5-shot | 82.38 1.12 | 10.44 0.89 | 18.53 1.41 | 20.08 0.16 | 90.91 0.00 | 32.90 0.22 | 17.14 0.54 | 99.11 0.00 | ||||||||||
+| Gemma2 2B IT | 0-shot | 95.83 0.00 | 25.56 0.00 | 40.35 0.00 | 23.86 0.00 | 95.45 0.00 | 38.18 0.00 | 39.27 0.00 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 92.80 0.11 | 57.33 0.89 | 70.88 0.72 | 31.92 0.44 | 81.82 0.00 | 45.92 0.46 | 58.40 0.59 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 93.33 0.00 | 62.22 0.00 | 74.67 0.00 | 34.62 0.00 | 81.82 0.00 | 48.65 0.00 | 61.66 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 88.80 0.17 | 26.44 0.44 | 40.75 0.55 | 22.30 0.10 | 86.36 0.00 | 35.45 0.13 | 38.10 0.34 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 92.33 1.05 | 42.89 0.89 | 58.57 1.04 | 26.78 0.72 | 85.45 1.82 | 40.79 1.04 | 49.68 1.04 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 85.00 0.00 | 18.89 0.00 | 30.91 0.00 | 20.65 0.00 | 86.36 0.00 | 33.33 0.00 | 32.12 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 100.00 0.00 | 8.00 0.44 | 14.81 0.76 | 20.99 0.08 | 100.00 0.00 | 34.70 0.11 | 24.76 0.43 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 73.07 1.24 | 15.11 0.89 | 25.04 1.30 | 18.20 0.15 | 77.27 0.00 | 29.46 0.20 | 27.25 0.75 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 19.64 0.00 | 100.00 0.00 | 32.84 0.00 | 16.42 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 71.43 0.00 | 5.56 0.00 | 10.31 0.00 | 19.05 0.00 | 90.91 0.00 | 31.50 0.00 | 20.90 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 88.54 0.07 | 68.67 0.44 | 77.35 0.31 | 33.18 0.31 | 63.64 0.00 | 43.62 0.27 | 60.48 0.29 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 85.24 0.95 | 6.44 0.44 | 11.98 0.78 | 20.59 0.40 | 94.55 1.82 | 33.82 0.65 | 15.27 0.48 | 96.25 0.36 | |||||||||||
+| DSM | Gemma2 2B IT | 5-shot | 54.00 8.00 | 11.11 7.41 | 17.89 9.97 | 39.45 0.86 | 86.67 4.44 | 54.16 0.16 | 32.97 1.20 | 99.56 0.89 | ||||||||||
+| Gemma2 2B IT | 0-shot | 64.71 0.00 | 40.74 0.00 | 50.00 0.00 | 42.86 0.00 | 66.67 0.00 | 52.17 0.00 | 51.09 0.00 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 64.24 1.21 | 29.63 7.41 | 40.14 6.60 | 41.83 1.31 | 75.56 4.44 | 53.74 0.21 | 46.94 3.20 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 62.64 0.44 | 80.74 1.48 | 70.54 0.85 | 49.09 1.82 | 27.78 0.00 | 35.47 0.49 | 53.01 0.67 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 70.59 0.00 | 44.44 0.00 | 54.55 0.00 | 46.43 0.00 | 72.22 0.00 | 56.52 0.00 | 55.53 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 65.52 0.00 | 70.37 0.00 | 67.86 0.00 | 50.00 0.00 | 44.44 0.00 | 47.06 0.00 | 57.46 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 55.30 1.52 | 22.96 1.48 | 32.44 1.73 | 38.47 0.46 | 72.22 0.00 | 50.20 0.39 | 41.32 1.06 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 69.19 0.89 | 81.48 0.00 | 74.83 0.51 | 62.09 1.10 | 45.56 2.22 | 52.54 1.85 | 63.69 1.18 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 35.71 0.00 | 83.33 0.00 | 50.00 0.00 | 25.00 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 68.75 0.00 | 81.48 0.00 | 74.58 0.00 | 61.54 0.00 | 44.44 0.00 | 51.61 0.00 | 63.09 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 59.09 0.00 | 48.15 0.00 | 53.06 0.00 | 39.13 0.00 | 50.00 0.00 | 43.90 0.00 | 48.48 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 61.34 5.04 | 37.04 0.00 | 46.12 1.33 | 40.46 2.35 | 64.44 6.67 | 49.69 3.73 | 47.90 2.53 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 57.14 0.00 | 14.81 0.00 | 23.53 0.00 | 41.94 0.00 | 72.22 0.00 | 53.06 0.00 | 25.53 0.00 | 84.44 0.00 | |||||||||||
+| AL | Gemma2 2B IT | 5-shot | 57.14 0.00 | 18.18 0.00 | 27.59 0.00 | 47.06 0.00 | 84.21 0.00 | 60.38 0.00 | 43.98 0.00 | 100.00 0.00 | ||||||||||
+| Gemma2 2B IT | 0-shot | 63.64 0.00 | 31.82 0.00 | 42.42 0.00 | 50.00 0.00 | 78.95 0.00 | 61.22 0.00 | 51.82 0.00 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 75.00 0.00 | 27.27 0.00 | 40.00 0.00 | 51.52 0.00 | 89.47 0.00 | 65.38 0.00 | 52.69 0.00 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 66.67 0.00 | 63.64 0.00 | 65.12 0.00 | 60.00 0.00 | 63.16 0.00 | 61.54 0.00 | 63.33 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 61.33 2.67 | 27.27 0.00 | 37.74 0.48 | 48.71 0.65 | 80.00 2.11 | 60.55 1.10 | 49.15 0.79 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 60.00 0.00 | 95.45 0.00 | 73.68 0.00 | 83.33 0.00 | 26.32 0.00 | 40.00 0.00 | 56.84 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 4.55 0.00 | 8.70 0.00 | 47.50 0.00 | 100.00 0.00 | 64.41 0.00 | 36.55 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 36.67 6.67 | 5.45 1.82 | 9.48 2.95 | 44.98 0.48 | 89.47 0.00 | 59.86 0.43 | 34.67 1.69 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 50.00 0.00 | 45.45 0.00 | 47.62 0.00 | 42.86 0.00 | 47.37 0.00 | 45.00 0.00 | 46.31 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 45.00 0.00 | 94.74 0.00 | 61.02 0.00 | 30.51 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 60.00 0.00 | 81.82 0.00 | 69.23 0.00 | 63.64 0.00 | 36.84 0.00 | 46.67 0.00 | 57.95 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 33.33 0.00 | 4.55 0.00 | 8.00 0.00 | 44.74 0.00 | 89.47 0.00 | 59.65 0.00 | 33.82 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 66.67 0.00 | 27.27 0.00 | 38.71 0.00 | 50.00 0.00 | 84.21 0.00 | 62.75 0.00 | 50.73 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 43.87 0.56 | 90.53 2.11 | 59.10 0.96 | 29.55 0.48 | 100.00 0.00 | |||||||||||
+| AD | Gemma2 2B IT | 5-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 47.37 0.00 | 100.00 0.00 | 64.29 0.00 | 32.14 0.00 | 100.00 0.00 | ||||||||||
+| Gemma2 2B IT | 0-shot | 64.29 0.00 | 45.00 0.00 | 52.94 0.00 | 54.17 0.00 | 72.22 0.00 | 61.90 0.00 | 57.42 0.00 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 60.87 0.00 | 70.00 0.00 | 65.12 0.00 | 60.00 0.00 | 50.00 0.00 | 54.55 0.00 | 59.83 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 67.25 1.16 | 80.00 0.00 | 73.07 0.68 | 71.81 0.76 | 56.67 2.22 | 63.33 1.67 | 68.20 1.17 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 54.55 0.00 | 30.00 0.00 | 38.71 0.00 | 48.15 0.00 | 72.22 0.00 | 57.78 0.00 | 48.24 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 50.91 1.82 | 30.00 0.00 | 37.74 0.48 | 46.55 0.80 | 67.78 2.22 | 55.19 1.29 | 46.47 0.89 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 13.33 26.67 | 2.00 4.00 | 3.48 6.96 | 46.47 1.05 | 94.44 0.00 | 62.28 0.93 | 32.88 3.94 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 66.67 0.00 | 10.00 0.00 | 17.39 0.00 | 48.57 0.00 | 94.44 0.00 | 64.15 0.00 | 40.77 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 66.67 0.00 | 20.00 0.00 | 30.77 0.00 | 50.00 0.00 | 88.89 0.00 | 64.00 0.00 | 47.38 0.00 | 100.00 0.00 | |||||||||||
+| AC | Gemma2 2B IT | 5-shot | 72.00 2.67 | 24.00 0.89 | 36.00 1.33 | 50.45 0.60 | 88.95 1.05 | 64.38 0.76 | 33.46 0.70 | 98.80 0.00 | ||||||||||
+| Gemma2 2B IT | 0-shot | 97.50 5.00 | 15.56 0.00 | 26.82 0.20 | 49.87 0.27 | 99.47 1.05 | 66.43 0.47 | 46.63 0.34 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 68.00 16.00 | 13.78 0.89 | 22.84 2.04 | 47.26 1.37 | 91.58 4.21 | 62.34 2.16 | 42.59 2.10 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 54.55 0.00 | 26.67 0.00 | 35.82 0.00 | 47.46 0.00 | 73.68 0.00 | 57.73 0.00 | 31.18 0.00 | 97.59 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 84.89 3.56 | 26.22 5.33 | 39.85 6.96 | 52.08 1.72 | 94.74 0.00 | 67.20 1.46 | 53.52 4.21 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 55.93 0.24 | 52.44 1.78 | 54.11 0.87 | 47.55 0.09 | 51.05 2.11 | 49.22 1.00 | 51.67 0.07 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 49.35 0.00 | 100.00 0.00 | 66.09 0.00 | 44.81 0.00 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 66.05 0.85 | 46.67 0.00 | 54.69 0.29 | 53.12 0.36 | 71.58 1.05 | 60.98 0.62 | 57.84 0.45 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 54.20 0.30 | 88.89 0.00 | 67.34 0.23 | 45.56 2.22 | 11.05 1.05 | 17.78 1.52 | 42.56 0.88 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 67.48 1.05 | 37.78 0.00 | 48.43 0.27 | 51.55 0.34 | 78.42 1.05 | 62.21 0.58 | 55.32 0.43 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 90.67 18.67 | 8.89 4.44 | 15.33 5.67 | 47.12 0.76 | 96.32 7.37 | 63.22 2.37 | 39.28 1.65 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 57.50 0.00 | 51.11 0.00 | 54.12 0.00 | 48.84 0.00 | 55.26 0.00 | 51.85 0.00 | 52.98 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 52.75 1.02 | 80.00 4.44 | 63.56 2.12 | 40.00 5.00 | 15.26 1.05 | 21.94 0.56 | 42.75 0.78 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 64.26 0.88 | 24.00 0.89 | 34.94 1.08 | 50.00 0.00 | 73.68 0.00 | 59.57 0.00 | 31.51 0.36 | 87.71 0.48 |
 Appendix C Full Results on D-all and A-all Datasets
-Tables 12, 13 and 14 show results for all models on D-all and A-all datasets.
+Tables˜14, 15 and 16 show results for all models on D-all and A-all datasets.
 Among models with SFT Gemma2 2B IT is the best on D-all, while on A-all the best are encoder-based models – RuRoBERTa and BERT.
 While in general prompting performs worse than SFT on D-all, the best model (Gemma 2 9B IT) shows comparable performance with AutoML models.
-On the contrary, on A-all prompting shows better results than SFT.
+On the contrary, prompting on the A-all shows better results than SFT.
 The best models in various prompting settings are Qwen2 7B IT, SaigaLlama3 8B, and Gemma2 2B IT.
@@
 F1-pathology | F1-macro | ||||||||
-| D-all | Gemma2 2B IT | 97.141.04 | 92.330.94 | 94.670.72 | 52.986.33 | 76.446.36 | 62.335.6 | 78.53.13 | ||||||||
-| Gemma2 9B IT | 97.650.82 | 91.581.24 | 94.510.68 | 47.628.42 | 77.866.23 | 58.696.91 | 76.63.77 | |||||||||
-| Qwen2 7B IT | 97.030.96 | 91.090.64 | 93.960.17 | 44.644.94 | 72.975.26 | 54.982.77 | 74.471.36 | |||||||||
-| SaigaLlama3 8B | 96.731.49 | 90.881.0 | 93.711.01 | 43.456.65 | 70.1610.04 | 53.417.35 | 73.564.13 | |||||||||
-| Vikhr 7B IT 0.4 | 96.831.2 | 90.461.23 | 93.531.01 | 40.488.16 | 68.8611.28 | 50.738.61 | 72.134.79 | |||||||||
-| Vikhr 7B IT 5.4 | 92.021.28 | 90.181.16 | 91.091.16 | 41.677.04 | 47.327.9 | 44.287.35 | 67.694.24 | |||||||||
-| VikhrGemma 2B IT | 97.441.68 | 91.830.8 | 94.540.66 | 49.45.98 | 79.2511.48 | 60.043.8 | 77.292.09 | |||||||||
-| RuRoBERTa | 92.480.63 | 95.51.16 | 93.960.42 | 68.195.05 | 54.764.45 | 60.482.31 | 77.221.27 | |||||||||
-| RuBioRoBERTa | 90.742.52 | 96.931.66 | 93.691.04 | 58.526.99 | 41.6719.2 | 48.6122.33 | 71.1511.6 | |||||||||
-| RuBERT | 92.260.52 | 96.221.56 | 94.190.68 | 72.1210.1 | 52.983.81 | 60.62.87 | 77.391.72 | |||||||||
-| BERT | 90.163.31 | 96.631.69 | 93.221.43 | 50.5228.15 | 37.524.38 | 42.526.53 | 67.8613.9 | |||||||||
-| Linguistic features | 88.10.6 | 88.13.4 | 88.01.6 | 51.65.8 | 50.74.1 | 50.82.6 | 69.41.9 | |||||||||
-| TF-IDF | 89.51.0 | 84.21.0 | 86.80.6 | 47.51.8 | 59.14.2 | 52.62.5 | 69.71.5 | |||||||||
-| A-all | Gemma2 2B IT | 57.094.28 | 57.613.21 | 57.273.15 | 51.116.0 | 50.613.62 | 50.774.48 | 54.023.39 | ||||||||
-| Gemma2 9B IT | 56.324.04 | 55.441.46 | 55.832.54 | 47.563.14 | 48.471.77 | 47.951.96 | 51.891.63 | |||||||||
-| Qwen2 7B IT | 56.133.95 | 56.431.57 | 56.232.59 | 49.782.85 | 49.512.04 | 49.591.86 | 52.911.79 | |||||||||
-| SaigaLlama3 8B | 56.324.83 | 58.561.32 | 57.332.79 | 53.783.74 | 51.592.03 | 52.561.83 | 54.951.57 | |||||||||
-| Vikhr 7B IT 0.4 | 56.132.43 | 57.083.28 | 56.582.68 | 50.894.95 | 49.923.4 | 50.384.09 | 53.483.32 | |||||||||
-| Vikhr 7B IT 5.4 | 55.367.22 | 56.892.4 | 55.954.71 | 51.564.79 | 50.13.07 | 50.652.68 | 53.32.82 | |||||||||
-| VikhrGemma 2B IT | 53.262.94 | 54.792.65 | 53.972.34 | 48.894.91 | 47.342.63 | 48.063.59 | 51.012.62 | |||||||||
-| RuRoBERTa | 61.533.69 | 60.1513.8 | 59.684.47 | 54.390.97 | 54.4418.08 | 52.2711.87 | 55.973.88 | |||||||||
+| D-all | Gemma2 2B IT | 92.33 0.94 | 97.14 1.04 | 94.67 0.72 | 76.44 6.36 | 52.98 6.33 | 62.33 5.60 | 78.50 3.13 | ||||||||
+| Gemma2 9B IT | 91.58 1.24 | 97.65 0.82 | 94.51 0.68 | 77.86 6.23 | 47.62 8.42 | 58.69 6.91 | 76.60 3.77 | |||||||||
+| Qwen2 7B IT | 91.09 0.64 | 97.03 0.96 | 93.96 0.17 | 72.97 5.26 | 44.64 4.94 | 54.98 2.77 | 74.47 1.36 | |||||||||
+| SaigaLlama3 8B | 90.88 1.00 | 96.73 1.49 | 93.71 1.01 | 70.16 10.04 | 43.45 6.65 | 53.41 7.35 | 73.56 4.13 | |||||||||
+| Vikhr 7B IT 0.4 | 90.46 1.23 | 96.83 1.20 | 93.53 1.01 | 68.86 11.28 | 40.48 8.16 | 50.73 8.61 | 72.13 4.79 | |||||||||
+| Vikhr 7B IT 5.4 | 90.18 1.16 | 92.02 1.28 | 91.09 1.16 | 47.32 7.90 | 41.67 7.04 | 44.28 7.35 | 67.69 4.24 | |||||||||
+| VikhrGemma 2B IT | 91.83 0.80 | 97.44 1.68 | 94.54 0.66 | 79.25 11.48 | 49.40 5.98 | 60.04 3.80 | 77.29 2.09 | |||||||||
+| RuRoBERTa | 92.48 0.63 | 95.50 1.16 | 93.96 0.42 | 68.19 5.05 | 54.76 4.45 | 60.48 2.31 | 77.22 1.27 | |||||||||
+| RuBioRoBERTa | 90.74 2.52 | 96.93 1.66 | 93.69 1.04 | 58.50 26.99 | 41.67 19.20 | 48.61 22.33 | 71.15 11.60 | |||||||||
+| RuBERT | 92.26 0.52 | 96.22 1.56 | 94.19 0.68 | 72.12 10.10 | 52.98 3.81 | 60.60 2.87 | 77.39 1.72 | |||||||||
+| BERT | 90.16 3.31 | 96.63 1.69 | 93.22 1.43 | 50.52 28.15 | 37.50 24.38 | 42.50 26.53 | 67.86 13.90 | |||||||||
+| Linguistic features | 88.10 0.60 | 88.10 3.40 | 88.00 1.60 | 51.60 5.80 | 50.70 4.10 | 50.80 2.60 | 69.40 1.90 | |||||||||
+| TF-IDF | 89.50 1.00 | 84.20 1.00 | 86.80 0.60 | 47.50 1.80 | 59.10 4.20 | 52.60 2.50 | 69.70 1.50 | |||||||||
+| A-all | Gemma2 2B IT | 57.61 3.21 | 57.09 4.28 | 57.27 3.15 | 50.61 3.62 | 51.11 6.00 | 50.77 4.48 | 54.02 3.39 | ||||||||
+| Gemma2 9B IT | 55.44 1.46 | 56.32 4.04 | 55.83 2.54 | 48.47 1.77 | 47.56 3.14 | 47.95 1.96 | 51.89 1.63 | |||||||||
+| Qwen2 7B IT | 56.43 1.57 | 56.13 3.95 | 56.23 2.59 | 49.51 2.04 | 49.78 2.85 | 49.59 1.86 | 52.91 1.79 | |||||||||
+| SaigaLlama3 8B | 58.56 1.32 | 56.32 4.83 | 57.33 2.79 | 51.59 2.03 | 53.78 3.74 | 52.56 1.83 | 54.95 1.57 | |||||||||
+| Vikhr 7B IT 0.4 | 57.08 3.28 | 56.13 2.43 | 56.58 2.68 | 49.92 3.40 | 50.89 4.95 | 50.38 4.09 | 53.48 3.32 | |||||||||
+| Vikhr 7B IT 5.4 | 56.89 2.40 | 55.36 7.22 | 55.95 4.71 | 50.10 3.07 | 51.56 4.79 | 50.65 2.68 | 53.30 2.82 | |||||||||
+| VikhrGemma 2B IT | 54.79 2.65 | 53.26 2.94 | 53.97 2.34 | 47.34 2.63 | 48.89 4.91 | 48.06 3.59 | 51.01 2.62 | |||||||||
+| RuRoBERTa | 61.53 3.69 | 60.15 13.80 | 59.68 4.47 | 54.39 0.97 | 54.44 18.08 | 52.27 11.87 | 55.97 3.88 | |||||||||
 | RuBioRoBERTa | 55.16 1.73 | 73.18 19.96 | 61.66 6.68 | 32.94 23.36 | 30.22 22.11 | 31.31 22.37 | 46.48 8.32 | |||||||||
-| RuBERT | 57.181.93 | 51.535.61 | 54.093.67 | 49.691.98 | 55.334.34 | 52.252.28 | 53.172.1 | |||||||||
-| BERT | 59.441.74 | 46.933.66 | 52.332.1 | 50.430.98 | 62.674.81 | 55.822.21 | 54.081.08 | |||||||||
-| Linguistic features | 53.42.9 | 39.719.3 | 42.017.9 | 46.12.4 | 60.220.4 | 51.07.8 | 46.56.2 | |||||||||
-| TF-IDF | 51.80.7 | 40.015.6 | 43.310.8 | 44.32.1 | 56.417.8 | 48.68.5 | 46.02.4 |
+| RuBERT | 57.18 1.93 | 51.53 5.61 | 54.09 3.67 | 49.69 1.98 | 55.33 4.34 | 52.25 2.28 | 53.17 2.10 | |||||||||
+| BERT | 59.44 1.74 | 46.93 3.66 | 52.33 2.10 | 50.43 0.98 | 62.67 4.81 | 55.82 2.21 | 54.08 1.08 | |||||||||
+| Linguistic features | 53.40 2.90 | 39.70 19.30 | 42.00 17.90 | 46.10 2.40 | 60.20 20.40 | 51.00 7.80 | 46.50 6.20 | |||||||||
+| TF-IDF | 51.80 0.70 | 40.00 15.60 | 43.30 10.80 | 44.30 2.10 | 56.40 17.80 | 48.60 8.50 | 46.00 2.40 |
 | Corpus | Model | Mode |
@@
 F1-pathology | F1-macro | ||||||||
-| D-all | Gemma2 2B IT | 5-shot MMLU | 7.36 | 100.0 | 13.71 | 100.0 | 15.64 | 27.05 | 20.38 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 30.06 | 100.0 | 46.23 | 100.0 | 19.72 | 32.94 | 39.58 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 97.55 | 89.33 | 93.26 | 32.14 | 69.23 | 43.90 | 68.58 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 68.1 | 94.87 | 79.29 | 78.57 | 29.73 | 43.14 | 61.21 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 22.09 | 94.74 | 35.82 | 92.86 | 16.99 | 28.73 | 32.28 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 52.15 | 95.51 | 67.46 | 85.71 | 23.53 | 36.92 | 52.19 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 0.61 | 100.0 | 1.22 | 100.0 | 14.74 | 25.69 | 13.45 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 41.1 | 93.06 | 57.02 | 82.14 | 19.33 | 31.29 | 44.16 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 89.57 | 88.48 | 89.02 | 32.14 | 34.62 | 33.33 | 61.18 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 2.45 | 100.0 | 4.79 | 100.0 | 14.97 | 26.05 | 15.42 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 57.06 | 96.88 | 71.81 | 89.29 | 26.32 | 40.65 | 56.23 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 31.29 | 91.07 | 46.58 | 82.14 | 17.04 | 28.22 | 37.4 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 38.04 | 84.93 | 52.54 | 60.71 | 14.41 | 23.29 | 37.92 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 98.77 | 85.19 | 91.48 | 0.0 | 0.0 | 0.0 | 45.74 | |||||||||
-| A-all | Gemma2 2B IT | 5-shot MMLU | 10.34 | 64.29 | 17.82 | 93.33 | 47.3 | 62.78 | 40.3 | ||||||||
-| Gemma2 2B IT | 0-shot MMLU | 11.49 | 66.67 | 19.61 | 93.33 | 47.62 | 63.06 | 41.34 | |||||||||
-| Gemma2 9B IT | 5-shot MMLU | 73.56 | 57.14 | 64.32 | 36.0 | 54.0 | 43.2 | 53.76 | |||||||||
-| Gemma2 9B IT | 0-shot MMLU | 39.08 | 65.38 | 48.92 | 76.0 | 51.82 | 61.62 | 55.27 | |||||||||
-| Qwen2 7B IT | 5-shot MMLU | 56.32 | 62.82 | 59.39 | 61.33 | 54.76 | 57.86 | 58.63 | |||||||||
-| Qwen2 7B IT | 0-shot MMLU | 62.07 | 62.79 | 62.43 | 57.33 | 56.58 | 56.95 | 59.69 | |||||||||
-| SaigaLlama3 8B | 5-shot MMLU | 10.34 | 69.23 | 18.0 | 94.67 | 47.65 | 63.39 | 40.7 | |||||||||
-| SaigaLlama3 8B | 0-shot MMLU | 34.48 | 69.77 | 46.15 | 82.67 | 52.1 | 63.92 | 55.04 | |||||||||
-| Vikhr 7B IT 0.4 | 5-shot MMLU | 96.55 | 53.85 | 69.14 | 4.0 | 50.0 | 7.41 | 38.27 | |||||||||
-| Vikhr 7B IT 0.4 | 0-shot MMLU | 22.99 | 62.5 | 33.61 | 84.0 | 48.46 | 61.46 | 47.54 | |||||||||
-| Vikhr 7B IT 5.4 | 5-shot MMLU | 13.79 | 70.59 | 23.08 | 93.33 | 48.28 | 63.64 | 43.36 | |||||||||
-| Vikhr 7B IT 5.4 | 0-shot MMLU | 33.33 | 59.18 | 42.65 | 73.33 | 48.67 | 58.51 | 50.58 | |||||||||
-| VikhrGemma 2B IT | 5-shot MMLU | 94.25 | 54.67 | 69.2 | 9.33 | 58.33 | 16.09 | 42.65 | |||||||||
-| VikhrGemma 2B IT | 0-shot MMLU | 98.85 | 53.42 | 69.35 | 0.0 | 0.0 | 0.0 | 34.68 |
+| D-all | Gemma2 2B IT | 5-shot MMLU | 100.00 0.00 | 7.36 0.00 | 13.71 0.00 | 15.64 0.00 | 100.00 0.00 | 27.05 0.00 | 20.38 0.00 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 100.00 0.00 | 30.55 0.25 | 46.80 0.29 | 19.83 0.06 | 100.00 0.00 | 33.10 0.08 | 39.95 0.18 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 88.83 0.25 | 96.56 0.49 | 92.53 0.36 | 59.56 4.84 | 29.29 1.43 | 39.26 2.32 | 65.90 1.34 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 94.87 0.00 | 68.10 0.00 | 79.29 0.00 | 29.73 0.00 | 78.57 0.00 | 43.14 0.00 | 61.21 0.00 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 94.74 0.00 | 22.09 0.00 | 35.82 0.00 | 16.99 0.00 | 92.86 0.00 | 28.73 0.00 | 32.28 0.00 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 95.66 0.08 | 54.11 0.98 | 69.12 0.83 | 24.30 0.38 | 85.71 0.00 | 37.86 0.47 | 53.49 0.65 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 100.00 0.00 | 1.10 0.25 | 2.18 0.48 | 14.80 0.03 | 100.00 0.00 | 25.78 0.05 | 13.98 0.26 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 92.98 0.04 | 40.61 0.25 | 56.53 0.24 | 19.20 0.06 | 82.14 0.00 | 31.12 0.08 | 43.83 0.16 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 88.97 0.24 | 90.06 0.25 | 89.51 0.24 | 37.69 1.54 | 35.00 1.43 | 36.30 1.48 | 62.90 0.86 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 100.00 0.00 | 2.45 0.00 | 4.79 0.00 | 14.97 0.00 | 100.00 0.00 | 26.05 0.00 | 15.42 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 96.88 0.00 | 57.06 0.00 | 71.81 0.00 | 26.32 0.00 | 89.29 0.00 | 40.65 0.00 | 56.23 0.00 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 90.94 0.06 | 30.80 0.25 | 46.01 0.28 | 16.94 0.05 | 82.14 0.00 | 28.08 0.07 | 37.05 0.18 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 84.93 0.00 | 38.04 0.00 | 52.54 0.00 | 14.41 0.00 | 60.71 0.00 | 23.29 0.00 | 37.92 0.00 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 85.25 0.03 | 99.26 0.25 | 91.72 0.12 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 45.86 0.06 | |||||||||
+| A-all | Gemma2 2B IT | 5-shot MMLU | 62.86 0.71 | 6.67 1.84 | 11.99 2.92 | 46.86 0.22 | 95.47 1.07 | 62.86 0.04 | 37.42 1.44 | ||||||||
+| Gemma2 2B IT | 0-shot MMLU | 66.67 0.00 | 11.49 0.00 | 19.61 0.00 | 47.62 0.00 | 93.33 0.00 | 63.06 0.00 | 41.34 0.00 | |||||||||
+| Gemma2 9B IT | 5-shot MMLU | 57.52 0.19 | 63.45 5.06 | 60.24 2.04 | 51.94 1.03 | 45.60 4.80 | 48.36 2.58 | 54.30 0.27 | |||||||||
+| Gemma2 9B IT | 0-shot MMLU | 63.85 0.77 | 38.16 0.46 | 47.77 0.58 | 51.09 0.36 | 74.93 0.53 | 60.76 0.43 | 54.26 0.50 | |||||||||
+| Qwen2 7B IT | 5-shot MMLU | 65.42 1.30 | 45.29 5.52 | 53.28 3.06 | 53.22 0.77 | 72.00 5.33 | 61.08 1.61 | 57.18 0.73 | |||||||||
+| Qwen2 7B IT | 0-shot MMLU | 62.67 0.06 | 64.83 1.38 | 63.72 0.65 | 57.51 0.47 | 55.20 1.07 | 56.32 0.32 | 60.02 0.17 | |||||||||
+| SaigaLlama3 8B | 5-shot MMLU | 80.51 5.64 | 11.26 0.46 | 19.76 0.88 | 48.46 0.41 | 96.80 1.07 | 64.59 0.60 | 42.18 0.74 | |||||||||
+| SaigaLlama3 8B | 0-shot MMLU | 67.91 0.93 | 33.56 0.46 | 44.92 0.62 | 51.43 0.34 | 81.60 0.53 | 63.09 0.41 | 54.01 0.51 | |||||||||
+| Vikhr 7B IT 0.4 | 5-shot MMLU | 56.68 1.42 | 80.00 8.28 | 66.10 1.52 | 54.26 2.13 | 28.53 12.27 | 35.58 14.09 | 50.84 6.29 | |||||||||
+| Vikhr 7B IT 0.4 | 0-shot MMLU | 64.91 1.21 | 22.07 0.46 | 32.93 0.34 | 48.79 0.16 | 86.13 1.07 | 62.29 0.41 | 47.61 0.04 | |||||||||
+| Vikhr 7B IT 5.4 | 5-shot MMLU | 62.12 4.24 | 22.07 4.14 | 32.16 4.54 | 48.08 0.10 | 83.73 4.80 | 61.04 1.30 | 46.60 1.62 | |||||||||
+| Vikhr 7B IT 5.4 | 0-shot MMLU | 59.18 0.00 | 33.33 0.00 | 42.65 0.00 | 48.67 0.00 | 73.33 0.00 | 58.51 0.00 | 50.58 0.00 | |||||||||
+| VikhrGemma 2B IT | 5-shot MMLU | 53.67 0.50 | 97.93 1.84 | 69.32 0.06 | 11.67 23.33 | 1.87 3.73 | 3.22 6.44 | 36.27 3.19 | |||||||||
+| VikhrGemma 2B IT | 0-shot MMLU | 53.42 0.00 | 98.85 0.00 | 69.35 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 34.68 0.00 |
 | Corpus | Model | Mode |
@@
 F1-pathology | F1-macro |
-| D-all | Gemma2 2B IT | 5-shot | 19.63 | 94.12 | 32.49 | 92.86 | 16.56 | 28.11 | 30.3 | 100.0 | ||||||||||
-| Gemma2 2B IT | 0-shot | 42.33 | 97.18 | 58.97 | 92.86 | 21.67 | 35.14 | 47.05 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 31.9 | 98.11 | 48.15 | 96.43 | 19.57 | 32.53 | 40.34 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 76.69 | 93.28 | 84.18 | 67.86 | 33.33 | 44.71 | 64.44 | 100.0 | |||||||||||
-| Qwen2 7B IT | 5-shot | 21.47 | 94.59 | 35.0 | 92.86 | 16.88 | 28.57 | 31.79 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 54.6 | 94.68 | 69.26 | 82.14 | 23.71 | 36.8 | 53.03 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 5.52 | 100.0 | 10.47 | 100.0 | 15.38 | 26.67 | 18.57 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 15.95 | 96.3 | 27.37 | 96.43 | 16.46 | 28.12 | 27.75 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 39.88 | 92.86 | 55.79 | 82.14 | 19.01 | 30.87 | 43.33 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 2.45 | 100.0 | 4.79 | 100.0 | 14.97 | 26.05 | 15.42 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 57.06 | 96.88 | 71.81 | 89.29 | 26.32 | 40.65 | 56.23 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 31.29 | 91.07 | 46.58 | 82.14 | 17.04 | 28.22 | 37.4 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 18.4 | 93.75 | 30.77 | 92.86 | 16.35 | 27.81 | 29.29 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 17.18 | 96.55 | 29.17 | 92.86 | 16.88 | 28.57 | 19.25 | 95.81 | |||||||||||
-| A-all | Gemma2 2B IT | 5-shot | 20.69 | 72.0 | 32.14 | 90.67 | 49.64 | 64.15 | 48.15 | 100.0 | ||||||||||
-| Gemma2 2B IT | 0-shot | 26.44 | 69.7 | 38.33 | 86.67 | 50.39 | 63.73 | 51.03 | 100.0 | |||||||||||
-| Gemma2 9B IT | 5-shot | 45.98 | 66.67 | 54.42 | 73.33 | 53.92 | 62.15 | 58.28 | 100.0 | |||||||||||
-| Gemma2 9B IT | 0-shot | 45.98 | 60.61 | 52.29 | 65.33 | 52.13 | 57.99 | 36.76 | 98.77 | |||||||||||
-| Qwen2 7B IT | 5-shot | 37.93 | 63.46 | 47.48 | 74.67 | 50.91 | 60.54 | 54.01 | 100.0 | |||||||||||
-| Qwen2 7B IT | 0-shot | 67.82 | 60.82 | 64.13 | 49.33 | 56.92 | 52.86 | 58.49 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 5-shot | 14.94 | 68.42 | 24.53 | 92.0 | 48.25 | 63.3 | 43.92 | 100.0 | |||||||||||
-| SaigaLlama3 8B | 0-shot | 33.33 | 63.04 | 43.61 | 77.33 | 50.0 | 60.73 | 52.17 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 5-shot | 78.16 | 54.4 | 64.15 | 24.0 | 48.65 | 32.14 | 48.15 | 100.0 | |||||||||||
-| Vikhr 7B IT 0.4 | 0-shot | 21.84 | 63.33 | 32.48 | 85.33 | 48.48 | 61.84 | 47.16 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 5-shot | 13.79 | 70.59 | 23.08 | 93.33 | 48.28 | 63.64 | 43.36 | 100.0 | |||||||||||
-| Vikhr 7B IT 5.4 | 0-shot | 33.33 | 59.18 | 42.65 | 73.33 | 48.67 | 58.51 | 50.58 | 100.0 | |||||||||||
-| VikhrGemma 2B IT | 5-shot | 42.53 | 56.92 | 48.68 | 62.67 | 48.96 | 54.97 | 34.55 | 99.38 | |||||||||||
-| VikhrGemma 2B IT | 0-shot | 16.09 | 60.87 | 25.45 | 82.67 | 48.44 | 61.08 | 28.85 | 93.21 |
+| D-all | Gemma2 2B IT | 5-shot | 94.25 0.07 | 20.12 0.25 | 33.16 0.34 | 16.65 0.04 | 92.86 0.00 | 28.23 0.06 | 30.70 0.20 | 100.00 0.00 | ||||||||||
+| Gemma2 2B IT | 0-shot | 96.06 0.56 | 41.84 0.25 | 58.29 0.34 | 21.00 0.33 | 90.00 1.43 | 34.05 0.54 | 46.17 0.44 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 98.17 0.03 | 32.88 0.49 | 49.26 0.56 | 19.80 0.12 | 96.43 0.00 | 32.85 0.16 | 41.06 0.36 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 93.28 0.00 | 76.69 0.00 | 84.18 0.00 | 33.33 0.00 | 67.86 0.00 | 44.71 0.00 | 64.44 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 94.59 0.00 | 21.47 0.00 | 35.00 0.00 | 16.88 0.00 | 92.86 0.00 | 28.57 0.00 | 31.79 0.00 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 95.64 0.48 | 56.56 0.98 | 71.08 0.91 | 25.17 0.73 | 85.00 1.43 | 38.84 1.02 | 54.96 0.96 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 7.48 0.98 | 13.91 1.72 | 15.66 0.14 | 100.00 0.00 | 27.08 0.21 | 20.50 0.97 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 96.18 0.06 | 15.46 0.25 | 26.64 0.37 | 16.38 0.04 | 96.43 0.00 | 28.01 0.06 | 27.32 0.21 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 93.02 0.08 | 40.86 0.49 | 56.78 0.49 | 19.26 0.13 | 82.14 0.00 | 31.21 0.17 | 43.99 0.33 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 100.00 0.00 | 2.45 0.00 | 4.79 0.00 | 14.97 0.00 | 100.00 0.00 | 26.05 0.00 | 15.42 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 96.88 0.00 | 57.06 0.00 | 71.81 0.00 | 26.32 0.00 | 89.29 0.00 | 40.65 0.00 | 56.23 0.00 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 90.94 0.06 | 30.80 0.25 | 46.01 0.28 | 16.94 0.05 | 82.14 0.00 | 28.08 0.07 | 37.05 0.18 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 93.90 0.08 | 18.90 0.25 | 31.46 0.35 | 16.43 0.04 | 92.86 0.00 | 27.93 0.06 | 29.69 0.20 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 96.73 0.09 | 18.16 0.49 | 30.58 0.70 | 17.59 0.35 | 95.71 1.43 | 29.71 0.57 | 20.10 0.43 | 95.81 0.00 | |||||||||||
+| A-all | Gemma2 2B IT | 5-shot | 69.78 1.11 | 20.69 0.00 | 31.92 0.11 | 49.63 0.00 | 89.60 0.53 | 63.88 0.14 | 35.14 6.50 | 99.51 0.25 | ||||||||||
+| Gemma2 2B IT | 0-shot | 71.44 0.87 | 26.44 0.00 | 38.59 0.13 | 50.69 0.15 | 87.73 0.53 | 64.26 0.27 | 51.42 0.20 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 5-shot | 67.80 0.57 | 38.62 3.68 | 49.09 2.66 | 52.52 0.70 | 78.67 2.67 | 62.96 0.40 | 56.02 1.13 | 100.00 0.00 | |||||||||||
+| Gemma2 9B IT | 0-shot | 60.61 0.00 | 45.98 0.00 | 52.29 0.00 | 52.13 0.00 | 65.33 0.00 | 57.99 0.00 | 36.76 0.00 | 98.77 0.00 | |||||||||||
+| Qwen2 7B IT | 5-shot | 67.93 2.23 | 34.25 1.84 | 45.47 1.01 | 51.52 0.30 | 81.07 3.20 | 62.98 1.22 | 54.22 0.11 | 100.00 0.00 | |||||||||||
+| Qwen2 7B IT | 0-shot | 60.01 0.41 | 69.66 0.92 | 64.47 0.17 | 56.72 0.10 | 46.13 1.60 | 50.87 0.99 | 57.67 0.41 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 5-shot | 69.34 0.46 | 17.70 1.38 | 28.18 1.83 | 48.79 0.27 | 90.93 0.53 | 63.50 0.10 | 45.84 0.96 | 100.00 0.00 | |||||||||||
+| SaigaLlama3 8B | 0-shot | 61.30 0.87 | 32.41 0.46 | 42.41 0.60 | 49.31 0.34 | 76.27 0.53 | 59.90 0.42 | 51.15 0.51 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 5-shot | 56.94 1.27 | 68.05 5.06 | 61.86 1.14 | 51.63 1.49 | 40.00 8.00 | 44.69 6.27 | 53.28 2.56 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 0.4 | 0-shot | 63.04 0.15 | 20.00 0.92 | 30.36 1.06 | 48.22 0.13 | 86.40 0.53 | 61.89 0.03 | 46.12 0.52 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 5-shot | 62.12 4.24 | 22.07 4.14 | 32.16 4.54 | 48.08 0.10 | 83.73 4.80 | 61.04 1.30 | 46.60 1.62 | 100.00 0.00 | |||||||||||
+| Vikhr 7B IT 5.4 | 0-shot | 59.18 0.00 | 33.33 0.00 | 42.65 0.00 | 48.67 0.00 | 73.33 0.00 | 58.51 0.00 | 50.58 0.00 | 100.00 0.00 | |||||||||||
+| VikhrGemma 2B IT | 5-shot | 51.38 2.77 | 24.14 9.20 | 32.22 8.23 | 46.04 1.46 | 74.40 5.87 | 56.71 0.87 | 41.01 3.23 | 99.88 0.25 | |||||||||||
+| VikhrGemma 2B IT | 0-shot | 60.17 0.35 | 17.01 0.46 | 26.52 0.53 | 48.11 0.16 | 81.60 0.53 | 60.53 0.28 | 29.02 0.09 | 93.70 0.25 |
+Appendix D Psychologists Evaluation Setup
+During the LLM generation interpretation, psychologists used the following rating scale for explanations, given by LLM: 1 - completely erroneous, 2 - mostly erroneous, 3 - partially erroneous, partially correct, 4 - mostly correct, 5 - completely correct.
+On this scale, the psychologists rated LLM explanations as 2.84 out of 5 with Fleiss’ , which shows fair inter-annotator agreement.
+To better categorize errors in LLM explanations, psychologists also provided a list of the most common errors in each generation with detailed explanations of the error type.
+This list is provided in Table˜17.
+| Error name | Description | Example |
+| 1.
+Tautology | The final part of the explanation repeats the initial part in other words without any proof.
+| This text indicates that the author has depression, as he describes his thoughts and feelings, which are characteristic of depressive disorders.
+|
+| 2.
+Groundless generalization | The patient’s experience from the text is defined as a sign of depression, while this experience on its own, without the context, is not specific to depression.
+| The desire to return to the lost state of happiness and happy life, which is also a sign of depression.
+|
+| 3.
+False conclusion | The false inference is derived from the text statement.
+| The author also mentions that his parents, who he considers to be positive, were unable to correct his behavior.
+In the original text there is no signs of parents intention to correct authors behaviour.
+|
+| 4.
+Confabulation | The explanation contains evidence, which is not mentioned in the context.
+| The text describes a deep dissatisfaction with the world, people and their actions included in the text, and also expresses a desire to be happy not in the text and enjoy the little things not in the text.
+|
+| 5.
+Distortion of medical understanding of depression | Misconception about depression.
+| She also expresses a desire to … plan long-term plans, which also indicates the presence of a depressive disorder.
+The long-term planning cannot be considered as a sign of depression.
+|
+| 6.
+Incompleteness of selected signs of depression | Of the several significant signs of depression only one or two signs (mostly minor) are highlighted.
+| Thank you for this test, so that I could repeat all this to myself once again and think about the rope.
+This significant sign does not mentioned in the explanation.
+|
+We also conducted an additional experiment with a clinically informed prompt to explore how the additional information from psychologists can affect the quality of explanations.
+To do so, we asked psychologists to select parts of the text that can serve as depression markers for texts that were used in a 5-shot prompt for the DE dataset.
+After we performed a 5-shot evaluation of the best model with a modified 5-shot prompt, which now includes depression markers along with the target label.
+The results are presented in Table˜18.
+| Corpus | Model | Mode |
+F1-healthy |
+F1-pathology | F1-macro | ||||||||
+| DE | 0-shot | Vikhr 7B IT 5.4 | 71.43 0.00 | 5.56 0.00 | 10.31 0.00 | 19.05 0.00 | 90.91 0.00 | 31.50 0.00 | 20.90 0.00 | ||||||||
+| 5-shot | Vikhr 7B IT 5.4 | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | |||||||||
+| 5-shot clinically informed | Vikhr 7B IT 5.4 | 87.64 0.00 | 86.67 0.00 | 87.15 0.00 | 47.83 0.00 | 50.00 0.00 | 48.89 0.00 | 68.02 0.00 |
+The adapted prompt leads to a slightly better F1-macro score; however, such a prompt requires qualified psychologists to label several texts for each investigated dataset.
+These results provide an interesting direction for future work and show that the overall quality of LLMs in mental disorders detection tasks can be improved with the help of psychologists.
+To further investigate the quality of model explanations with clinically informed prompts, we asked psychologists to conduct the evaluation of results as in Section˜5.4.
+During evaluation, the psychologists set the average score to 2.47 out of 5, which is lower than for prompting without a clinically informed prompt.
+Due to the significant difference in the explanations of LLM, the mistakes were categorized into other groups, namely: (1) confabulation, (2) undifferentiation (inability to select symptomatic parts of the text), and (3) incompleteness of explanation.
+With a clinically informed prompt, each of the explanations contains at least one of the mentioned mistakes.
+The most common types of mistakes are confabulation and undifferentiation, which are found in 47% of the explanations.
+The incompleteness of the explanation occurs in 40% of the explanations.
+Overall, these results show that even with the clinically informed prompt, modern LLMs are unable to generate explanations that meet the requirements of clinicians.
+Appendix E Feature Importance Ablation
+To further deepen the explanation analysis, we conducted feature ablation for the best LLM on the DE dataset.
+For this purpose, we extracted the top-3 words by their importance on target label generation (and on full explanation generation) for each text in the test set.
+The words were extracted with the Feature Ablation method from the Captum framework Kokhlikyan et al.
+(2020), which calculates feature ablation based on differences in predictions with and without features.
+The list of unique words with the biggest mean feature importance is shown at Figure˜1 (importance for target label generation) and Figure˜2 (importance for full explanation generation).
+For both target class and explanation generation, the most important features in texts with predicted pathology class contain a significant amount of words with negative meanings (such as "disappointing", "negative", "bitter", "angry", "disgusting", "darkness", etc.), as well as words with direct pathology description ("depression", "fear").
+The detailed analysis from the trained clinicians reveals that most of the features with high importance for depression class are connected with fear, suffering, and unhealthy conditions.
+On the other hand, the most important features of a healthy class contain mostly positive semantics, such as "humanity", "unselfishness", and "kind".
+Appendix F Disaggregated Results for DSM Dataset
+The scores for the Depression-Social Media (DSM) dataset were aggregated based on the obtained BDI questionnaire scores.
+To further investigate how these results vary between subgroups with various scores within one group (inside sub-splits of healthy or pathology class), we provided Tables˜19 and 20 with disaggregated scores.
+As one can see, for the healthy class, the F1-healthy scores are mostly consistent for all models, with the exception of RuBioRoBERTa, which has a significantly lower F1-healthy score for the split with BDI Score from 3 to 6.
+For pathology class, F1-pathology scores vary more significantly, which can be described with a bigger variability of initial scores, especially for the split with scores from 34 to 63.
+| Mode | Model | BDI Score 0-3 | BDI Score 3-6 | BDI Score 7-10 | ||||||||||||||||
+F1-healthy |
+F1-healthy |
+F1-healthy | ||||||||||||
+| 5-shot | SaigaLlama3 8B | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 100.00 0.00 | 6.25 0.00 | 11.76 0.00 | ||||||||||
+| 5-shot | Vikhr 7B IT 0.4 | 100.00 0.00 | 83.33 0.00 | 90.91 0.00 | 100.00 0.00 | 80.00 0.00 | 88.89 0.00 | 100.00 0.00 | 81.25 0.00 | 89.66 0.00 | ||||||||||
+| 5-shot MMLU | SaigaLlama3 8B | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 100.00 0.00 | 6.25 0.00 | 11.76 0.00 | ||||||||||
+| 5-shot MMLU | Vikhr 7B IT 5.4 | 100.00 0.00 | 83.33 0.00 | 90.91 0.00 | 100.00 0.00 | 80.00 0.00 | 88.89 0.00 | 100.00 0.00 | 81.25 0.00 | 89.66 0.00 | ||||||||||
+| LoRA | Qwen2 7B IT | 100.00 0.00 | 80.56 11.45 | 88.79 7.00 | 100.00 0.00 | 83.33 21.34 | 89.15 15.15 | 100.00 0.00 | 65.62 10.05 | 78.80 7.34 | ||||||||||
+| SFT | RuBioRoBERTa | 100.00 0.00 | 86.11 11.45 | 92.12 6.78 | 100.00 0.00 | 50.00 10.00 | 66.07 8.93 | 100.00 0.00 | 58.33 4.66 | 73.57 3.79 | ||||||||||
+| Mode | Model | BDI Score 30-34 | BDI Score 34-63 | ||||||||||
+F1-pathology |
+F1-pathology | ||||||||
+| 5-shot | SaigaLlama3 8B | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | ||||||
+| 5-shot | Vikhr 7B IT 0.4 | 100.00 0.00 | 52.00 4.00 | 68.33 3.33 | 100.00 0.00 | 37.50 0.00 | 54.55 0.00 | ||||||
+| 5-shot MMLU | SaigaLlama3 8B | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | 100.00 0.00 | ||||||
+| 5-shot MMLU | Vikhr 7B IT 5.4 | 100.00 0.00 | 50.00 0.00 | 66.67 0.00 | 100.00 0.00 | 37.50 0.00 | 54.55 0.00 | ||||||
+| LoRA | Qwen2 7B IT | 100.00 0.00 | 58.33 14.62 | 72.59 11.90 | 100.00 0.00 | 41.67 11.79 | 57.87 11.50 | ||||||
+| SFT | RuBioRoBERTa | 100.00 0.00 | 55.00 5.00 | 70.83 4.17 | 100.00 0.00 | 27.08 4.66 | 42.42 5.42 | ||||||
+Appendix G Textual Examples
+For a better understanding of the investigated task, we presented several samples from the DE dataset - one per class.
+The samples were paraphrased and anonymized, and then translated into English.
+The examples are presented in Table˜21.
+| Label | Text (In English) |
+| Pathology | Hello!
+My name is NAME, I am already AGE years old.
+I live with my family: my husband and our wonderful daughter.
+My profession is a midwife in a women’s clinic, which I truly love.
+The team in our department is small, but we are all very close to each other.
+If any problems arise, we solve them very quickly.
+Until recently, I was an active person, with a large circle of friends who delighted me with various entertainments: going to the theater, watching movies, walking around the city or just meeting for a cup of coffee.
+Communicating with them brought me great pleasure.
+However, recently I began to distance myself from everyone, both at work and in my personal environment.
+This was the result of my fear, which turned out to be much more powerful than I expected.
+Fear takes away not only emotional strength, but also physical.
+At present, it seems to me that it is better to remain alone, avoiding meetings with other people.
+Although this makes my life less bright, I cannot cope with this condition yet.
+Also, I used to be into bead embroidery, but now I can’t even bring myself to do a simple task.
+Luckily, we have our cute house cat, who helps me cope with stress and is a great antidepressant.
+This is my current lifestyle.
+|
+| Healthy | When I was a child, I accidentally found out that there were people who didn’t like me.
+They were my classmates, and I was very upset when I tried to be friends with them, but my attempt only increased the hostility.
+I tried to attract their attention with gifts, invitations to visit and other ways, but each new attempt ended in failure and even more trouble.
+Then I realized that no matter how you behave, there will always be people who don’t like you, sometimes for no apparent reason.
+But it is important to understand that this is normal and you shouldn’t try to live up to the expectations of others.
+You should value yourself for who you are.
+Many years have passed since then, and now I have many friends and acquaintances, some leave, and new ones come.
+Each person is unique and beautiful in their own way.
+I especially liked the expression: “Each person is a small cosmos.” It is really profound.
+Inside each person there is a whole universe consisting of his life experience, mistakes, disappointments, joys, defeats and small victories.
+If a person allows others to open up, it can be incredibly beautiful and exciting.
+In a world where we encounter many people every day, I want to believe that everyone respects each other, despite the fact that everyone may not like them.
+Even those we find unattractive may be dear to someone else.
+So there is no need to worry about not being loved, because we can find a common language and fill each other with love for life.
+We are all small universes living together in one big world, trying to get to know each other every day.
+|